### v2

# The $q$-Schur algebras in type $D$, I: Fundamental multiplication formulas

Abstract.
By embedding the Hecke algebra $\check{H}_{\boldsymbol{q}}$ of type $D$ into the Hecke algebra $H_{{\boldsymbol{q}},1}$ of type $B$ with unequal parameters $({\boldsymbol{q}},1)$, the ${\boldsymbol{q}}$-Schur algebras $S^{\kappa}_{\boldsymbol{q}}(n,r)$ of type $D$ is naturally defined as the endomorphism algebra of the tensor space with the $\check{H}_{\boldsymbol{q}}$-action restricted from the $H_{{\boldsymbol{q}},1}$-action that defines the $({\boldsymbol{q}},1)$-Schur algebra $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ of type $B$. We investigate the algebras $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ and $S^{\kappa}_{\boldsymbol{q}}(n,r)$ both algebraically and geometrically and describe their natural bases, dimension formulas and weight idempotents. Most importantly, we use the geometrically derived two sets of the fundamental multiplication formulas in $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ to derive multi-sets (9 sets in total!) of the fundamental multiplication formulas in $S^{\kappa}_{\boldsymbol{q}}(n,r)$.
Key words and phrases:
Hecke algebra, $q$-Schur algebra, partial flag variety, orthogonal group, special orthogonal group
1991 Mathematics Subject Classification:
16T20, 17B37, 20C08, 20C33, 20G43
Supported partially by the UNSW Science FRG (J. Du) and ARC DP220102861 (Daniel Chan)
Supported partially by NSF grant DMS1801915 and Simons Travel Support for Mathematicians.
1. Introduction
The $q$-Schur algebras of type $A$ originated from two sources. One originates from quantum group theory and the other from representation theory of finite general linear groups. In [17], Jimbo used $q$-Schur algebras to describe the (quantum) Schur–Weyl duality, while in [9] Dipper–James used them to investigate representations of finite general linear groups in non-defining characteristic. Thus, $q$-Schur algebras play a central role in connecting representations of quantum linear groups with those of finite general linear groups.
Although one naturally expected thirty years ago the existence of such a connection for other finite classical groups, the Schur–Weyl–Brauer duality and its quantum version for classical types other than $A$ cannot be used to establish such a connection. This is because the Brauer algebras and their quantum analogs, the so-called BMW algebras (see [7, Th. 10.2.5] and the references therein), have a very different representation theory from those of Hecke algebras of type $B/C/D$. Thus, we hypothesized the existence of a ‘Schur-Weyl-Hecke duality’ for other classical types. This puzzle has finally been solved after thirty years!
In their study of quantum symmetric pairs and their canonical basis theory, Bao–Wang [2] discovered a new Schur–Weyl duality which is called here the Schur–Weyl–Hecke duality. The Schur–Weyl–Hecke duality of type $B/C$ involves the twin $i$-quantum groups ${\bf U}^{\jmath}(n)$ and ${\bf U}^{\imath}(n)$ and the Hecke algebras of types $B$ and $C$. These $i$-quantum groups appear in the quantum symmetric pairs of type AIII with no black nodes in their Satake diagrams. The $q$-Schur algebras involved have numerous applications; see, e.g., [4, 12, 18, 20, 21, 23]. Note that, for the type $D$ case, a commuting action of the $i$-quantum groups ${\bf U}^{\jmath}(n)$, ${\bf U}^{\imath}(n)$ and Hecke algebras of type $D$ on tensor spaces was first observed by Ehrig and Stroppel [13] in their study of maximal parabolic category ${\mathcal{O}}$/isotropic Grassmannians of $\mathfrak{so}_{2n}$. See also Bao’s work [1]. It is natural to ask if there is a Schur–Weyl–Hecke duality in type $D$?
In [14], Z. Fan and the second-named author investigated a natural geometric setting of the $q$-Schur algebras of type $D$ (see [19, Appendix] for the associated algebraic setting) and their limiting algebra $\mathcal{K}$, following the work [4]. Moreover, certain $i$-quantum groups are also introduced so that the type $D$ $q$-Schur algebras are their homomorphic images. However, it has been noticed that multiplication formulas in the type $D$ $q$-Schur algebras discussed in [14, §4] are incomplete (see [14, Prop. 4.3.2]). In other words, more cases need to be considered in the determination of the natural basis when the algebra is defined by $q$-permutation modules of the type $D$ Hecke algebra (see [19, Appendix]). Hence, the dimension formulas need to be revised and multiplication formulas about these new basis elements should be computed as well.
The aim of the paper is to fill these gaps in [14] and [19] and to set up a foundation for a possible solution to the aforementioned question. It is interesting to note that, on the way to achieve the goal, we discovered more. The strategy here is new as we develop two parallel theories: a type $B$ theory in unequal parameters and a type $D$ theory. We then use a splitting process via either weight idempotents or geometric splitting to link the two.
We now describe it in more details. We start with the algebraic approach developed in [11], where the Hecke algebra $\check{H}_{\boldsymbol{q}}$ of type $D_{r}$ is embedded in the Hecke algebra $H_{{\boldsymbol{q}},1}$ of type $B_{r}$ in unequal parameters. Interestingly, there is a geometric construction for $H_{q,1}$, where $q$ is a prime power, using complete isotropic flag variety over the even-order orthogonal group ${\text{\rm O}}_{2r}(q)$. Then, we may present the $({\boldsymbol{q}},1)$-Schur algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ of type $B$ both algebraically, using $q$-permutation modules, and geometrically, using pairs of isotropic partial flag variety over ${\text{\rm O}}_{2r}(q)$ and their associated convolution algebra (see Theorem 3.10). In this way, we may define the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ of type $D$ by restriction from $H_{{\boldsymbol{q}},1}$ to $\check{H}_{\boldsymbol{q}}$ and identify it with the algebra $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined in [19] (see Proposition 4.2). Next, using a criterion for splitting an ${\text{\rm O}}_{2r}(q)$-orbit into two ${\text{\rm SO}}_{2r}(q)$-orbits, a new parametrization of ${\text{\rm SO}}_{2r}(q)$-orbits is obtained (see (4.8.3)). We then introduce the natural basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$, indexed by (labelled) double cosets in the type $D$ Weyl group, and the defining orbital basis for the convolution algebra over ${\text{\rm SO}}_{2r}(q)$, indexed by ${\text{\rm SO}}_{2r}(q)$-orbits. Theorem 4.15 gives a geometric setting for $S_{q}^{\textsc{d}}(n,r)$ via a basis correspondence. This precise basis correspondence (or identification) in loc. cit. is the key to finding the fundamental multiplication formulas in these generic $q$-Schur algebras of type $D$ (Theorems 6.26.4 and 7.17.3) through a splitting process of the corresponding formulas of type $B$ given in Theorem 5.2.
More precisely, we summarize the fundamental multiplication formulas in the paper as follows. (See notational list below.)
Theorem 1.1.
Let $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ and $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ be the ${\boldsymbol{q}}$-Schur algebras of type $B$ and $D$, respectively, over the polynomial ring $\mathbb{Z}[{\boldsymbol{q}}]$ with their respective bases $\{e_{A}\mid A\in\Xi\}$ and $\{\phi_{\check{\mathbb{A}}}\mid{\check{\mathbb{A}}}\in\check{\Xi}\}$. Suppose $B,C\in\Xi$ such that $B-E_{h,h+1}^{\theta}$ and $C-E_{h+1,h}^{\theta}$ $(h\in[1,n])$ are diagonal and, in $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$,
$e_{B}e_{A}=\delta_{\text{\rm co}(B),\text{\rm ro}(A)}\sum_{1\leq p\leq N}g_{h,A,p}e_{{}_{h}\!A_{p}},\quad e_{C}e_{A}=\delta_{\text{\rm co}(C),\text{\rm ro}(A)}\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}e_{{}_{h}\!A_{\overline{p}}},$
for some $g_{h,A,p},g^{\prime}_{{}_{h,A,\overline{p}}}\in\mathbb{Z}[{\boldsymbol{q}}]$. (See (3.7.1) for the notation ${}_{h}\!A_{p},{}_{h}\!A_{\overline{p}}$.) If ${\check{\mathbb{A}}}$, ${\check{\mathbb{B}}}$, and $\check{\mathbb{C}}$ in $\check{\Xi}$ are obtained by a certain splitting process from $A,B,$ and $C$, respectively, then the structure constants for $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ appearing in $\phi_{{\check{\mathbb{B}}}}\phi_{\check{\mathbb{A}}}$ (resp., $\phi_{\check{\mathbb{C}}}\phi_{\check{\mathbb{A}}}$) are some of the $g_{h,A,p}$ (resp. $g^{\prime}_{{}_{h,A,\overline{p}}}$ or $\frac{1}{2}g^{\prime}_{{}_{h,A,\overline{p}}}$).
We organize the paper as follows. In Section 2, we first introduce, following [10, (3.3.2)(2)] and [11, Rem. 2.6(1)], the algebraic definition for the type $D$ $q$-Schur algebras $S^{\kappa}_{{\boldsymbol{q}}}(n,r)$ associated with (or covering) the type $B$ algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ in unequal parameters. We then reinterpret $S^{\kappa}_{{\boldsymbol{q}}}(n,r)$ as the algebra $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined by Heck algebras of type $D$ and their $q$-permutation modules (see [19, (A.3.3)]). In Section 3, we present the geometric definition of $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$, largely following [4, 14]. In Section 4, we develop a precise labelling for those ${\text{\rm O}}_{2r}(q)$-orbits which split into two ${\text{\rm SO}}_{2r}(q)$-orbits. This allows us to label the natural basis for $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ both algebraically in terms of double cosets and geometrically in terms of ${\text{\rm SO}}_{2r}(q)$-orbits. Finally, in Section 5, the multiplication formulas in the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ of type $B$ (with unequal parameters) will be derived geometrically via convolution products as in [4]. In the last Sections 6 and 7, fundamental multiplication formulas for the ${\boldsymbol{q}}$-Schur algebra of type $D$ are derived using either weight idempotents or orbit splitting techniques. This includes especially those missing ones in [14] (see Remarks 4.16 and 7.4).
In a forthcoming paper, we will use the multiplication formulas discovered here to seek the existence of the Schur–Weyl–Hecke duality in type $D$.
Some notation hints: There are two parallel notations for types $B$ and $D$. The following table lists some of them for reference. Let ${\epsilon}\in\{+,-\}$.
(1) (2) (3) (4) (5) (6) (7) (8) (9)
[…]
where
  • (1)
    Weyl groups $W=W^{\textsc{b}}$ and ${\check{W}}=W^{\textsc{d}}$;
  • (2)
    Index sets of orbits of parabolic subgroups
    or isotropic flags in ${\mathcal{X}}$;
  • (3)
    Hecke algebras in unequal/equal parameter;
  • (4)
    ${\boldsymbol{q}}$-Schur algebras;
  • (5)
    Natural bases for ${\boldsymbol{q}}$-Schur algebras;
  • (6)
    Index sets for natural bases;
  • (7)
    Finite orthogonal/special orthogonal groups $G(q)={\text{\rm O}}_{2r}(q)/{\check{G}}(q)={\text{\rm SO}}_{2r}(q)$;
  • (8)
    $G(q)$-/${\check{G}}(q)$-orbits in ${\mathcal{X}}\times{\mathcal{X}}$, set of pairs of $n$-step isotropic flags;
  • (9)
    Weight functions defined in (2.3.2) and (4.12.5).
Throughout the paper, let $n$ be a positive integer and let $N=2n+1$.
Let ${\mathcal{A}}:=\mathbb{Z}[{\boldsymbol{q}}]$ be the integral polynomial ring in indeterminate ${\boldsymbol{q}}$. For any integer $m\geq 0$, define the Gaussian integer $\left[\!\left[m\right]\!\right]=\frac{{\boldsymbol{q}}^{m}-1}{{\boldsymbol{q}}-1}$.
For integers $a<b$, let $[a,b]:=\{x\in\mathbb{Z}\mid a\leq x\leq b\}$.
1.1. Acknowledgements
The second-named author is grateful to Chun-Ju Lai for pointing out a gap in the parametrization of ${\text{\rm SO}}_{2r}(q)$-orbits in [14]; see [19, (A.4.2)], Remark 4.16, and (4.8.3).
The authors express their gratitude to the University of New South Wales and the University at Buffalo for their hospitality during the preparation of this paper. Additionally, the third author acknowledges the financial support provided by the China Scholarship Council.
We also thank Professor Weiqiang Wang for his comments on an early version of the paper.
We sincerely thank the referee for valuable comments and suggestions.
2. The ${\boldsymbol{q}}$-Schur algebras of type $D$: an algebraic setting
We start with the definitions of the Weyl groups, Hecke algebras, and ${\boldsymbol{q}}$-Schur algebres of type $B/D$. By identifying $W$ and its parabolic subroups with certain fixed-point subgroups of the symmetric group ${\mathfrak{S}}_{2r}$, we then index the natural (or double coset) basis for the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$ of type $B$ by a certain matrix set $\Xi_{N,2r}$. This matrix set will soon be used in Section 3 to label the ${\text{\rm O}}_{2r}(q)$-orbits in ${\mathcal{X}}\times{\mathcal{X}}$ for an $n$-step isotropic flag variety ${\mathcal{X}}$.
2.1. The Weyl groups $W/{\check{W}}$, Hecke algebras $H_{{\boldsymbol{q}},1}/\check{H}_{\boldsymbol{q}}$, and ${\boldsymbol{q}}$-Schur algebras $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)/S^{\kappa}_{\boldsymbol{q}}(n,r)$
Let $W=W^{\textsc{b}}=W(B_{r})$ ($r\geq 2$) and ${\check{W}}=W^{\textsc{d}}=W(D_{r})$ ($r\geq 4$) denote the Weyl groups of type $B_{r}$ and $D_{r}$ associated with the Dynkin diagrams:
$B_{r}$:123r-1r
[…]
Then we have the following sequence of three Coxeter systems:
$({\bar{W}},\bar{S})\leq({\check{W}},{\check{S}})\leq(W,S),$ (2.0.1)
where $S=\{s_{1},\cdots,s_{r-1},s_{r}\}$, ${\check{S}}=\{s_{1},\cdots,s_{r-1},\varsigma_{r}\}$ with $\varsigma_{r}:=s_{r}s_{r-1}s_{r}$, and $\bar{S}=\{s_{1},\cdots,s_{r-1}\}$. Note that $\bar{W}$ is isomorphic to the symmetric group ${\mathfrak{S}}_{r}$ on $r$ letters.
Let $t_{r}=s_{r}$, $t_{r-i}=s_{r-i}t_{r-i+1}s_{r-i}$, $u_{i}=t_{r}t_{r-i}$ for $11\leq i\leq r-$ 1, and let $C=\langle t_{1},\cdots,t_{r}\rangle$ and $\check{C}=\langle u_{1},\cdots,u_{r-1}\rangle$. Then, we have $W=C\rtimes\bar{W}$ and ${\check{W}}=\check{C}\rtimes\bar{W}$. Moreover, we often identify ${\bar{W}}$ with ${\mathfrak{S}}_{r}$ in the sequel.
Let $\ell$ (resp. ${\check{\ell}}$) be the length function on ${{W}}$ (resp. ${\check{W}}$) with respect to $S$ (resp. ${\check{S}}$) and $n_{r}$ the function giving the number of $s_{r}$ in a reduced expression of an element of ${{W}}$ (cf. [10, (2.1.1)]). Clearly,
${\check{W}}=\{w\in W\mid n_{r}(w)\in 2\mathbb{N}\}.$ (2.0.2)
The inner automorphism ${\mathfrak{f}}:w\mapsto s_{r}ws_{r}$ on ${{W}}$ induces by restriction a graph automorphism ${\mathfrak{f}}$ on ${\check{W}}$ which flips the Coxeter graph:
${\mathfrak{f}}:{\check{W}}\longrightarrow{\check{W}},\;\;w\longmapsto s_{r}ws_{r}.$ (2.0.3)
Note that ${\mathfrak{f}}$ fixes each element of $\bar{S}$, and interchanges the two parabolic copies of ${\mathfrak{S}}_{r}$ in ${\check{W}}$.
Let $H_{{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}}=H_{{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}}^{\textsc{b}}$ be the 2-parameter Hecke algebra of type $B_{r}$, over the polynomial ring ${\mathcal{A}}^{\prime}=\mathbb{Z}[{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}]$ in two variables, generated by $T_{1},\ldots,T_{r-1},T_{r}$ ($T_{i}=T_{s_{i}}$) which satisfy the relations:
$(1)$ $T_{i}T_{j}=T_{j}T_{i},\;\;|i-j|\geq 2,$ (2.0.4)
$(2)$ $T_{j}T_{j+1}T_{j}=T_{j+1}T_{j}T_{j+1}\;\;(1\leq j<r-1);$
$(3)$ $T_{i}^{2}=({\boldsymbol{q}}-1)T_{i}+{\boldsymbol{q}}\;\;(1\leq i<r),$
$(4)$ $T_{r-1}T_{r}T_{r-1}T_{r}=T_{r}T_{r-1}T_{r}T_{r-1},\;\;$
$(5)$ $T_{r}^{2}=({\boldsymbol{q}}^{\prime}-1)T_{r}+{\boldsymbol{q}}^{\prime}.$
Putting $T_{w}=T_{i_{1}}\cdots T_{i_{l}}$, where $w=s_{i_{1}}\cdots s_{i_{l}}\in W$ is a reduced expression, we obtain a basis $\{T_{w}\mid w\in W\}$ for $H_{{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}}$.
Let $H_{{\boldsymbol{q}},1}=H_{{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}}\otimes_{{\mathcal{A}}^{\prime}}\mathbb{Z}[{\boldsymbol{q}}]$ be the algebra obtained by base change from ${\mathcal{A}}^{\prime}$ to ${\mathcal{A}}:=\mathbb{Z}[{\boldsymbol{q}}]$ under the specialization of ${\boldsymbol{q}}^{\prime}=1$. Then $H_{{\boldsymbol{q}},1}$ is presented by $T_{1},\ldots,T_{r-1},T_{r}$ and relations
[…]
The $\mathbb{Z}[{\boldsymbol{q}}]$-algebra $H_{{\boldsymbol{q}},1}$ contains a subalgebra $\check{H}_{\boldsymbol{q}}:=H^{\textsc{d}}_{\boldsymbol{q}}$ which is isomorphic to the Hecke algebra of type $D_{r}$. Here $\check{H}_{\boldsymbol{q}}$ is generated by $T_{1},\ldots,T_{r-1},T_{\varsigma}:=T_{r}T_{r-1}T_{r}$ with defining relations (1)–(3) in (2.0.4) together with:
$T_{\varsigma}^{2}=({\boldsymbol{q}}-1)T_{\varsigma}+{\boldsymbol{q}},\;\;T_{\varsigma}T_{i}=T_{i}T_{\varsigma}\;(i\neq r-2),\;\;T_{r-2}T_{\varsigma}T_{r-2}=T_{\varsigma}T_{r-2}T_{\varsigma}.$
We will use the same notation as above describing basis $\{T_{w}\mid w\in W\}$ for $H_{{\boldsymbol{q}},1}$ and basis $\{T_{w}\mid w\in{\check{W}}\}$ for $\check{H}_{\boldsymbol{q}}$. We often write $H=H_{{\boldsymbol{q}},1}$ and $\check{H}=H_{\boldsymbol{q}}$ for notational simplicity (see §4.1).
Both $H_{{\boldsymbol{q}},1}$ and $\check{H}_{\boldsymbol{q}}$ admit an algebra anti-involution (anti-automorphism of order 2):
$\tau:T_{w}\longmapsto T_{w^{-1}}.$ (2.0.6)
We use the following indexing set to label (standard) parabolic subgroups of $W$:
$\Lambda(n+1,r)=\{\lambda=(\lambda_{1},\ldots,\lambda_{n},\lambda_{n+1})\in\mathbb{N}^{n+1}\mid\lambda_{1}+\cdots+\lambda_{n+1}=r\}.$
For $\lambda=(\lambda_{1},\ldots,\lambda_{n},\lambda_{n+1})\in\Lambda(n+1,r)$, define the parabolic subgroup of $W$ associated with $\lambda$ as the subgroup
$W_{\lambda}=\langle S\backslash\{s_{\lambda_{1}+\cdots+\lambda_{i}}\mid i\in[1,n]\}\rangle,$ (2.0.7)
and let $x_{\lambda}=\sum_{w\in W_{\lambda}}T_{w}$. The $x_{\lambda}H_{{\boldsymbol{q}},1}$ is a right $H_{{\boldsymbol{q}},1}$-module with basis $x_{\lambda}T_{d}$, for $d\in{\mathcal{D}}_{\lambda}$, where ${\mathcal{D}}_{\lambda}$ is the set of distinguished right coset representative of $W_{\lambda}$. Let ${\mathcal{D}}_{\lambda\mu}:={\mathcal{D}}_{\lambda}^{-1}\cap{\mathcal{D}}_{\mu}$, for $\lambda,\mu\in\Lambda(n+1,r)$. This is the set of distinguished representatives of $W_{\lambda}$-$W_{\mu}$ double cosets.
Following [10, (3.3.2)(2)] (or [3] in the context of $i$-quantum groups), we define
$S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)=S^{\jmath}_{{\boldsymbol{q}},1}(n,r)=\text{\rm End}_{H_{{\boldsymbol{q}},1}}\Big{(}\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H_{{\boldsymbol{q}},1}\Big{)}.$ (2.0.8)
Remark 2.1.
The double subscripts in the notation $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$ indicate this unequal parameter setting. It is different from the equal-parameter algebra $\mathbf{S}^{\jmath}$ or ${\mathcal{S}}^{\jmath}_{\mathcal{Z}}(n,r)$ investigated in [4] or in [12].
Following [11, Rem. 2.6(1)], we define, by restriction, the ${\boldsymbol{q}}$-Schur algebra of type $D$ [^1] An alternative definition is given in Definition 2.10:
$S_{{\boldsymbol{q}}}^{\kappa}(n,r)=\text{\rm End}_{\check{H}_{\boldsymbol{q}}}\big{(}\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H_{{\boldsymbol{q}},1}|_{\check{H}_{\boldsymbol{q}}}\big{)}$ (2.1.1)
Let ${\boldsymbol{q}}=\upsilon^{2}$ and ${\mathcal{Z}}:=\mathbb{Z}[\upsilon,\upsilon^{-1}]$. By base change to ${\mathcal{Z}}$, we obtain the Hecke algebras ${\mathcal{H}}_{\upsilon,1},\check{\mathcal{H}}_{\upsilon}$ and the $\upsilon$-Schur algebras ${\mathcal{S}}_{\upsilon,1}^{\textsc{b}}(n,r)$ and ${\mathcal{S}}_{\upsilon}^{\kappa}(n,r)$:
${\mathcal{H}}_{\upsilon,1}:=H_{{\boldsymbol{q}},1}\otimes{\mathcal{Z}},\;\;\check{\mathcal{H}}_{\upsilon}:=\check{H}_{\boldsymbol{q}}\otimes{\mathcal{Z}},\;\;{\mathcal{S}}_{\upsilon,1}^{\textsc{b}}(n,r):=S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\otimes_{{\mathcal{A}}}{\mathcal{Z}},\;\;{\mathcal{S}}_{\upsilon}^{\kappa}(n,r)=S_{{\boldsymbol{q}}}^{\kappa}(n,r)\otimes_{{\mathcal{A}}}{\mathcal{Z}}.$ (2.1.2)
Though we will not consider these ${\mathcal{Z}}$-algebras in the rest of the paper, they play important roles for the study of the Schur–Weyl–Hecke duality mentioned in the introduction.
2.2. Identifying $W,W_{\lambda}$ with certain fixed-point subgroups of ${\mathfrak{S}}_{2r}$
We may embed $W$ into ${\mathfrak{S}}_{2r}$ as a fixed-point subgroup $({\mathfrak{S}}_{2r})^{\theta}$, where $\theta$ is the involution
$\theta:{\mathfrak{S}}_{2r}\longrightarrow{\mathfrak{S}}_{2r},(i,j)\longmapsto(2r+1-i,2r+1-j).$
More precisely, the map
$\iota:W$ $\longrightarrow{\mathfrak{S}}_{2r},$ (2.1.3)
$s_{i}$ $\longmapsto(i,i+1)(2r+1-i,2r-i)\;\;(1\leq i<r),\quad s_{r}\longmapsto(r,r+1)$
is an injective group homomorphism. By identifying $W(=W^{\textsc{b}})$ with its image $\iota(W)$, $W$ acts on the set $[1,2r]$ and, hence, on the power set of $[1,2r]$. With this identification, $W$ consists of permutations
$\begin{pmatrix}1&2&\cdots&r&r+1&\cdots&2r-1&2r\cr i_{1}&i_{2}&\cdots&i_{r}&i_{r+1}&\cdots&i_{2r-1}&i_{2r}\end{pmatrix},$ (2.1.4)
where $i_{j}+i_{2r+1-j}=2r+1$ for all $j$. In other words, $W$ is generated by
$s_{i}=\sigma_{i}\sigma_{2r-i}\;(1\leq i\leq r-1),\;\;s_{r}=\sigma_{r},$
where $\sigma_{i}=(i,i+1)$, for $i=1,2,\ldots,2r-1$ form the Coxeter generators of ${\mathfrak{S}}_{2r}$.
Moreover, the restriction of $\iota$ to $\check{W}(=W^{\textsc{d}})$ has the image which is the subgroup generated by
$s_{i}=\sigma_{i}\sigma_{2r-i},\;\;(1\leq i<r),\quad\varsigma=\sigma_{r}\sigma_{r-1}\sigma_{r+1}\sigma_{r}=(r-1,r+1)(r,r+2).$ (2.1.5)
We often identify ${\check{W}}$ as a subgroup of ${\mathfrak{S}}_{2r}$ via $\iota$. Thus, ${\check{W}}$ acts on the power set of $[1,2r]$.
For any composition $\mu=(\mu_{1},\ldots,\mu_{N})$ of $2r$, define the associated partial sum sequence
${\widetilde{\mu}}=({\widetilde{\mu}}_{1},{\widetilde{\mu}}_{2},\cdots,{\widetilde{\mu}}_{N}),\;\;\text{ where }{\widetilde{\mu}}_{i}=\mu_{1}+\cdots+\mu_{i}\;({\widetilde{\mu}}_{0}=0).$ (2.1.6)
Let ${\mathfrak{S}}_{\mu}$ be the Young subgroup of ${\mathfrak{S}}_{2r}$ associated with $\mu$. This is the stabilizer subgroup of the standard Young tabloid [^2] If we identify $\mu$ with its Young diagram consisting of $\mu_{1}$ boxes in row 1, and $\mu_{2}$ boxes in row 2 and so on, then a $\mu$-tableau is obtained by filling the numbers $1,2,\ldots,2r$ into boxed. A $\mu$-tabloid is a row-equivalent class of $\mu$-tableaux. $R^{\mu}=R^{\mu}_{1}\times\cdots\times R_{N}^{\mu}$, where $R^{\mu}_{i}:=[\widetilde{\mu}_{i-1}+1,\widetilde{\mu}_{i-1}+\mu_{i}]$, for all $i=1,2,\ldots,N$. In other words,
${\mathfrak{S}}_{\mu}=\bigcap_{i=1}^{N}\text{Stab}_{{\mathfrak{S}}_{2r}}(R_{i}^{\mu}),$ (2.1.7)
consisting of permutations which stabilize each of the subsets $R^{\mu}_{i}$.
For $\lambda\in\Lambda(n+1,r)$, let
$\widehat{\lambda}:=(\lambda_{1},\cdots,\lambda_{n},2\lambda_{n+1},\lambda_{n},\cdots,\lambda_{1}).$
This defines an injective map
^$\widehat{\ }:\Lambda(n+1,r)\longrightarrow\Lambda(2n+1,2r),\;\;\lambda\longmapsto\widehat{\lambda}.$ (2.1.8)
Let $\widehat{\Lambda}(n+1,r)$ denote its image. We now describe the parabolic subgroup $W_{\lambda}$ defined in (2.0.7) in terms of stabilizers of the subsets $R_{i}^{\lambda}:=R_{i}^{\widehat{\lambda}}$, for $1\leq i\leq{\mathsf{m}}$, where ${\mathsf{m}}={\mathsf{m}}(\lambda)$ denotes the maximal index $i$ satisfying $\lambda_{i}\neq 0$.
Let
${\mathcal{D}}(n,r)$ :=$:=\{(\lambda,d,\mu)\mid\lambda,\mu\in\Lambda(n+1,r),d\in{\mathcal{D}}_{\lambda,\mu}\},$ (2.1.9)
$\Xi_{N,2r}$ :=$:=\{A=(a_{i,j})\in\text{Mat}_{N,N}(\mathbb{N})\mid a_{i,j}=a_{N+1-i,N+1-j},\sum_{i,j}a_{i,j}=2r\}.$
Lemma 2.2.
Let $\lambda,\mu\in\Lambda(n+1,r)$, $w\in{\mathfrak{S}}_{2r}$, $d\in{\mathcal{D}}_{\lambda,\mu}$, and ${\widehat{d}}=\iota(d)$ (see (2.1.3)).
  • (1)
    $W_{\lambda}=({\mathfrak{S}}_{{\widehat{\lambda}}})^{\theta}=\bigcap_{i=1}^{\mathsf{m}}\text{\rm Stab}_{W}(R^{\lambda}_{i}).$
  • (2)
    Suppose the double coset ${\mathfrak{S}}_{\widehat{\lambda}}w{\mathfrak{S}}_{\widehat{\mu}}$ defines a $(2n+1)\times(2n+1)$ matrix $A=(a_{i,j})$ over $\mathbb{N}$ whose entries sum to $2r$. Then ${\mathfrak{S}}_{\widehat{\lambda}}w{\mathfrak{S}}_{\widehat{\mu}}$ is stabilized by $\theta$ if and only if $a_{i,j}=a_{N+1-i,N+1-j}$ for all $i,j\in[1,N]$ ($N=2n+1$). In particular, for ${\widehat{d}}=\iota(d)\in{\mathfrak{S}}_{2r}$, $\theta$ stabilizes ${\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}}{\mathfrak{S}}_{\widehat{\mu}}$ and the double coset $W_{\lambda}dW_{\mu}=({\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}}{\mathfrak{S}}_{\widehat{\mu}})^{\theta}$.
  • (3)
    There is a bijection
    $\mathfrak{d}:{\mathcal{D}}(n,r)\longrightarrow\Xi_{N,2r},\;\;(\lambda,d,\mu)\longmapsto A=(|R^{\widehat{\lambda}}_{i}\cap dR^{\widehat{\mu}}_{j}|).$ (2.2.1)
    Moreover, we have $s_{r}\in W_{\lambda}\cap{}^{d}W_{\mu}$ if and only if $|R^{\widehat{\lambda}}_{n+1}\cap dR^{\widehat{\mu}}_{n+1}|\geq 2$.
Proof.
The equality $W_{\lambda}={\mathfrak{S}}_{\widehat{\lambda}}\cap W=({\mathfrak{S}}_{{\widehat{\lambda}}})^{\theta}$ is clear from the definition of the embedding $\iota$. To see the second equality, we observe
$w(R^{\lambda}_{i})=R^{\lambda}_{i},\text{ for all }1\leq i\leq{\mathsf{m}}\iff w(R^{\widehat{\lambda}}_{i})=R^{\widehat{\lambda}}_{i},w(R^{{\widehat{\lambda}}}_{2n+2-i})=R^{{\widehat{\lambda}}}_{2n+2-i},\text{ for all }1\leq i\leq{\mathsf{m}}.$
Thus, we have
$\bigcap_{i=1}^{\mathsf{m}}\text{\rm Stab}_{W}(R^{\lambda}_{i})=\bigcap_{i=1}^{N}\text{Stab}_{W}(R_{i}^{{\widehat{\lambda}}_{i}})=\bigcap_{i=1}^{N}\big{(}\text{Stab}_{{\mathfrak{S}}_{2r}}(R_{i}^{{\widehat{\lambda}}_{i}})\cap W\big{)}={\mathfrak{S}}_{{\widehat{\lambda}}}\cap W=({\mathfrak{S}}_{{\widehat{\lambda}}})^{\theta}.$
proving (1).
The remaining but the final assertion is well-known; see, e.g., [12, Lem. 3.1]. We now prove the last assertion.
For $\lambda\in\Lambda(n+1,r)$ and $w\in W$, let $R^{\widehat{\lambda}}=R^{\widehat{\lambda}}_{1}\times\cdots\times R^{\widehat{\lambda}}_{N}$ and $w.R^{\widehat{\lambda}}:=w(R^{\widehat{\lambda}}_{1})\times\cdots\times w(R^{\widehat{\lambda}}_{N})$, which are called ${\widehat{\lambda}}$-tabloids. This defines a $W$-set $X_{N}:=\{w.R^{\widehat{\lambda}}\mid\lambda\in\Lambda(n+1,r),w\in W\}$, consisting of all ${\widehat{\lambda}}$-tabloids. This $W$-action on $X_{N}$ induces a diagonal action of $W$ on $X_{N}\times X_{N}$ such that each orbit has a representative $(R^{\widehat{\lambda}},dR^{\widehat{\mu}})$ for some $\lambda,\mu\in\Lambda(n+1,r)$ and $d\in{\mathcal{D}}_{\lambda,\mu}$. For $R,R^{\prime}\in X_{N}$ with $R=R_{1}\times\ldots\times R_{N}$, $R^{\prime}=R^{\prime}_{1}\times\ldots\times R^{\prime}_{N}$, let
$R\wedge R^{\prime}=(R_{1}\cap R^{\prime}_{1})\times(R_{1}\cap R^{\prime}_{2})\times\cdots\times(R_{1}\cap R^{\prime}_{N})\times\ldots\times(R_{N}\cap R^{\prime}_{1})\times(R_{N}\cap R^{\prime}_{2})\times\cdots\times(R_{N}\cap R^{\prime}_{N}).$
Define $X_{N^{2}}:=\{R\wedge R^{\prime}\mid R,R^{\prime}\in X_{N}\}$. Then the map
$\wedge:X_{N}\times X_{N}\longrightarrow X_{N^{2}},\;\;(R,R^{\prime})\longmapsto R\wedge R^{\prime}$
is bijective and $W$-equivariant. Thus, $\text{Stab}_{W}(R,R^{\prime})=\text{Stab}_{W}(R\wedge R^{\prime})=\prod_{i,j\in[1,N]}\text{Stab}_{W}(R_{i}\cap R^{\prime}_{j}).$ Hence, for $(\lambda,d,\mu)\in{\mathcal{D}}(n,r)$, $R=R^{\widehat{\lambda}}$ and $R^{\prime}=dR^{\widehat{\mu}}$, since $\text{Stab}_{W}(R,R^{\prime})=W_{\lambda}\cap{}^{d}W_{\mu}$, it follows that
$s_{r}\in W_{\lambda}\cap{}^{d}W_{\mu}$ $\iff s_{r}\in\text{Stab}_{W}(R^{\widehat{\lambda}}\wedge dR^{\widehat{\mu}})$
$\iff s_{r}\in\text{Stab}_{W}(R^{\widehat{\lambda}}_{n+1}\wedge dR^{\widehat{\mu}}_{n+1})$
$\iff r,r+1\in R^{\widehat{\lambda}}_{n+1}\cap dR^{\widehat{\mu}}_{n+1}\iff|R^{\widehat{\lambda}}_{n+1}\cap dR^{\widehat{\mu}}_{n+1}|\geq 2,$
as desired. ∎
Remark 2.3.
In the proof above, Lemma 2.2(1) implies that we may replace the $W$-set $X_{N}$ of ${\widehat{\lambda}}$-tabloids by its truncated version, the set of all $\lambda$-tabloids $w.R^{\lambda}=w(R^{\lambda}_{1})\times\cdots\times w(R^{\lambda}_{{\mathsf{m}}(\lambda)})$ ($\lambda\in\Lambda(n+1,r),w\in W$), where $R^{\lambda}_{1}\sqcup\cdots\sqcup R^{\lambda}_{\mathsf{m}}=[1,r+\lambda_{n+1}]$.
For a matrix $A=(a_{i,j})\in\Xi_{N,2r}$, form two sequences via its rows and columns and their associated partial sum sequences as defined in (2.1.6):
${\mathbb{r}}(A)$ $=(a_{1,1},a_{1,2},\cdots\cdots,a_{1,N},\cdots,a_{N,1},a_{N,2},\cdots,a_{N,N}),\quad\widetilde{\mathbb{r}}_{A}=(\widetilde{a}^{r}_{1,1},\cdots\cdots,\widetilde{a}^{r}_{N,N})$ (2.3.1)
${\mathbb{c}}(A)$ $=(a_{1,1},a_{2,1},\cdots\cdots,a_{N,1},\cdots,a_{1,N},a_{2,N},\cdots,a_{N,N}),\quad\widetilde{\mathbb{c}}_{A}=(\widetilde{a}^{c}_{1,1},\cdots\cdots,\widetilde{a}^{c}_{N,N}).$
If $A=\mathfrak{d}(\lambda,d,\mu)$, then we set $d_{A}=d$, ${\widehat{\lambda}}=\text{\rm ro}(A)$, and ${\widehat{\mu}}=\text{\rm co}(A)$. Thus, we have
$\text{\rm ro}(A)=\Big{(}\sum_{j=1}^{N}a_{1,j},\sum_{j=1}^{N}a_{2,j},\ldots,\sum_{j=1}^{N}a_{N,j}\Big{)},\quad\text{\rm co}(A)=\Big{(}\sum_{i=1}^{N}a_{i,1},\sum_{i=1}^{N}a_{i,2},\ldots,\sum_{i=1}^{N}a_{i,N}\Big{)}.$ (2.3.2)
Note that ${\mathfrak{S}}_{\widehat{\lambda}}\cap{\widehat{d}_{A}}{\mathfrak{S}}_{\widehat{\mu}}{\widehat{d}_{A}}^{-1}={\mathfrak{S}}_{{\bf r}(A)}$ and ${\widehat{d}_{A}}^{-1}{\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}_{A}}\cap{\mathfrak{S}}_{\widehat{\mu}}={\mathfrak{S}}_{{\bf c}(A)}$.
By deleting row $n+1$ and column $n+1$ (i.e., the central row and column of $A$), $A$ is divided into four $n\times n$ submatrices: the upper/lower right corner matrices $A^{\text{\normalsize$\llcorner$}}/A_{\text{\normalsize$\ulcorner$}}$ and the upper/lower left corner matrices $A^{\text{\normalsize$\lrcorner$}}/A_{\text{\normalsize$\urcorner$}}$. The following number will be very useful as it tells when the element $d_{A}$ is in ${\check{W}}$. Let
$|A^{\text{\normalsize$\llcorner$}}|=\sum_{i,j\in[1,n]}a_{i,n+1+j}=|A_{\text{\normalsize$\urcorner$}}|$ (2.3.3)
be the entry sum of the upper right corner matrix.
Corollary 2.4.
Maintain the notation introduced above. For $A=(a_{i,j})\in\Xi_{N,2r}$, $d_{A}\in W$ is the following permutation $\left(\begin{smallmatrix}\widetilde{a}^{c}_{i-1,j}+1&\widetilde{a}^{c}_{i-1,j}+2&\cdots&\widetilde{a}^{c}_{i-1,j}+a_{i,j}\\ \widetilde{a}^{r}_{i,j-1}+1&\widetilde{a}^{r}_{i,j-1}+2&\cdots&\widetilde{a}^{r}_{i,j-1}+a_{i,j}\end{smallmatrix}\right)$ in ${\mathfrak{S}}_{2r}$, for all $(i,j)$ with $a_{i,j}>0$. Moreover, we have $d_{A}\in{\check{W}}$ if and only if $|A^{\text{\normalsize$\llcorner$}}|\in 2\mathbb{N}$.
Proof.
The first assertion follows from [8, Ex. 8.2], Thus, if $d_{A}$ permutes $(1,2,\ldots,2r)$ to $(i_{1},i_{2},\ldots,i_{2r})$ as displayed in (2.1.4), then $\{i_{1},i_{2},\ldots,i_{r}\}\cap\{r+1,r+2,\ldots,2r\}$ has cardinality $|A^{\text{\normalsize$\llcorner$}}|$. However, by [10, (2.1.1)], $|A^{\text{\normalsize$\llcorner$}}|=n_{r}(d_{A})$. Now, the last assertion follows from (2.0.2). ∎
Each triple $(\lambda,d,\mu)\in{\mathcal{D}}(n,r)$ defines a natural basis element in $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$. We denote this element by $e_{A}$ if $A=\mathfrak{d}(\lambda,d,\mu)$. By definition,
$e_{A}(x_{\nu}h)=\delta_{\mu,\nu}\Big{(}\sum_{w\in W_{\lambda}dW_{\mu}}T_{w}\Big{)}h.$ (2.4.1)
Proposition 2.5.
The set $\{e_{A}\mid A\in\Xi_{2n+1,2r}\}$ forms a basis for $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$ which we call its natural basis.
2.3. Parabolic subgroups of ${\check{W}}$ as stabilizers of tabloids associated with signed compositions
We now extend Lemma 2.2 above to (standard) parabolic subgroups of ${\check{W}}$; compare [19]. Let $\lambda\in\Lambda(n+1,r)$ have maximal index ${\mathsf{m}}$.
The $\lambda_{n+1}=0$ case. If $\lambda_{n+1}=0$, then ${\mathsf{m}}\leq n$. Define the signed composition $\lambda^{+}$ and $\lambda^{-}$ by setting
${\check{R}}^{\lambda^{+}}_{i}$ $={\check{R}}^{\lambda^{-}}_{i}=R_{i}^{\lambda}=[\lambda_{i-1}+1,\lambda_{i-1}+\lambda_{i}]\;\;\text{for }1\leq i\leq{\mathsf{m}}-1,$ (2.5.1)
${\check{R}}_{\mathsf{m}}^{\lambda^{+}}$ $=[\widetilde{\lambda}_{{\mathsf{m}}-1}+1,\widetilde{\lambda}_{{\mathsf{m}}-1}+\lambda_{\mathsf{m}}]=\{\widetilde{\lambda}_{{\mathsf{m}}-1}+1,\cdots,r-1,r\},$
${\check{R}}_{\mathsf{m}}^{\lambda^{-}}$ $=[\widetilde{\lambda}_{{\mathsf{m}}-1}+1,\widetilde{\lambda}_{{\mathsf{m}}}-1]\cup\{r+1\}=\{\widetilde{\lambda}_{{\mathsf{m}}-1}+1,\cdots,r-1,r+1\}.$
Form the tabloids
${\check{R}}^{\lambda^{+}}:={\check{R}}^{\lambda^{+}}_{1}\sqcup\cdots\sqcup{\check{R}}^{\lambda^{+}}_{\mathsf{m}},\qquad{\check{R}}^{\lambda^{-}}:={\check{R}}^{\lambda^{-}}_{1}\sqcup\cdots\sqcup{\check{R}}^{\lambda^{-}}_{\mathsf{m}}.$
Define, for ${\epsilon}\in\{+,-\}$,
${\check{W}}_{\lambda^{\epsilon}}:={\text{\rm Stab}}_{\check{W}}({\check{R}}_{1}^{\lambda^{\epsilon}})\cap{\text{\rm Stab}}_{\check{W}}({\check{R}}_{2}^{\lambda^{\epsilon}})\cap\cdots\cap{\text{\rm Stab}}_{\check{W}}({\check{R}}_{\mathsf{m}}^{\lambda^{\epsilon}}).$
The $\lambda_{n+1}\neq 0$ case. If $\lambda_{n+1}\neq 0$, then define ${\lambda^{\bullet}}$ by setting
${\check{R}}_{i}^{{\lambda^{\bullet}}}$ =$=\begin{cases}[\widetilde{\lambda}_{i-1}+1,\widetilde{\lambda}_{i-1}+\lambda_{i}]=R^{\lambda}_{i},&\text{ if }1\leq i\leq n,\\ [\widetilde{\lambda}_{n}+1,\widetilde{\lambda}_{n}+2\lambda_{n+1}],&\text{ if }i=n+1.\end{cases}$ (2.5.2)
Form the tabloid ${\check{R}}^{{\lambda^{\bullet}}}:={\check{R}}^{{\lambda^{\bullet}}}_{1}\sqcup\cdots\sqcup{\check{R}}^{{\lambda^{\bullet}}}_{n+1}$ and define the subgroup of ${\check{W}}$:
${\check{W}}_{{\lambda^{\bullet}}}:=\mathop{\bigcap}\limits_{i=1}^{n+1}{\rm Stab}_{\check{W}}({\check{R}}_{i}^{{\lambda^{\bullet}}}).$
The following result is established in [19, Lem. A.2.1] with different indexing.
Lemma 2.6.
For $\lambda\in\Lambda(n+1,r)$ and ${\epsilon}\in\{{\bullet},+,-\}$, the subgroups ${\check{W}}_{\lambda^{\epsilon}}$ are parabolic subgroups of ${\check{W}}=\langle s_{1},\ldots,s_{r-1},\varsigma_{r}\rangle$ with $\varsigma_{r}=s_{r}s_{r-1}s_{r}$. More precisely, we have
  • (1)
    If $\lambda_{n+1}=0$ (so ${\mathsf{m}}\leq n$) and $\lambda_{\mathsf{m}}>1$, then
    ${\check{W}}_{\lambda^{+}}$ $=\langle\check{S}-\{s_{{\widetilde{\lambda}_{1}}},s_{\widetilde{\lambda}_{2}},\cdots,s_{\widetilde{\lambda}_{{\mathsf{m}}-1}},\varsigma_{r}\}\rangle,$ (2.6.1)
    ${\check{W}}_{\lambda^{-}}$ $=\langle\check{S}-\{s_{\widetilde{\lambda}_{1}},s_{\widetilde{\lambda}_{2}},\cdots,s_{\widetilde{\lambda}_{{\mathsf{m}}-1}},s_{r-1}\}\rangle.$
  • (2)
    If $\lambda_{n+1}=0$ and $\lambda_{\mathsf{m}}=1$, then ${\check{R}}_{\mathsf{m}}^{\lambda^{+}}=\{r\}$ and ${\check{R}}_{\mathsf{m}}^{\lambda^{-}}=\{r+1\}$ and
    ${\check{W}}_{\lambda^{+}}={\check{W}}_{\lambda^{-}}=\langle{\check{S}}-\{s_{\widetilde{\lambda}_{1}},s_{\widetilde{\lambda}_{2}},\cdots,s_{\widetilde{\lambda}_{{\mathsf{m}}-1}},s_{r-1},\varsigma_{r}\}\rangle.$ (2.6.2)
  • (3)
    If $\lambda_{n+1}\neq 0$, then
    ${\check{W}}_{\lambda^{\bullet}}=\begin{cases}\langle\check{S}-\{s_{\widetilde{\lambda}_{1}},s_{\widetilde{\lambda}_{2}},\cdots,s_{\widetilde{\lambda}_{n-1}},s_{r-1},\varsigma_{r}\}\rangle,&\text{ if }\lambda_{n+1}=1,\\ \langle\check{S}-\{s_{\widetilde{\lambda}_{1}},s_{\widetilde{\lambda}_{2}},\cdots,s_{\widetilde{\lambda}_{n-1}}\}\rangle,&\text{ if }\lambda_{n+1}>1.\end{cases}$ (2.6.3)
Moreover, every parabolic subgroup of ${\check{W}}$ is one of the forms above.
For the parabolic subgroups $W_{\lambda}$ of $W^{\textsc{b}}$ defined in (2.0.7), we naturally define
$\check{W}_{\lambda}=\begin{cases}W_{\lambda},&\text{ if }\lambda_{n+1}=0;\\ W_{\lambda}\cap\check{W},&\text{ if }\lambda_{n+1}\neq 0.\end{cases}$ (2.6.4)
We may describe the parabolic subgroups of $W^{\textsc{d}}$ given in Lemma 2.6 in terms of the subgroups ${\check{W}}_{\lambda}$. Recall the flipping graph automorphism ${\mathfrak{f}}$ in (2.0.3).
Corollary 2.7.
Let $\lambda\in\Lambda(n+1,r)$.
  1. (1)
    If $\lambda_{n+1}=0$ and $\lambda_{\mathsf{m}}>1$, then ${\check{W}}_{\lambda^{+}}=W_{\lambda}$ and ${\check{W}}_{\lambda^{-}}=s_{r}{\check{W}}_{\lambda}s_{r}={\mathfrak{f}}({\check{W}}_{\lambda}).$
  2. (2)
    If $\lambda_{n+1}=0$ and $\lambda_{\mathsf{m}}=1$, then ${\check{W}}_{\lambda^{+}}=W_{\lambda}={\mathfrak{f}}(W_{\lambda})={\check{W}}_{\lambda^{-}}.$
  3. (3)
    If $\lambda_{n+1}\neq 0$, then ${\check{W}}_{\lambda^{\bullet}}={\check{W}}_{\lambda}={\mathfrak{f}}({\check{W}}_{\lambda^{\bullet}})$.
In particular, ${\check{W}}_{\lambda}$ is a parabolic subgroup of ${\check{W}}$. [^3] The reader should not be confused the notation ${\check{W}}_{\lambda}$ with ${\check{W}}_{\lambda^{\epsilon}}$. Note that ${\check{W}}_{\lambda}={\check{W}}_{\lambda^{+}}$ or ${\check{W}}_{\lambda}={\check{W}}_{\lambda^{\bullet}}$, depending on $\lambda_{n+1}=0$ or not. Moreover, for every parabolic subgroup $\langle J\rangle$ of $\check{W}$ ($J\subseteq\check{S}$), there exists a $\lambda\in\Lambda(n+1,r)$ such that $\langle J\rangle=\check{W}_{\lambda}$ or $\langle J\rangle=s_{r}\check{W}_{\lambda}s_{r}={\mathfrak{f}}({\check{W}}_{\lambda})$.
Remarks 2.8.
(1) The case (2) above shows that distinct tabloids may have the same stabilizer subgroups. This phenomenon is similar to the type $A$ fact that distinct compositions may define the same Young subgroup of the symmetric group.
(2) Consider the parabolic subgroup $W_{\mu}=\langle s_{1},\ldots,s_{r-1}\rangle$ of $W$, where $\mu={(r,0^{n})}$. This is also a parabolic subgroup of $\check{W}$. That is $W_{\mu}=\check{W}_{\mu}$. Clearly, $\varsigma=s_{r}s_{r-1}s_{r}\in{\mathcal{D}}_{\mu,\mu}$, but $s_{r-1}=s_{r}\varsigma s_{r}$ is not distinguished in the double coset $\check{W}_{\mu}(s_{r}ds_{r})\check{W}_{\mu}=\check{W}_{\mu}$.
2.4. The ${\boldsymbol{q}}$-Schur algebras of type $D$
We are now ready to introduce the $q$-Schur algebra of type $D$ in terms of ${\boldsymbol{q}}$-permutation modules of $\check{H}_{\boldsymbol{q}}$ (see [19, (A.3.3)] and compare [11, Def. 2.5]). Let
$\Lambda^{\textsc{d}}=\Lambda^{\textsc{d}}(n,r):=\Lambda^{\bullet}(n+1,r)\sqcup\Lambda^{\circ}_{+}(n+1,r)\sqcup\Lambda^{\circ}_{-}(n+1,r),$ (2.8.1)
where $\Lambda^{\bullet}(n+1,r):=\{\lambda^{\bullet}\mid\lambda\in\Lambda(n+1,r),\lambda_{n+1}\neq 0\}$ and $\Lambda^{\circ}_{\pm}(n+1,r):=\{\lambda^{\pm}\mid\lambda\in\Lambda^{\circ}(n+1,r),\lambda_{n+1}=0\}$.
Remark 2.9.
We remark that, if ${\mathcal{P}}({\check{S}})$ denotes the power set of ${\check{S}}$ and $n\geq r$, then the following map is surjective:
$\Lambda^{\textsc{d}}(n,r)\longrightarrow{\mathcal{P}}({\check{S}}),\;\;\alpha\longmapsto{\check{S}}\cap{\check{W}}_{\alpha}.$
As promised in footnote 1, the following is an alternative definition for the ${\boldsymbol{q}}$-Schur algebra of type $D$.
Definition 2.10.
[[19]] For $\alpha\in\Lambda^{\textsc{d}}$, let ${\check{x}}_{\alpha}=\sum_{w\in{\check{W}}_{\alpha}}T_{w}$ and define the ${\boldsymbol{q}}$-Schur (or $\upsilon$-Schur) algebra of type $D$:
$S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ $:=\text{\rm End}_{\check{H}_{\boldsymbol{q}}}\Big{(}\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}{\check{H}_{\boldsymbol{q}}}\Big{)}\quad(\text{over }{\mathcal{A}}=\mathbb{Z}[{\boldsymbol{q}}]);$
${\mathcal{S}}_{\upsilon}^{\textsc{d}}(n,r)$ $:=\text{\rm End}_{{\check{\mathcal{H}}}_{\upsilon}}\Big{(}\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}{\check{\mathcal{H}}}_{\upsilon}\Big{)}\quad(\text{over }{\mathcal{Z}}=\mathbb{Z}[\upsilon,\upsilon^{-1}]).$
The algebras defined above involve only $q$-permutation modules of type $D$ Hecke algebras. They are isomorphic to the algebras defined in (2.1.1) and (2.1.2). This will be proved in Proposition 4.2.
For every $\alpha\in\Lambda^{\textsc{d}}(n,r)$, the identity map $1_{\alpha}$ on ${\check{x}}_{\alpha}\check{H}$ extends to an idempotent, called an weight idempotent in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. We record the following.
Lemma 2.11.
The set $\{1_{\alpha}\mid\alpha\in\Lambda^{\textsc{d}}(n,r)\}$ forms a set of idempotents such that $1=\sum_{\alpha\in\Lambda^{\textsc{d}}(n,r)}1_{\alpha}$ is the identity element of $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$.
We will introduce the natural basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ and discuss its geometric intepretation in Section 4.
3. A geometric setting for the ${\boldsymbol{q}}$-Schur algebras $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$
In [4], a geometric framework is introduced for the $q$-Schur algebra $S^{\jmath}_{q}(n,r)$ of type $B$ in equal parameter. We now modify their construction using $(2n+1)$-step filtrations for the $2r$-dimensional space $\mathbb{F}_{q}^{2r}$ (see [14]) and the action by the orthogonal group ${\text{\rm O}}_{2r}(q)$. This results in a $q$-Schur algebra in unequal parameters. Restricting the action to the subgroup ${\text{\rm SO}}_{2r}(q)$ leads to a geometric setting for the $q$-Schur algebra of type $D$; see Theorems 3.10 and 4.15 in the next two sections.
3.1. A geometric setting for the endomorphism algebra of a permutation module
Let $G$ be a finite group acting on a finite set $X$ (not necessarily faithfully or transitively), i.e., $X$ is a $G$-set, and let $\mathcal{R}$ be a commutative ring with 1. Let $\mathfrak{F}_{G}(X,{\mathcal{R}})$ be the ${\mathcal{R}}$-free module of all functions $f:X\to{\mathcal{R}}$ which are constant on the orbits of $G$. If $X^{\prime}$ is another $G$-set and $\pi:X\to X^{\prime}$ is a $G$-set map, then $\pi$ induces ${\mathcal{R}}$-module maps $\pi_{!}:\mathfrak{F}_{G}(X,{\mathcal{R}})\to\mathfrak{F}_{G}(X^{\prime},{\mathcal{R}})$ and $\pi^{*}:\mathfrak{F}_{G}(X^{\prime},{\mathcal{R}})\to\mathfrak{F}_{G}(X,{\mathcal{R}})$ defined by
$(\pi_{!}f)(x^{\prime})$ $=\sum_{x\in\pi^{-1}(x^{\prime})}f(x),\;\;\text{ for all }f\in\mathfrak{F}_{G}(X,{\mathcal{R}}),x^{\prime}\in X^{\prime};$
$(\pi^{*}f^{\prime})(x)$ $=f^{\prime}(\pi(x)),\;\;\text{ for all }f^{\prime}\in\mathfrak{F}_{G}(X^{\prime},{\mathcal{R}}),x\in X.$
The $G$-action is extended diagonally to $X\times X$. Consider three projection maps $\pi_{i,j}:X\times X\times X\to X\times X$ sending $(x_{1},x_{2},x_{3})$ to $(x_{i},x_{j})$ for $(i,j)\in\{(1,2),(2,3),(1,3)\}$. Then $\mathfrak{F}_{G}(X\times X,{\mathcal{R}})$ becomes an associative algebra with multiplication defined by $f*f^{\prime}=(\pi_{1,3})_{!}(\pi_{1,2}^{*}f\cdot\pi_{2,3}^{*}f^{\prime})$. In other words,
$(f*f^{\prime})(x,y)=\sum_{z\in X}f(x,z)f^{\prime}(z,y),\;\;\text{ for all }(x,y)\in X\times X.$
Let $\Omega$ be the set of all $G$-orbits on $X\times X$. For each orbit ${\mathcal{O}}$, there is an associated (characteristic) orbital function $f_{\mathcal{O}}$ defined by setting
$f_{\mathcal{O}}(x,y)=\begin{cases}1,&\text{if }(x,y)\in{\mathcal{O}};\\ 0,&\text{elsewhere.}\end{cases}$ (3.0.1)
Then $\mathfrak{F}_{G}(X\times X,{\mathcal{R}})$ has a basis $f_{\mathcal{O}},{\mathcal{O}}\in\Omega$, and
$f_{\mathcal{O}}*f_{{\mathcal{O}}^{\prime}}=\sum_{{\mathcal{O}}^{\prime\prime}}c_{{\mathcal{O}},{\mathcal{O}}^{\prime},{\mathcal{O}}^{\prime\prime}}f_{{\mathcal{O}}^{\prime\prime}},\;\;\text{ where }\;c_{{\mathcal{O}},{\mathcal{O}}^{\prime},{\mathcal{O}}^{\prime\prime}}=\#\{z\in X\mid(x,z)\in{\mathcal{O}},(z,y)\in{\mathcal{O}}^{\prime}\},$ (3.0.2)
for $(x,y)\in{\mathcal{O}}^{\prime\prime}$. (This number is independent of the selection of $(x,y)\in{\mathcal{O}}^{\prime\prime}$.) The identity element $\bf 1$ is the sum of the idempotent functions $f_{\mathcal{O}}$ associated with ${\mathcal{O}}$ of the form $G.(x,x)$.
Let ${\mathcal{R}}X$ be the free ${\mathcal{R}}$-module with basis $X$. Then ${\mathcal{R}}X$ becomes an ${\mathcal{R}}G$-module via the $G$-action on $X$. The following result is well-known.
Lemma 3.1.
Let $G$ be a finite group.
(1) If $H,K$ are subgroups of $G$, then there is a bijection between the set $H\backslash G/K$ of all double cosets $HgK$ ($g\in G$) and the set $\Omega$ of all $G$-orbits in the $G$-set $G/H\times G/K$ with diagonal action.
(2) For a finite $G$-set $X$, the endomorphism algebra $\text{\rm End}_{\mathcal{R}G}(\mathcal{R}X)^{\text{\rm op}}$ is isomorphic to $\mathfrak{F}_{G}(X\times X,{\mathcal{R}})$.
3.2. Finite orthogonal groups and flag varieties
We now look at the case, where $G=O_{2r}(q)$ or $SO_{2r}(q)$ is the finite orthogonal or special orthogonal group, and $X=\mathcal{F}_{n,r}^{\jmath}$ is the set of $n$-step filtrations (or partial flag variety) of isotropic subspaces.
Let $\mathbb{F}_{q}$ be a finite field of $q$ elements and of odd characteristic. [^4] The odd characteristic guarantees that ${\text{\rm O}}_{2r}(q)={\text{\rm O}}^{+}_{2r}(q)$ and ${\text{\rm SO}}_{2r}(q)={\text{\rm SO}}^{+}_{2r}(q)$. See [15, §1.3.16] and [15, Th. 1.7.8, §1.7.9]. On the $2r$-dimensional vector space $\mathbb{F}_{q}^{2r}$, we fix a nondegenerate symmetric bilinear form $\langle-,-\rangle_{J}$ whose associated $2r\times 2r$-matrix is
[…]
(thus, ${\mathsf{J}}_{ij}=\delta_{i,2r+1-j}$ for any $1\leq i,j\leq 2r$). In other words, for $x,y\in\mathbb{F}_{q}^{2r}$,
$\langle x,y\rangle_{{\mathsf{J}}}=x^{t}{\mathsf{J}}y=x_{1}y_{2r}+x_{2}y_{2r-1}+\cdots+x_{2r}y_{1}.$ (3.1.1)
The orthogonal group is defined by
$\mathrm{O}_{2r}(q):=\{g\in{\text{\rm GL}}_{2r}(q)\ |\ J=g^{t}Jg\}=\{g\in{\text{\rm GL}}_{2r}(q)\mid\langle gx,gy\rangle_{\mathsf{J}}=\langle x,y\rangle_{\mathsf{J}}\}.$
This is a group with a split BN-pair, where $B=\widehat{B}(q)\cap{\text{\rm O}}_{2r}(q)$ with $\widehat{B}(q)$ being the Borel subgroup of ${\text{\rm GL}}_{2r}(q)$ consisting of upper triangular matrices and $N=\{\dot{w}\mid w\in W\}$ (see, e.g., [15, p.80]), where
$\dot{w}$ is the permutation matrix $\dot{w}=(\delta_{k,i_{l}})$ if $w$ is given as in (2.1.4), sending $k$ to $i_{k}$. (3.1.2)
We also have the special orthogonal group
$\mathrm{SO}_{2r}(q):=\{g\in{\text{\rm O}}_{2r}(q)\mid{\text{\rm\,det\,}}(g)=1\}.$
Note that $\mathrm{SO}_{2r}(q)$ is also a group with a split BN-pair; see, e.g., [15, Th. 1.7.8].
By convention, $W^{\perp}$ stands for the orthogonal complement of a vector subspace $W\subset\mathbb{F}_{q}^{2r}$ with respect to the bilinear form (3.1.1). We call a vector subspace $W\subset\mathbb{F}_{q}^{2r}$ isotropic if $W\subseteq W^{\perp}$. Note that a maximal isotropic subspace $M$ in $\mathbb{F}_{q}^{2r}$ has dimension $r$ and $M^{\perp}=M$. Also, the restriction of $\langle\;\quad\rangle_{\mathsf{J}}$ to $W^{\perp}/W$ is a form (3.1.1) with rank $2r^{\prime}$, where $r^{\prime}=r-\text{\rm dim\,}W$.
We fix a positive integer $n$ and let $N=2n+1$. For a sequence of $n$ isotropic subspaces
$W_{1}\subseteq W_{2}\subseteq\cdots\subseteq W_{n},$ (3.1.3)
we may extend it to an $N$-step filtration
[…]
by setting $F_{i}=W_{i}$ and $F_{n+i}=W_{N-n-i}^{\perp}$, for all $1\leq i\leq n$. We call (3.1.4) an $n$-step isotropic flag (extended to an $N$-step flag) and (3.1.3) a pure $n$-step isotropic flag. If $n=r$ and dim $\text{\rm dim\,}W_{i}=i$, then (3.1.3) is called a complete isotropic flag. In this case, since $F_{r}=F_{r+1}$, (3.1.4) becomes a ($2r$-step) complete isotropic flag after reindexing.
Consider the following finite sets:
  • The $n$-step isotropic flag variety is the set
    $\mathcal{F}_{n,r}^{\jmath}={\mathcal{F}}_{n}^{\jmath}(\mathbb{F}_{q}^{2r})=\{0=F_{0}\subseteq F_{1}\subseteq F_{2}\subseteq\cdots\subseteq F_{N}\equiv\mathbb{F}_{q}^{2r}\mid F_{i}\subseteq F_{i}^{\perp},F_{i}^{\perp}=F_{N-i},\forall i\in[1,n]\},$
    consisting of all $n$-step isotropic flags (extended to $N$-steps) in $\mathbb{F}_{q}^{2r}$.
  • The complete isotropic flag variety is the set
    $\mathcal{B}_{r}={\mathcal{B}}_{r}(\mathbb{F}_{q}^{2r})=\{0\subset\mathsf{F}_{1}\subset\mathsf{F}_{2}\subset\cdots\subset\mathsf{F}_{2r}\equiv\mathbb{F}_{q}^{2r}\mid\mathsf{F}_{i}\subseteq\mathsf{F}_{i}^{\perp},\text{dim}\mathsf{F}_{i}=i,\mathsf{F}_{i}^{\perp}=\mathsf{F}_{2r-i},\forall i\in[1,r]\},$
    consisting of all complete ($r$-step) isotropic flags in $\mathbb{F}_{q}^{2r}$.
We often set ${\mathcal{X}}:=\mathcal{F}_{n,r}^{\jmath}$ and ${\mathcal{B}}:=\mathcal{B}_{r}$ for notational simplicity. Both admit naturally an ${\mathrm{O}}_{2r}(q)$-action from the left. Moreover, ${\mathrm{O}}_{2r}(q)$ acts transitively on $\mathcal{B}$.
3.3. Parabolic subgroups of ${\text{\rm O}}_{2r}(q)$
Consider the natural basis $e_{1},e_{2},\ldots,e_{2r}$ for $\mathbb{F}_{q}^{2r}$ and form the (standard) pure isotropic complete flag $\mathsf{F}^{s}$ with
$\mathsf{F}^{s}_{1}=\langle e_{1}\rangle,\mathsf{F}^{s}_{2}=\langle e_{1},e_{2}\rangle,\ldots,\mathsf{F}^{s}_{r}=\langle e_{1},e_{2},\ldots,e_{r}\rangle.$
Then we have $B(q)=\text{Stab}_{G(q)}(\mathsf{F}^{s})$. Moreover, the mapping $G(q)\to\mathcal{B}_{r}$ sending $g$ to $gF$ induces a bijection $G(q)/B(q)\to\mathcal{B}_{r}$ between the two $G$-sets. Thus, by Lemma 3.1(1), the $G$-orbits in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$ correspond bijectively to the set of double coset $B\setminus G/B$. Hence, the Bruhat decomposition, every $G$-orbit $G.(\mathsf{F},\mathsf{F}^{\prime})$ in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$ has the from
$G.(\mathsf{F},\mathsf{F}^{\prime})=G.(\mathsf{F}^{s},\dot{w}\mathsf{F}^{s}),\text{ for some }w\in W^{\textsc{b}},$ (3.1.5)
where $\dot{w}$ is corresponding permutation matrix in $G$. For example, $\dot{s}_{i}$ is the matrix obtained from the identity matrix $I_{2r}$ by swapping its rows $i$ and $i+1$, and rows $2r+1-i$ and $2r-i$. Hence, putting $\dot{W}:=\{\dot{w}\mid w\in W^{\textsc{b}}\}\subset\Xi_{2r,2r}$, and $({\mathcal{B}}_{r}\times{\mathcal{B}}_{r})/{{\mathrm{O}}_{2r}(q)}$ to be the set of ${{\mathrm{O}}_{2r}(q)}$-orbits in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$, we have a bijective map
${\mathfrak{p}}:(\mathcal{B}_{r}\times\mathcal{B}_{r})/{{\mathrm{O}}_{2r}(q)}\longrightarrow\dot{W}.$ (3.1.6)
In general, for $\lambda\in\Lambda(n+1,r)$, where $2n+1\leq 2r$, let $P_{\widehat{\lambda}}(q)$ be the standard parabolic subalgebra of ${\text{\rm GL}}_{2r}(q)$ associated with $\widehat{\lambda}$, consisting of upper quasi-triangular matrices with blocks of sizes $\widehat{\lambda}_{i}$ on the diagonal. Let
$P_{\lambda}(q)=P_{\widehat{\lambda}}(q)\cap G(q).$
Then $G(q)$ acts on the set $G(q)/P_{\lambda}(q)$ of left cosets $gP_{\lambda}(q)$ in $G(q)$.
Associated with $\lambda$, there is a pure isotropic $n$-step partial flag $F^{\lambda}$ such that
$F^{\lambda}_{1}=\langle e_{1},\ldots,e_{\widetilde{\lambda}_{1}}\rangle,F^{\lambda}_{2}=\langle e_{1},\ldots,e_{\widetilde{\lambda}_{2}}\rangle,\ldots,F^{\lambda}_{n}=\langle e_{1},\ldots,e_{\widetilde{\lambda}_{n}}\rangle.$ (3.1.7)
Then $P_{\lambda}(q)=\text{Stab}_{G(q)}(F^{\lambda})$ and there is a $G(q)$-set isomorphism $G(q)/P_{\lambda}(q)\cong G(q).F^{\lambda}$.
Lemma 3.2.
Maintain the notation above. We have $P_{\lambda}(q)\subseteq{\text{\rm SO}}_{2r}(q)$ if and only if $\lambda_{n+1}=0$.
3.4. Parametrizing ${\mathrm{O}}_{2r}(q)$-orbits
The set ${\mathcal{F}}_{n,r}^{\jmath}$ is a finite $G(q)$-set. If ${\mathcal{F}}_{n,r}^{\jmath}/G(q)$ denotes the set of $G(q)$-orbits in ${\mathcal{F}}_{n,r}^{\jmath}$, then there is a surjective map onto the set $\widehat{\Lambda}(n+1,r)$:
${\bf dim}:{\mathcal{F}}_{n,r}^{\jmath}\longrightarrow\widehat{\Lambda}(n+1,r),\;\;F\longmapsto{\bf dim}(F),\;\;\text{where}$ (3.2.1)
${\bf dim}(F):=($ $\text{\rm dim\,}F_{1}/F_{0},\text{\rm dim\,}F_{2}/F_{1},\ldots,\text{\rm dim\,}F_{n}/F_{n-1},\text{\rm dim\,}F_{n+1}/F_{n},\ldots,\text{\rm dim\,}F_{N}/F_{N-1})$
is called the dimension sequence (or dimension vector), induces a bijective map from ${\mathcal{F}}_{n,r}^{\jmath}/G(q)$ to $\widehat{\Lambda}(n+1,r)$. Thus, the $G$-orbits are indexed by the set $\Lambda(n+1,r)$.
For each pair $(F,F^{\prime})$ in ${\mathcal{F}}_{n,r}^{\jmath}\times{\mathcal{F}}_{n,r}^{\jmath}$, let $F_{i,j}=F_{i-1}+F_{i}\cap F^{\prime}_{j}$ and $F^{\prime}_{i,j}=F^{\prime}_{i-1}+F^{\prime}_{i}\cap F_{j}$. Then the filtration $F$ is refined to
[…]
$=F_{2,0})\subseteq F_{2,1}\subseteq F_{2,2}\subseteq\cdots\subseteq F_{2,N}(=F_{2}$
[…]
$=F_{N,0})\subseteq F_{N,1}\subseteq F_{N,2}\subseteq\cdots\subseteq F_{N,N}(=F_{N}).$
For simplicity, we write $F_{\!\centerdot\centerdot}$ for $(F,F^{\prime})_{\!\centerdot\centerdot}$ and $F^{\prime}_{\!\centerdot\centerdot}$ for $(F^{\prime},F)_{\!\centerdot\centerdot}$.
Lemma 3.3.
The subspace filtration (3.2.2) is an $n(N+1)$-step isotropic (partial) flag. In other words, $F_{\!\centerdot\centerdot}\in{\mathcal{F}}^{\jmath}_{n(N+1),r}$.
Proof.
If we set $F_{\!\centerdot\centerdot}=(V_{1}\subseteq V_{2}\subseteq\cdots\subseteq V_{N^{2}})$, then $F_{i,j}=V_{(i-1)N+j}$. We need to prove that $(F_{i,j})^{\perp}=V_{N^{2}-(i-1)N-j}=F_{N-i+1,N-j}$ since $N^{2}-(i-1)N-j=(N-i)N+N-j$. This can be checked easily.
We now prove that all $V_{i}$, for $1\leq i\leq n(N+1)$, are isotropic subspaces of $\mathbb{F}_{q}^{2r}$. Since $F_{i,j}\subseteq V_{i}$, for all $i\in[1,n]$, are clearly isotropic, it remains to prove that $F_{n+1,j}$, for $j\in[1,n]$ are isotropic. This is because $(F_{n}+(F_{n+1}\cap F^{\prime}_{j}))^{\perp}=F_{n}^{\perp}\cap(F_{n+1}^{\perp}+F_{j}^{\prime\perp})=F_{n+1}\cap(F_{n}+F^{\prime}_{N-j})=F_{n}+F_{n+1}\cap F^{\prime}_{N-j}\supseteq F_{n+1,j}.$
The dimension vector dim associated with $F_{\!\centerdot\centerdot}$ (see (3.2.1)) gives rise to a dimension matrix
${\mathfrak{m}}(F,F^{\prime}):={\bf dim}(F,F^{\prime})_{\!\centerdot\centerdot}=(a_{i,j}),\quad\text{where }a_{ij}=\text{\rm dim\,}(F_{i,j}/F_{i,j-1}).$ (3.3.1)
Note that dim ($\text{\rm dim\,}(F_{i,j}/F_{i,j-1})=\text{\rm dim\,}(F^{\prime}_{j,i}/F^{\prime}_{j,i-1})$, which follows from the vector space isomorphism
$F_{i,j}/F_{i,j-1}\cong\frac{F_{i}\cap F_{j}^{{}^{\prime}}}{F_{i-1}\cap F_{j}^{{}^{\prime}}+F_{i}\cap F_{j-1}^{{}^{\prime}}}\cong F^{\prime}_{j,i}/F^{\prime}_{j,i-1}$
by the Zassenhaus Lemma.
For $A\in\Xi_{N,2r}$ with ${\widehat{\lambda}}=\rm{ro}(A)$ and ${\widehat{\mu}}=\text{\rm co}(A)$, we partition $[1,2r]$ into $N^{2}$ subsets $I_{j,l}$ for $11\leq j,l\leq N=2n+$ 1, where
$I_{j,l}=I_{j,l}(A):=\big{[}\widetilde{a}^{r}_{j,l-1}+1,\widetilde{a}^{r}_{j,l-1}+a_{j,l}\big{]}=R^{\widehat{\lambda}}_{j}\cap d_{A}R^{\widehat{\mu}}_{l},$ (3.3.2)
where $\widetilde{a}^{r}_{j,l-1}$ is defined in (2.3.1) (and (2.1.6)).
Lemma 3.4.
(1) If $w\in W^{\textsc{b}}$ is the permutation of $[1,2r]$ sending $k$ to $i_{k}$ and suppose that the complete flags $(\mathsf{F},\mathsf{F}^{w})\in\mathcal{B}_{r}\times\mathcal{B}_{r}$ have the form $\mathsf{F}_{j}=\langle v_{1},\ldots,v_{j}\rangle$ and $\mathsf{F}^{w}_{j}=\langle v_{i_{1}},\ldots,v_{i_{j}}\rangle$, for $1\leq j\leq 2r$, then ${\mathfrak{m}}(\mathsf{F},\mathsf{F}^{w})=\dot{w}$.
[…]
Proof.
(1) By the transitivity of the action of ${\text{\rm O}}_{2r}(q)$ in ${\mathcal{B}}_{r}$, there exists a $g\in{\text{\rm O}}_{2r}(q)$ such that $(\mathsf{F},\mathsf{F}^{\prime})=g.(\mathsf{F}^{s},\dot{w}\mathsf{F}^{s})$, for some $w\in W$ (see (3.1.5)), where $\mathsf{F}^{\sf s}$ is the standard flag, i.e., $\mathsf{F}^{s}_{i}=\langle e_{1},\ldots,e_{i}\rangle$. We may simply take $(\mathsf{F},\mathsf{F}^{w})=(\mathsf{F}^{s},\dot{w}\mathsf{F}^{s})$ and assume $v_{i}=e_{i}$ for all $i\in[1,2r]$.
We first prove that $\mathsf{F}^{w}=(\mathsf{F}^{\sf s})^{w}$ is isotropic. For any $1\leq m\leq 2r,$ by the definition of $w\in W$ (see (2.1.4)), we have $i_{m}+i_{2r+1-m}=2r+1$ which, by (3.1.1), implies that $e_{i_{m}}\perp e_{i_{p}}$ for all $p\in[1,2r]\setminus\{2r-m+1\}$. Thus, $e_{i_{m}}\perp e_{i_{k}}$ for all $m\in[1,j]$ ($j\leq r$) and $k\in[1,2r-j]$ since $m+k<2r+1$. Hence, ${\mathsf{F}^{w}_{j}}^{\perp}\supseteq\mathsf{F}^{w}_{2r-j}$. Conversely, suppose there is a nonzero element $v=\sum_{m\in[1,2r]}c_{m}e_{i_{m}}\in{\mathsf{F}^{w}_{j}}^{\perp}\setminus\mathsf{F}^{w}_{2r-j}$. Then there must exist some $k\geq 2r+1-j$ such that $c_{k}\neq 0$. Thus, $e_{i_{2r+1-k}}\in\mathsf{F}^{w}_{j}$ and
$\langle v,e_{i_{2r+1-k}}\rangle=c_{k}\langle e_{i_{k}},e_{i_{2r+1-k}}\rangle\neq 0$
which contradicts to $v\in{\mathsf{F}^{w}_{j}}^{\perp}$. Hence, we must have ${\mathsf{F}^{w}_{j}}^{\perp}\subseteq\mathsf{F}^{w}_{2r-j}$, proving ${\mathsf{F}^{w}_{j}}^{\perp}=\mathsf{F}^{w}_{2r-j}$. So, $\mathsf{F}^{\prime}$ is an isotropic flag.
By the definition of ${\mathfrak{m}}(\mathsf{F},\mathsf{F}^{\prime})=A=(a_{k,l})$, we have $a_{k,l}=\text{\rm dim\,}\frac{\mathsf{F}_{k}\cap\mathsf{F}^{\prime}_{l}}{\mathsf{F}_{k-1}\cap\mathsf{F}^{\prime}_{l}+\mathsf{F}_{k}\cap\mathsf{F}^{\prime}_{l-1}}$. Thus,
$a_{k,l}$ $=\text{\rm dim\,}\frac{\langle v_{1},\cdots,v_{k}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l}}\rangle}{\langle v_{1},\cdots,v_{k-1}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l}}\rangle+\langle v_{1},\cdots,v_{k}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l-1}}\rangle}$
=$=\begin{cases}\text{\rm dim\,}\frac{\langle v_{1},\cdots,v_{k-1}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l}}\rangle+\langle v_{1},\cdots,v_{k}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l-1}}\rangle}{\langle v_{1},\cdots,v_{k-1}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l}}\rangle+\langle v_{1},\cdots,v_{k}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l-1}}\rangle},&\text{ if}k\neq i_{l},\\ \\ \text{\rm dim\,}\frac{\langle v_{1},\cdots,v_{k-1}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l-1}}\rangle+\langle v_{k}\rangle}{\langle v_{1},\cdots,v_{k-1}\rangle\cap\langle v_{i_{1}},\cdots,v_{i_{l-1}}\rangle},&\text{ if }k=i_{l},\end{cases}$
$=\delta_{k,i_{l}},$
which is the same as the matrix $\dot{w}$.
(2) Associated with $A\in\Xi_{N,2r}$, there is a permutation, a distinguished double coset representative, sending $k$ to $i_{k}$, where $(i_{1},i_{2},\cdots,i_{2r})$ is the permutation of $1,2,\ldots,2r$ obtained by concatenating $I_{1,1}$ with $I_{2,1}$, down column 1 and then down column 2, and so on. See Lemma 2.2 and [8, Exer.8.2(1)]. Then, by definition, $F^{A}_{i}=\langle v_{i_{1}},\cdots,v_{i_{\tilde{\mu}_{i}}}\rangle$, where $\mu=(\mu_{1},\ldots,\mu_{N})=\text{co}(A)$. By part (1), $F^{A}$ is isotropic as it can be obtained by dropping some steps from a corresponding complete isotropic flag.
Now, $F_{k}\cap F_{l}^{{}^{\prime}}$ is spanned by $\textsc{b}_{k,l}:=\{v_{k}\mid k\in\bigcup_{i\in[1,k],j\in[1,l]}I_{i,j}\}$. Thus, $F_{k-1}\cap F_{l}^{{}^{\prime}}+F_{k}\cap F_{l-1}^{{}^{\prime}}$ has basis $\textsc{b}_{k-1,l}\cup\textsc{b}_{k,l-1}$. Hence, $\textsc{b}_{k,l}\setminus(\textsc{b}_{k-1,l}\cup\textsc{b}_{k,l-1})=\{v_{k}\mid k\in I_{k,l}\}$. Consequently, $a_{k,l}=\text{\rm dim\,}\frac{F_{k}\cap F_{l}^{{}^{\prime}}}{F_{k-1}\cap F_{l}^{{}^{\prime}}+F_{k}\cap F_{l-1}^{{}^{\prime}}}$. The last assertion follows from Lemma 2.4. ∎
We observe from the proof that the $n(N+1)$-step isotropic flag $F_{\!\centerdot\centerdot}$ associated with $(F,F^{\prime})$ has the following filtration terms
$F_{i,j}=F_{i-1}+F_{i}\cap F^{\prime}_{j}=\langle v_{1},v_{2},\cdots,v_{\widetilde{a}^{r}_{i,j}}\rangle$ (3.4.1)
ordered on the lexicographic ordering of $[1,N]\times[1,N]$.
Example 3.5.
Let
$A=\begin{pmatrix}1&2&3\\ 3&0&{\bf 3}\\ {\bf 3}&{\bf 2}&{\bf 1}\end{pmatrix}.\quad$
Then $n=1$, $N=3$, $r=9$, ${\widehat{\lambda}}=\text{\rm ro}(A)=(6,6,6),{\widehat{\mu}}=\text{\rm co}(A)=(7,4,7)$, and $d_{A}=\big{(}\begin{smallmatrix}1&2&3&4&5&6&7&8&9&\cdots\\ 1&7&8&9&13&14&15&2&3&\cdots\end{smallmatrix}\big{)}.$
Breaking the natural basis $(e_{1},e_{2}\ldots,e_{18})$ into subsequences of lengths $a_{1,1},a_{1,2},a_{1,3},a_{2,1},\ldots,a_{3,3}$ and form the matrix of the subsequences
$A(e_{1},\ldots,e_{18})=\begin{pmatrix}e_{1}&(e_{2},e_{3})&(e_{4},e_{5},e_{6})\\ (e_{7},e_{8},e_{9})&-&(e_{10},e_{11},e_{12})\\ (e_{13},e_{14},e_{15})&(e_{16},e_{17})&e_{18}\\ \end{pmatrix}$
Then $F=F^{\widehat{\lambda}}$ with $F^{\widehat{\lambda}}_{i}$ being spanned by all rows $j$, $1\leq j\leq i$, and $F^{\prime}=d_{A}F^{\widehat{\mu}}$ with $F^{\prime}_{i}$ being spanned by all columns $j$, $1\leq j\leq i$:
$F_{1}$ $=\langle e_{1},\ldots,e_{6}\rangle,\;\;F_{2}=F_{1}+\langle e_{7},\ldots,e_{12}\rangle,\;\;F_{3}=F_{2}+\langle e_{13},\ldots,e_{18}\rangle,$
$F_{1}^{\prime}$ $=\langle e_{1},e_{7},e_{8},e_{9},e_{13},e_{14},e_{15}\rangle,\;\;F^{\prime}_{2}=F_{1}^{\prime}+\langle e_{2},e_{3},e_{16},e_{17}\rangle,\;\;F^{\prime}_{3}=F^{\prime}_{2}+\langle e_{4},e_{5},e_{6},e_{10},e_{11},e_{12},e_{18}\rangle,$
and $F_{\!\centerdot\centerdot}$ has the form
$F_{i,j}=F^{\widehat{\lambda}}_{i-1}+\text{subspaces spanned by subsequences oflengths $a_{i,1},\ldots,a_{i,j}$}=\langle e_{1},e_{2},\ldots,e_{\widetilde{a}^{r}_{i,j}}\rangle.$
For vector spaces $U,V$, $n$-step isotropic flags $F,F^{\prime}\in\mathcal{F}_{n,r}^{\jmath}$, and $h\in[1,n]$, we set
[…]
We set $E_{ij}^{\theta}=E_{ij}+E_{N+1-i,N+1-j},$ where $E_{ij}$ is the $N\times N$ matrix whose $(i,j)$-entry is $1$ and all other entries are 0.
Corollary 3.6.
Let $E,F,F^{\prime}\in\mathcal{F}_{n,r}^{\jmath}$ and $h\in[1,n]$. We have
[…]
Here $p$ satisfies the conditions $F_{h}\cap F_{j}^{\prime}=E_{h}\cap F_{j}^{\prime}$ for $j<p$ and $F_{h}\cap F_{j}^{\prime}\neq E_{h}\cap F_{j}^{{}^{\prime}}$ for $j\geq p$.
Proof.
The first assertion follows from Lemma 3.4(2) immediately.
[…]
Suppose $F^{\prime}=E^{A}$. Then $a_{kl}=\text{\rm dim\,}\frac{E_{k}\cap F_{l}^{\prime}}{E_{k-1}\cap F_{l}^{\prime}+E_{k}\cap F_{l-1}^{\prime}}$ for all $k,l$. We compute $a_{kl}^{\prime}=\text{\rm dim\,}\frac{F_{k}\cap F_{l}^{\prime}}{F_{k-1}\cap F_{l}^{\prime}+F_{k}\cap F_{l-1}^{\prime}}$. Clearly, we have
$a_{kl}=a^{\prime}_{kl}\;\;\text{ for all $l$ and $k\neq h,h+1,N-h,N+1-h$}.$ (3.6.1)
With the notation in the proof of Lemma 3.4, we have bases $\textsc{b}_{k,l}\subseteq{\mathcal{B}}$ for $E_{k}\cap F^{\prime}_{l}$ and $\textsc{b}^{\prime}_{k,l}\subseteq{\mathcal{B}}$ for $F_{k}\cap F^{\prime}_{l}$. Then
$a^{\prime}_{h,j}$ $=|\textsc{b}^{\prime}_{h,j}|-|\textsc{b}^{\prime}_{h,j-1}|-|\textsc{b}^{\prime}_{h-1,j}|+|\textsc{b}^{\prime}_{h-1,j-1}|$ (3.6.2)
$=|\textsc{b}_{h,j}|-|\textsc{b}_{h,j-1}|-|\textsc{b}_{h-1,j}|+|\textsc{b}_{h-1,j-1}|+|\textsc{b}^{\prime}_{h,j}|-|\textsc{b}^{\prime}_{h,j-1}|-|\textsc{b}_{h,j}|+|\textsc{b}_{h,j-1}|$
$=a_{h,j}+(|\textsc{b}^{\prime}_{h,j}|-|\textsc{b}_{h,j}|)-(|\textsc{b}^{\prime}_{h,j-1}|-|\textsc{b}_{h,j-1}|)\;\;\;\text{and similarly,}$,
$a^{\prime}_{h+1,j}$ $=a_{h+1,j}-(|\textsc{b}^{\prime}_{h,j}|-|\textsc{b}_{h,j}|)+(|\textsc{b}^{\prime}_{h,j-1}|-|\textsc{b}_{h,j-1}|).$
[…]
$|\textsc{b}_{h,j}|-|\textsc{b}^{\prime}_{h,j}|=\begin{cases}0,&\text{ if }j<p;\\ 1,&\text{ if }j\geq p.\end{cases}$ (3.6.3)
[…]
$I_{h,p}$ $=I^{\prime}_{h,p}\cup\{m\},\qquad I_{N+1-h,N+1-p}=I^{\prime}_{N+1-h,N+1-p}\cup\{{2r+1-m}\};$
$I_{h+1,p}$ $=I^{\prime}_{h+1,p}\setminus\{m\},\quad I_{N-h,N+1-p}=I^{\prime}_{N-h,N+1-p}\setminus\{{2r+1-m}\}\;\;(I_{j,l}=I^{\prime}_{j,l},\text{ for other}j,l),$
and set $E_{i}=\langle v_{j}\mid j\in\cup_{k=1}^{n}I_{i,k}\rangle$. Then $E_{h}=F_{h}+\langle v_{m}\rangle$ and $E_{i}=F_{i}$ for other $i$. One checks easily that $E$ is isotropic using the fact $\langle v_{m},v_{2r+1-m}\rangle_{\mathsf{J}}\neq 0$. Finally, by an argument around (3.6.1), (3.6.2) and (3.6.3), one checks easily that ${\mathfrak{m}}(E,F^{\prime})=A$. ∎
Let $(\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath})/G(q)$ denote the set of $G(q)$-orbits in $\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}$. Then, we obtain a map
${\mathfrak{m}}:(\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath})/{{\mathrm{O}}_{2r}(q)}\longrightarrow\Xi_{N,2r},\;\;G(q).(F,F^{\prime})\longmapsto{\mathfrak{m}}(F,F^{\prime}).$ (3.6.4)
The “equal parameter” counterpart of the following result can be found in [4, Lem. 2.1].
Corollary 3.7.
The map ${\mathfrak{m}}$ is bijective.
Proof.
The following proof modifies the proof of [4, Lem. 1.2]. First, the proof for a well-defined ${\mathfrak{m}}$ uses the argument there. The surjectivity follows from part (2) of the lemma above. Finally, the injectivity of ${\mathfrak{m}}$ follows from Lemma 3.1(2) (with $X={\mathcal{X}}$) and the bijection (2.2.1), noting the relation between $n$-step isotropic flags and parabolic subgroups introduced in the next subsection. ∎
For an orbit $(F,F^{\prime})\in\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}$, if ${\mathfrak{m}}(F,F^{\prime})=A$, we write ${\mathcal{O}}_{A}:=G.(F,F^{\prime})$. Similarly, for $(\mathsf{F},\mathsf{F}^{\prime})\in{\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$, if ${\mathfrak{m}}(F,F^{\prime})=\dot{w}$, we set ${\mathcal{O}}_{w}=G(q).(F,F^{\prime})$.
For any $A=(a_{i,j})\in\Xi_{N,2r}$, $h\in[1,n]$ and $p\in[1,N]$, define
${}_{h}\!A_{p}=A+E_{h,p}^{\theta}-E^{\theta}_{h+1,p}\;\;\text{ and }\;\;{}_{h}\!A_{\overline{p}}=A-E_{h,p}^{\theta}+E^{\theta}_{h+1,p}.$ (3.7.1)
Thus, ${}_{h}\!A_{p},{}_{h}\!A_{\overline{p}}\in\Xi_{N,2r}$ imply $a_{h+1,p}\geq 1$ and $a_{h,p}\geq 1$, respectively. For later use, we record the following result which follows from Corollary 3.6.
Proposition 3.8.
Let $h\in[1,n]$, $p\in[1,N]$, $A,B,C\in\Xi_{N,2r}$ with $B-E^{\theta}_{h,h+1}$ and $C-E^{\theta}_{h+1,h}$ diagonal. Let $(F,F^{\prime})\in{\mathcal{O}}_{A^{\prime}}$ with $A^{\prime}={{}_{h}\!A_{p}}$ (resp., ${}_{h}\!A_{\overline{p}}$) and $\alpha=\text{\rm ro}(A^{\prime})$. Suppose $\{v_{1},v_{2},\ldots,v_{2r}\}$ is a basis for $\mathbb{F}_{q}^{2r}$ such that $F_{i}=\langle v_{1},v_{2},\ldots,v_{\widetilde{\alpha}_{i}}\rangle$ for all $i\in[1,N]$ and $F^{\prime}=F^{A^{\prime}}$. Then the flag $E\in{\mathcal{F}}^{\jmath}_{n,r}$, obtained from $F$ by moving a vector $v_{m}$ from $\{v_{k}\mid k\in I^{\prime}_{h,p}\;(\text{resp., }I^{\prime}_{h+1,p})\}$ to $\{v_{k}\mid k\in I^{\prime}_{h+1,p}\;(\text{resp., }I^{\prime}_{h,p})\}$ and moving $v_{2r+1-m}$ from $\{v_{k}\mid k\in I^{\prime}_{N+1-h,N+1-p}\;(\text{resp., }I^{\prime}_{N-h,N+1-p})\}$ to $\{v_{k}\mid k\in I^{\prime}_{N-h,N+1-p}\;(\text{resp., }I^{\prime}_{N+1-h,N+1-p})\}$, satisfies ${\mathfrak{m}}(F,E)=B$ (resp., $C$) and ${\mathfrak{m}}(E,F^{\prime})=A$.
3.5. A geometric setting for $H^{\textsc{b}}_{{\boldsymbol{q}},1}$
Let $G(q)={\mathrm{O}}_{2r}(q)$ and let $\mathfrak{F}_{G(q)}(\mathcal{B}_{r}\times\mathcal{B}_{r})=\mathfrak{F}_{G(q)}(\mathcal{B}_{r}\times\mathcal{B}_{r},\mathbb{Z})$. Let $\tau_{w}$ be the orbital function associated with orbit ${\mathcal{O}}_{w}$; see (3.0.1). Then, for any $1\leq j\leq r$ and $\mathsf{F},\mathsf{F}^{\prime}\in\mathcal{B}_{r}$, $\tau_{j}=\tau_{s_{j}}\in\mathfrak{F}_{G(q)}(\mathcal{B}_{r}\times\mathcal{B}_{r})$ has the form
$\tau_{j}(\mathsf{F},\mathsf{F}^{{}^{\prime}})=\begin{cases}1,&\text{ if $\mathsf{F},\mathsf{F}^{\prime}$ differ only at $j$-th (and $(2r-j)$-th) subspace};\\ 0,&\text{ otherwise.}\end{cases}$
In other words, if two isotropic complete flags $\mathsf{F}$ and $\mathsf{F}^{\prime}$ differ only at the $j$-th subspace (i.e, $\mathsf{F}_{i}=\mathsf{F}^{\prime}_{i}$, for all $i\in[1,r]$, $i\neq j$, and $\mathsf{F}_{j}\neq\mathsf{F}^{\prime}_{j}$), then the orbit ${\mathcal{O}}_{j}=G(q).(\mathsf{F},\mathsf{F}^{\prime})$ has image ${\mathfrak{p}}(\mathsf{F},\mathsf{F}^{\prime})=\dot{s}_{j}$, for all $j\in[1,r]$ (i.e., ${\mathcal{O}}_{j}={\mathcal{O}}_{s_{j}}$). (Note that each $\mathsf{F}$ is an $r$-step isotropic filtration extended to a $2r$-step filtration.)
We also write ${\bf 1}=\tau_{w}$, for $w=1$. This is the function associated with the orbit ${\mathcal{O}}_{1}:=G(q).(\mathsf{F},\mathsf{F})$ for the identity $1\in W$.
Recall the Hecke algebra $H_{{\boldsymbol{q}},1}=H_{{\boldsymbol{q}},1}^{\textsc{b}}$ of type $B_{r}$ from §2. Let $H_{q,1}:=H_{{\boldsymbol{q}},1}|_{{\boldsymbol{q}}=q}:=H_{{\boldsymbol{q}},1}\otimes_{\mathcal{A}}\mathbb{Z}$ be the $\mathbb{Z}$-algebra obtained by specializing ${\boldsymbol{q}}$ to the prime power $q$. Then
$H_{q,1}\cong H_{{\boldsymbol{q}},1}/({\boldsymbol{q}}-q)H_{{\boldsymbol{q}},1}.$
Note that $H_{q,1}$ can be presented by (2.0.4) with ${\boldsymbol{q}}$ replaced by $q$. The following result can be proved directly, using convolution product, or via Lemma 3.1 (see [16]).
Proposition 3.9.
Let $G(q)={\mathrm{O}}_{2r}(q)$. There is a $\mathbb{Z}$-algebra isomorphism
$H_{q,1}\cong\mathfrak{F}_{G(q)}(\mathcal{B}_{r}\times\mathcal{B}_{r})$
defined by sending $T_{j}$ to $\tau_{j}$, for all $j\in[1,r]$.
Proof.
We only verify the degenerate relation $\tau_{r}^{2}=1$. Observe that, for $\mathsf{F},\mathsf{F}^{\prime}\in\mathcal{B}_{r}$,
$\tau_{r}^{2}(\mathsf{F},\mathsf{F}^{\prime})=\sum_{{\mathsf{F}^{\prime\prime}}\in\mathcal{B}_{r}}\tau_{r}(\mathsf{F},{\mathsf{F}^{\prime\prime}})\tau_{r}({\mathsf{F}^{\prime\prime}},\mathsf{F}^{\prime})\neq 0$
implies $\exists\mathsf{F}^{\prime\prime}$ s.t. $(\mathsf{F},\mathsf{F}^{\prime\prime})\in{\mathcal{O}}_{r}$ and $(\mathsf{F}^{\prime\prime},\mathsf{F}^{\prime})\in{\mathcal{O}}_{r}$. Thus, $\mathsf{F}_{r}\neq\mathsf{F}^{\prime\prime}_{r}\neq\mathsf{F}^{\prime}_{r}$ are all maximal isotropic. Suppose $\mathsf{F}_{r}=\mathsf{F}_{r-1}+\langle u\rangle$, $\mathsf{F}^{\prime\prime}_{r}=\mathsf{F}_{r-1}+\langle v\rangle$ and $\mathsf{F}^{\prime}_{r}=\mathsf{F}_{r-1}+\langle\alpha u+\beta v\rangle$. Then $\langle u,u\rangle_{\mathsf{J}}=0=\langle v,v\rangle_{\mathsf{J}}$ and $\langle u,v\rangle_{\mathsf{J}}=1$. So $0=\langle\alpha u+\beta v,\alpha u+\beta v\rangle=2\alpha\beta$, forcing $\beta=0$. Hence, $\mathsf{F}=\mathsf{F}^{\prime}$ and in this case $\tau_{r}^{2}(F,F)=1$. Consequently, $\tau_{r}^{2}={\bf 1}.$
3.6. A geometric setting for $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$
Let $S^{\textsc{b}}_{q,1}(n,r):=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)|_{{\boldsymbol{q}}=q}=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)\otimes_{\mathcal{A}}\mathbb{Z}$ be the $\mathbb{Z}$-algebra obtained by specializing ${\boldsymbol{q}}$ to the prime power $q$. Then
$S^{\textsc{b}}_{q,1}(n,r):=S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)|_{{\boldsymbol{q}}=q}\cong S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)/({\boldsymbol{q}}-q)S_{{\boldsymbol{q}},1}(n,r).$
Let $G(q)={\text{\rm O}}_{2r}(q)$, ${\mathcal{X}}=\mathcal{F}_{n,r}^{\jmath}$ and $\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}})=\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}},\mathbb{Z})$. Recall the natural basis $\{e_{A}|_{{\boldsymbol{q}}=q}\mid A\in\Xi_{N,2r}\}$ for $S^{\textsc{b}}_{q,1}(n,r)$ and the orbital basis $\{f_{\mathcal{O}}\}$ for $\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}})$
Theorem 3.10.
There is a $\mathbb{Z}$-algebra isomorphism
$S^{\textsc{b}}_{q,1}(n,r)\cong\mathfrak{F}_{G(q)}({\mathcal{X}}\times{\mathcal{X}}),$
sending $e_{A}|_{{\boldsymbol{q}}=q}$ to $f_{{\mathcal{O}}_{A}}$ for all $A\in\Xi_{N,2r}$.
The proof is standard by Lemma 3.1(2) and an argument similar to that of [8, Th. 13.15].
4. A natural basis and geometric setting for $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$
We describe the natural (or double coset, or orbital) basis for the ${\boldsymbol{q}}$-Schur algebra of type $D$ according to its algebraic or geometric setting.
4.1. The algebraic definition of the natural basis
Recall the Hecke algebra $H=H_{{\boldsymbol{q}},1}$ of type $B_{r}$ and its subalgebra $\check{H}=\check{H}_{\boldsymbol{q}}$ of type $D_{r}$. By restriction, every $H$-module $M$ is an $\check{H}$-module, denoted by $M|_{\check{H}}$. For any $\alpha\in\Lambda^{\textsc{d}}(n,r)$, let ${\check{x}}_{\alpha}=\sum_{w\in{\check{W}}_{\alpha}}T_{w}$.
Lemma 4.1.
For any $\lambda\in\Lambda(n+1,r)$, there are $\check{H}$-module isomorphisms:
$x_{\lambda}H|_{\check{H}}\cong\begin{cases}{\check{x}}_{\lambda^{+}}\check{H}\oplus{\check{x}}_{\lambda^{-}}\check{H},&\text{ if }\lambda_{n+1}=0;\\ {\check{x}}_{\lambda^{\bullet}}\check{H},&\text{ if }\lambda_{n+1}\neq 0.\end{cases}$
Proof.
Since $W=\check{W}\cup s_{r}\check{W}$ is a disjoint union, we have $H=\check{H}\oplus T_{r}\check{H}$. Thus,
$x_{\lambda}H=\begin{cases}x_{\lambda}\check{H}\oplus x_{\lambda}T_{r}\check{H},&\text{ if }\lambda_{n+1}=0;\\ x_{\lambda}\check{H},&\text{ if }\lambda_{n+1}\neq 0.\\ \end{cases}$
If $\lambda_{n+1}=0$, then $x_{\lambda}={\check{x}}_{\lambda^{+}}$ and $T_{r}x_{\lambda}T_{r}={\check{x}}_{\lambda^{-}}$ by Corollary 2.7(1)&(2). We have $x_{\lambda}\check{H}={\check{x}}_{\lambda^{+}}\check{H}$ and $x_{\lambda}T_{r}\check{H}=T_{r}{\check{x}}_{\lambda^{-}}\check{H}\cong{\check{x}}_{\lambda^{-}}\check{H}$. If $\lambda_{n+1}\neq 0$, by Corollary 2.7(3), we have either $x_{\lambda}={\check{x}}_{\lambda^{\bullet}}$ or $x_{\lambda}=(1+T_{r}){\check{x}}_{\lambda^{\bullet}}$. In both cases, we have $x_{\lambda}H|_{\check{H}}\cong{\check{x}}_{\lambda^{\bullet}}\check{H}$. ∎
Proposition 4.2.
There is an ${\mathcal{A}}$-algebra isomorphism $S_{\boldsymbol{q}}^{\kappa}(n,r)\cong S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. By base change, it extends to a ${\mathcal{Z}}$-algebra isomorphism ${\mathcal{S}}_{\upsilon}^{\kappa}(n,r)\cong{\mathcal{S}}_{\upsilon}^{\textsc{d}}(n,r).$
Proof.
By the above lemma, there is an $\check{H}$-module isomorphism
$g:\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H|_{\check{H}}\longrightarrow\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}\check{H}.$
This $\check{H}$-module isomorphism induces the required ${\mathcal{A}}$-algebra isomorphism
$\widetilde{g}:S_{\boldsymbol{q}}^{\textsc{d}}(n,r)\longrightarrow S_{\boldsymbol{q}}^{\kappa}(n,r),\;\;\phi\longmapsto g^{-1}\phi g.$
The proposition is proved. ∎
Similar to Corollary 2.5, the above isomorphism allows us to define the natural basis for $S_{\boldsymbol{q}}^{\kappa}(n,r)$.
For $\alpha,\beta\in\Lambda^{\textsc{d}}(n,r)$, let $\check{\mathcal{D}}_{\alpha,\beta}$ denote the set of distinguished ${\check{W}}_{\alpha}$-${\check{W}}_{\beta}$ double coset representatives. Let
$\check{\mathcal{D}}(n,r)=\{(\alpha,d,\beta)\mid\alpha,\beta\in\Lambda^{\textsc{d}}(n,r),d\in\check{\mathcal{D}}_{\alpha,\beta}\}.$ (4.2.1)
For each $(\alpha,d,\beta)\in\Lambda^{\textsc{d}}(n,r)$, define $\phi_{\alpha,\beta}^{d}\in S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ by setting
$\phi_{\alpha,\beta}^{d}({\check{x}}_{\gamma}h)=\delta_{\beta,\gamma}\sum_{w\in{\check{W}}_{\alpha}d{\check{W}}_{\beta}}T_{w}h=\delta_{\beta,\gamma}T_{{\check{W}}_{\alpha}d{\check{W}}_{\beta}}h,\text{ for all $h\in\check{H}$.}$
Corollary 4.3.
The set $\{\phi_{\alpha,\beta}^{d}\mid(\alpha,d,\beta)\in\check{\mathcal{D}}(n,r)\}$ forms a basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. We call this basis the natural basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. Moreover, the idempotents $1_{\alpha}$ considered in Lemma 2.11 take the form $1_{\alpha}=\phi_{\alpha,\alpha}^{1}$.
4.2. Parametrising ${\text{\rm SO}}_{2r}(q)$-orbits
We now construct the natural basis for $S_{\boldsymbol{q}}^{\kappa}(n,r)$ defined in (2.1.1) via the isomorphism in Proposition 3.10.
Let ${\check{G}}={\check{G}}(q)={\text{\rm SO}}_{2r}(q)$ (and $G=G(q)={\text{\rm O}}_{2r}(q)$). We start with parametrizing ${\check{G}}$-orbits in the complete flag variety ${\mathcal{B}}_{r}$ introduced in Section 3. Let $\mathsf{F}=\mathsf{F}^{s}$ be the standard complete flag with $\mathsf{F}_{i}=\langle e_{1},\ldots,e_{i}\rangle$ and let
$\mathcal{B}_{r}^{+}:={\check{G}}(q).\mathsf{F}^{s},\qquad\mathcal{B}_{r}^{-}:={\check{G}}(q)\dot{s}_{r}.\mathsf{F}^{s}.$
If $M_{r}=\langle e_{1},e_{2},\ldots,e_{r}\rangle$ is the standard maximal isotropic subspace in $\mathbb{F}_{q}^{2r}$ (of dimension $r$), then
$\mathcal{B}_{r}^{+}=\{\mathsf{F}\in\mathcal{B}_{r}\mid\text{\rm dim\,}(\mathsf{F}_{r}\cap M_{r})\equiv r\text{ mod }2\},\quad\mathcal{B}_{r}^{-}=\{\mathsf{F}\in\mathcal{B}_{r}\mid\text{\rm dim\,}(\mathsf{F}_{r}\cap M_{r})\equiv r-1\text{ mod }2\}.$ (4.3.1)
We have $\mathcal{B}_{r}=\mathcal{B}^{+}_{r}\sqcup\mathcal{B}_{r}^{-}$ and $\mathcal{B}_{r}^{+}\cong{\check{G}}(q)/B(q),\;\mathcal{B}_{r}^{-}\cong{\check{G}}(q)/B(q)^{\dot{s}_{r}}$, where $B(q)=\text{Stab}_{{\check{G}}(q)}(\mathsf{F}^{s})$ and $B(q)^{\dot{s}_{r}}=\text{Stab}_{{\check{G}}(q)}(\dot{s}_{r}\mathsf{F}^{s})$.
To parametrize ${\check{G}}(q)$-orbits in the $n$-step isotropic (partial) flag variety ${\mathcal{X}}:=\mathcal{F}_{n,2r}^{\jmath}$, we use the $G(q)$-orbits. Let ${\mathcal{X}}/G(q)$ denote the set of all $G(q)$-orbits. Then (3.2.1) implies a bijection
$f:{\mathcal{X}}/G(q)\longrightarrow\Lambda(n+1,r),\;\;G(q).F\longmapsto\big{(}\text{\rm dim\,}(F_{1}/F_{0}),\ldots,\text{\rm dim\,}(F_{n}/F_{n-1}),\text{\rm dim\,}(F_{n+1}/F_{n})/2\big{)}.$
Let ${\mathcal{O}}_{\lambda}=f^{-1}(\lambda)$. Then ${\mathcal{X}}$ is a disjoint union of orbits ${\mathcal{X}}=\sqcup_{\lambda\in\Lambda(n+1,r)}{\mathcal{O}}_{\lambda}$.
Lemma 4.4.
An orbit ${\mathcal{O}}_{\lambda}=G(q).F$ splits into two ${\check{G}}(q)$-orbits if and only if $\lambda_{n+1}=0$.
Proof.
Since $G(q)={\check{G}}(q)\sqcup{\check{G}}(q)\dot{s}_{r}$, it follows that $G(q).F={\check{G}}(q).F\cup{\check{G}}(q)\dot{s}_{r}.F$. Thus, by Lemma 3.2,
$\lambda_{n+1}>0\iff\dot{s}_{r}\in\text{Stab}_{G(q)}(F)\iff G(q).F={\check{G}}(q).F.$
Hence, $G(q).F={\check{G}}(q).F\sqcup{\check{G}}(q)\dot{s}_{r}.F\iff\lambda_{n+1}=0$. The lemma follows. ∎
Let
$\check{\mathcal{O}}_{\lambda^{\bullet}}={\mathcal{O}}_{\lambda},\;\;\check{\mathcal{O}}_{\lambda^{+}}={\check{G}}(q).\mathsf{F}^{s},\;\;\text{ and }\;\;\check{\mathcal{O}}_{\lambda^{-}}={\check{G}}(q)\dot{s}_{r}.\mathsf{F}^{s}.$
By the Lemma, we may decompose ${\mathcal{X}}$ into a disjoint union of ${\check{G}}(q)$-orbits:
${\mathcal{X}}=\Big{(}\bigsqcup_{\lambda\in\Lambda^{\bullet}(n+1,r)}\check{\mathcal{O}}_{\lambda^{\bullet}}\Big{)}\sqcup\Big{(}\bigsqcup_{\lambda\in\Lambda^{\circ}(n+1,r)}(\check{\mathcal{O}}_{\lambda^{+}}\sqcup\check{\mathcal{O}}_{\lambda^{-}})\Big{)}.$ (4.4.1)
We now describe the decomposition of ${\mathcal{X}}\times{\mathcal{X}}$ into ${\check{G}}(q)$-orbits.
If $G(q).(F,F^{{}^{\prime}})$ is a $G(q)$-orbit in ${\mathcal{X}}\times{\mathcal{X}}$, then it can be written as $G(q).(F,F^{{}^{\prime}})={\check{G}}(q).(F,F^{{}^{\prime}})\cup{\check{G}}(q)g.(F,F^{{}^{\prime}})$ for some $g\in G(q)\setminus{\check{G}}(q)$. Since the map ${\mathfrak{m}}$ is a bijection (Corollary 5.7), we shall denote a $G(q)$-orbit by $\mathcal{O}_{A}:=G(q).(F,F^{{}^{\prime}})$ if ${\mathfrak{m}}(F,F^{\prime})=A$, for some $A\in\Xi_{N,2r}$.
Proposition 4.5.
Let $A=(a_{ij})\in\Xi_{N,2r}$. The $G(q)$-orbit $\mathcal{O}_{A}$ splits into two ${\check{G}}(q)$-orbits if and only if $a_{n+1,n+1}=0$.
Proof.
Recall Lemma 3.3 and consider the map
$\Phi:\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}\longrightarrow\mathcal{F}_{n(N+1),r}^{\jmath},\ ({F},F^{{}^{\prime}})\longmapsto F_{\!\centerdot\centerdot}=(F,F^{\prime})_{\!\centerdot\centerdot}=(F_{i,j})_{1\leq i,j\leq N},$ (4.5.1)
where $F_{i,j}=F_{i-1}+F_{i}\cap F^{{}^{\prime}}_{j}$ is defined as in (3.2.2). This map is $G(q)$-equivariant since $\Phi(g(F,F^{{}^{\prime}}))=g\Phi(F,F^{{}^{\prime}})$ for any $g\in G(q).$ Hence, $\text{Stab}_{G(q)}(F,F^{\prime})$ is a subgroup of $\text{Stab}_{G(q)}(\{F_{i,j}\}_{i,j}).$ Thus, if we choose $(F,F^{\prime})\in{\mathcal{O}}_{A}$ such that $F=F^{\lambda}$ and $F^{\prime}=\dot{d}F^{\mu}$, where $\lambda=\text{\rm co}(A)$, $\mu=\text{\rm co}(A)$ and $d\in{\mathcal{D}}_{\lambda,\mu}$, then, with the notation in §3.4, $\text{Stab}_{G(q)}(F^{\lambda},\dot{w}F^{\mu})=P_{\lambda}\cap{}^{\dot{w}}P_{\mu}$ and $\text{Stab}_{G(q)}(\{F_{i,j}\}_{i,j})=P_{\nu}$, where $\nu$ is the composition defined by $W_{\nu}=W_{\lambda}\cap{}^{d}W_{\mu}$. By [6, Prop. 2.8.4], $U_{\lambda}(P_{\lambda}\cap{}^{\dot{w}}P_{\mu})=P_{\nu}$, where $U_{\lambda}$ is a unipotent subgroup of $P_{\lambda}$ in the Levi decomposition of $P_{\lambda}=L_{\lambda}U_{\lambda}$.
If $a_{n+1,n+1}\neq 0,$ then $a_{n+1,n+1}$ is positive even and $a_{n+1,n+1}=|I_{n+1,n+1}|\geq 2$ by the notation in Lemma 3.4(2). Moreover, $r,r+1\in I_{n+1,n+1}$. Hence, $\dot{s}_{r}\in\text{Stab}_{G(q)}(\{F_{i,j}\}_{i,j})$ is not contained in ${\check{G}}(q)$. In other words, there exists $\dot{s}_{r}\in G(q)\setminus{\check{G}}(q)$ such that $\dot{s}_{r}\in\text{Stab}_{G(q)}(F,F^{\prime})$. Hence, $\dot{s}_{r}.(F,F^{{}^{\prime}})=(F,F^{\prime})$ and so ${\mathcal{O}}_{A}={\check{G}}(q).(F,F^{{}^{\prime}})$.
If $a_{n+1,n+1}=0,$ then $\text{dim}({F}_{n+1,n+1}/{F}_{n+1,n})=a_{n+1,n+1}=0$. Thus,
$\text{Stab}_{G(q)}(F,F^{\prime})\leq\text{Stab}_{G(q)}(\{F_{i,j}\}_{i,j})\subseteq{\check{G}}(q).$
Hence, ${\mathcal{O}}_{A}$ splits into two ${\check{G}}(q)$-orbits
${\check{G}}(q).(F,F^{\prime})\;\;\text{ and }\;\;{\check{G}}(q)g.(F,F^{\prime}),$
where $g\in G(q)\setminus{\check{G}}(q)$. This finishes the proof. ∎
Recall the standard maximal isotropic subspace $M_{r}$ used in (4.3.1). By the proof above and the map $\Phi$ in (4.5.1), we define, for the orbit ${\mathcal{O}}_{A}$ containing $(F,F^{\prime})$,
$\mathcal{F}_{n(N+1),r}^{\jmath}(A)$ $=\Phi(\mathcal{O}_{A})=G(q).F_{\!\centerdot\centerdot}$ (4.5.2)
$d_{F,F^{\prime}}=\text{\rm dim\,}$ $\big{(}(F_{n}+F_{n+1}\cap F^{\prime}_{n+1})\cap M_{r}\big{)}.$
Restricting the map $\Phi$ to $\mathcal{O}_{A}$ yields a surjective $G(q)$-equivariant map $\Phi_{A}:\mathcal{O}_{A}\to\mathcal{F}_{n(N+1),r}^{\jmath}(A).$
Definition 4.6.
Suppose $A=(a_{i,j})\in\Xi_{N,2r}$ and ${\mathcal{O}}_{A}=G(q).(F,F^{\prime})$ for some $F,F^{\prime}\in{\mathcal{F}}_{n,r}^{\jmath}$ and $F_{\!\centerdot\centerdot}$ as in (3.2.2).
  • (1)
    If $a_{n+1,n+1}\neq 0$, define $\dot{\mathcal{O}}_{A}:={\mathcal{O}}_{A}$ to be the non-split $G(q)$-orbit.
  • (2)
    If $a_{n+1,n+1}=0$, then we label the two ${\check{G}}(q)$-orbits as follows. Since $F_{n+1,n}=F_{n+1,n+1}$ is a maximal isotropic subspace, the $G(q)$-orbit $\mathcal{F}_{n(N+1),r}^{\jmath}(A)=G(q).F_{\!\centerdot\centerdot}$ splits into two ${\check{G}}(q)$-orbits
    $\mathcal{F}_{n(N+1),r}^{\jmath}(A)=\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{+}\sqcup\ \mathcal{F}_{n(N+1),r}^{\jmath}(A)^{-}$
    where $\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{\pm}=\{\Phi(F,F^{\prime})\in\mathcal{F}_{n(N+1),r}^{\jmath}(A)\mid d_{F,F^{\prime}}\equiv r^{\pm}\text{ mod }2\}$, for $\begin{cases}r^{+}=r;\\ r^{-}=r-1.\end{cases}$ Define
    ${}^{+}\mathcal{O}_{A}:=\Phi_{A}^{-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{+}\right)\text{ and }{}^{-}\mathcal{O}_{A}:=\Phi_{A}^{-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A)^{-}\right).$
    This decomposition $\mathcal{O}_{A}={{}^{+}}\mathcal{O}_{A}\sqcup{{}^{-}}\mathcal{O}_{A}$ is further refined to
    $\mathcal{O}_{A}=\begin{cases}{}^{+}\mathcal{O}_{A}^{+}\sqcup{}^{-}\mathcal{O}_{A}^{-},&\text{ if }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=+,\\ {}^{+}\mathcal{O}_{A}^{-}\sqcup{}^{-}\mathcal{O}_{A}^{+},&\text{ if }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=-,\end{cases}$ (4.6.1)
    where $\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=\begin{cases}+,&\text{ if $|A^{\text{\normalsize$\llcorner$}}|$ is even;}\\ -,&\text{ otherwise.}\end{cases}$
We remark that there is an alternative interpretation for the above double signed labelling, using a similarly defined map for the transpose $A^{t}$ of $A$:
$\Phi^{\prime}_{A}:\mathcal{O}_{A}\longrightarrow\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t}),\ ({F},F^{{}^{\prime}})\longmapsto F^{\prime}_{\!\centerdot\centerdot}=\big{(}F^{\prime}_{i,j}:=F^{{}^{\prime}}_{i-1}+F^{{}^{\prime}}_{i}\cap F_{j}\big{)}_{1\leq i,j\leq N}.$
Thus, ${\mathcal{F}}_{n(N+1),r}^{\jmath}(A^{t})=G(q).F^{\prime}_{\!\centerdot\centerdot}$ splits into two orbits
$\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})$ $=\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{+}\sqcup\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{-},$
where $\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{\pm}=\{\Phi(F,F^{\prime})\in\mathcal{F}_{n(N+1),r}^{\jmath}(A)\mid d_{F^{\prime},F}\equiv r^{\pm}\text{ mod }2\}$. Thus, $\mathcal{O}_{A}=\mathcal{O}_{A}^{+}\sqcup\mathcal{O}_{A}^{-},$ where
$\mathcal{O}_{A}^{+}:=\Phi_{A}^{\prime-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{+}\right)\text{ and }\mathcal{O}_{A}^{-}:=\Phi_{A}^{\prime-1}\left(\mathcal{F}_{n(N+1),r}^{\jmath}(A^{t})^{-}\right).$
Lemma 4.7.
Maintain the notations given above for $A$ with $a_{n+1,n+1}=0$. We have
$\begin{cases}\mathcal{O}_{A}^{+}={}^{+}\mathcal{O}_{A},\ \mathcal{O}_{A}^{-}={}^{-}\mathcal{O}_{A},&\text{ if }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=+;\\ \mathcal{O}_{A}^{+}={}^{-}\mathcal{O}_{A},\ \mathcal{O}_{A}^{-}={}^{+}\mathcal{O}_{A},&\text{ if }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=-,\end{cases}$
and hence, ${}^{\epsilon_{1}}\mathcal{O}_{A}^{\epsilon_{2}}:=^{\epsilon_{1}}\mathcal{O}_{A}\cap\mathcal{O}_{A}^{\epsilon_{2}}$, for the four selections of $\epsilon_{1},\epsilon_{2}\in\{+,-\}$. Moreover, the $\pm$ signs ${\epsilon}_{1}$ and ${\epsilon}_{2}$ are determined by ${\epsilon}_{1}1=(-1)^{d_{F,F^{\prime}}-r}$ and ${\epsilon}_{2}1=(-1)^{d_{F^{\prime},F}-r}$.
Proof.
Let ${\widehat{\lambda}}=\text{\rm ro}(A)$ and ${\widehat{\mu}}=\text{\rm co}(A)$. We may assume $F=F^{\widehat{\lambda}}$ and $F^{\prime}=\dot{d}_{A}F^{\widehat{\mu}}$ and define $F_{i,j}$ (resp., $F^{\prime}_{i,j}$) with respect to $(F,F^{\prime})$ (resp., $(F^{\prime},F)$). Then, by (3.4.1), $F_{n+1,n+1}=M_{r}$, the standard maximal isotropic subspace. On the other hand,
$M_{F^{\prime},F}$ $:=F^{\prime}_{n+1,n+1}\cap M_{r}=(F^{\prime}_{n}+F^{\prime}_{n+1}\cap F_{n+1})\cap M_{r}$ (4.7.1)
$=\dot{d}_{A}\big{(}(F^{\prime}_{n}+F^{\prime}_{n+1}\cap\dot{d}_{A}^{-1}F_{n+1})\cap\dot{d}_{A}^{-1}M_{r}\big{)}=\dot{d}_{A}\big{(}F^{\prime}_{n+1,n+1}\cap\dot{d}_{A}^{-1}M_{r}\big{)}$
has dimension congruent to $r$ (mod 2) if $\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=+$, since $F^{\prime}_{n+1,n+1}=M_{r}$. This proves $\mathcal{O}_{A}^{+}={}^{+}\mathcal{O}_{A}$. The other cases can be proved similarly. ∎
For the convenience of type $D$ theory, we introduce the following “signed matrices”. Let
$\overset{\bullet}{\Xi}:=\{A\in\Xi_{N,2r}\mid a_{n+1,n+1}\neq 0\},\;\;\quad\overset{\circ}{\Xi}:=\{A\in\Xi_{N,2r}\mid a_{n+1,n+1}=0\}$
which partition $\Xi=\Xi_{N,2r}$ into a disjoint union $\Xi=\overset{\bullet}{\Xi}\sqcup\overset{\circ}{\Xi}$.
Definition 4.8.
(1) For $A\in\overset{\circ}{\Xi}$ and $\epsilon_{1},\epsilon_{2}\in\{+,-\}$, we use ${}^{\epsilon_{1}}A^{\epsilon_{2}}$ to label the ${\check{G}}(q)$-orbit ${}^{\epsilon_{1}}\mathcal{O}_{A}^{\epsilon_{2}}$ and write
$\check{\mathcal{O}}(^{\epsilon_{1}}A^{\epsilon_{2}}):=^{\epsilon_{1}}\!\!\mathcal{O}_{A}^{\epsilon_{2}};$ (4.8.1)
(2) For $A\in\overset{\bullet}{\Xi}$, we use $\dot{A}$ to label the ${\check{G}}$-orbit ${\dot{\mathcal{O}}}_{A}={\mathcal{O}}_{A}$ and write
$\check{\mathcal{O}}(\dot{A}):=\dot{{\mathcal{O}}}_{A}.$ (4.8.2)
Hence, as an ${\check{G}}(q)$-set, ${\mathcal{X}}\times{\mathcal{X}}$ is a disjoint union of orbits:
${\mathcal{X}}\times{\mathcal{X}}$ $=\Big{(}\bigsqcup_{A\in\overset{\bullet}{\Xi}}\dot{\mathcal{O}}_{A}\Big{)}\sqcup\Big{(}\bigsqcup_{A\in\overset{\circ}{\Xi}_{+}}({}^{+}\mathcal{O}_{A}^{+}\sqcup{}^{-}\mathcal{O}_{A}^{-})\Big{)}\sqcup\Big{(}\bigsqcup_{A\in\overset{\circ}{\Xi}_{-}}({}^{+}\mathcal{O}_{A}^{-}\sqcup{}^{-}\mathcal{O}_{A}^{+})\Big{)}$ (4.8.3)
$=\Big{(}\bigsqcup_{A\in\overset{\bullet}{\Xi}}\check{\mathcal{O}}(\dot{A})\Big{)}\sqcup\Big{(}\bigsqcup_{A\in\overset{\circ}{\Xi},{\epsilon},{\epsilon}^{\prime}\in\{+,-\}}\check{\mathcal{O}}({}^{\epsilon}\!\!A^{{\epsilon}^{\prime}})\Big{)}.$
We further introduce the following notation. Recall the row/column sum vectors in (2.3.2).
Definition 4.9.
Let $\Xi:=\Xi_{N,2r}$ and, for any $A=(a_{i,j})\in\Xi$, let ${\widehat{\lambda}}=\text{\rm ro}(A)$ and ${\widehat{\mu}}=\text{\rm co}(A)$.
(1) We further partition $\overset{\circ}{\Xi}$ into a disjoint union $\overset{\circ}{\Xi}=\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$, where
$\overset{\circ\!\circ\!\circ}{\Xi}$ =$=\{A\in\Xi_{N,2r}|\lambda_{n+1}=0,a_{n+1,n+1}=0\text{ and }\mu_{n+1}=0\},$
$\overset{\circ\!\circ\!\bullet}{\Xi}$ =$=\{A\in\Xi_{N,2r}|\lambda_{n+1}=0,a_{n+1,n+1}=0\text{ and }\mu_{n+1}\neq 0\},$
$\overset{\bullet\!\circ\!\circ}{\Xi}$ =$=\{A\in\Xi_{N,2r}|\lambda_{n+1}\neq 0,a_{n+1,n+1}=0\text{ and }\mu_{n+1}=0\},$
$\overset{\bullet\!\circ\!\bullet}{\Xi}$ =$=\{A\in\Xi_{N,2r}|\lambda_{n+1}\neq 0,a_{n+1,n+1}=0\text{ and }\mu_{n+1}\neq 0\}.$
Thus, we have $\Xi_{N,2r}=\overset{\bullet}{\Xi}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$.
[…]
$\overset{\star\circ*}{\Xi}_{\scriptscriptstyle{+}}$ :=$:=\{{{}^{+}{A}^{+}},{{}^{-}{A}^{-}}\mid A\in\overset{\star\circ*}{\Xi}\text{ and }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=+\},\quad$ (4.9.1)
$\overset{\star\circ*}{\Xi}_{\scriptscriptstyle{-}}$ :=$:=\{{{}^{+}{A}^{-}},{{}^{-}{A}^{+}}\mid A\in\overset{\star\circ*}{\Xi}\text{ and }\text{\rm sgn}(A^{\text{\normalsize$\llcorner$}})=-\},$
and
${\overset{\bullet}{\Xi}\centerdot}:=\{\dot{A}\mid A\in\overset{\bullet}{\Xi}\},\quad\overset{\circ}{\Xi}_{+}:=\overset{\circ\!\circ\!\circ}{\Xi}_{\!+}{\sqcup}\overset{\circ\!\circ\!\bullet}{\Xi}_{\!+}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}_{\!+}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}_{\!+},\quad\overset{\circ}{\Xi}_{-}:=\overset{\circ\!\circ\!\circ}{\Xi}_{\!-}{\sqcup}\overset{\circ\!\circ\!\bullet}{\Xi}_{\!-}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}_{\!-}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}_{\!-}.$
(Compare [19, (A.4.2)]. Define
$\check{\Xi}(n,r):={\overset{\bullet}{\Xi}\centerdot}\mathop{\sqcup}(\overset{\circ\!\circ\!\circ}{\Xi}_{\!+}{\sqcup}\overset{\circ\!\circ\!\circ}{\Xi}_{\!-})\sqcup(\overset{\circ\!\circ\!\bullet}{\Xi}_{\!+}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}_{\!-})\sqcup(\overset{\bullet\!\circ\!\circ}{\Xi}_{\!+}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}_{\!-})\sqcup(\overset{\bullet\!\circ\!\bullet}{\Xi}_{\!+}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}_{\!-})={\overset{\bullet}{\Xi}\centerdot}\sqcup\overset{\circ}{\Xi}_{+}\sqcup\overset{\circ}{\Xi}_{-}.$ (4.9.2)
Remark 4.10.
Consider the diagonal embedding ${\mathcal{X}}\to{\mathcal{X}}\times{\mathcal{X}},F\mapsto(F,F)$. This map is ${\check{G}}(q)$-equivariant as well as $G(q)$-equivariant. Thus, the parametrization in (4.4.1) is compatible with that in Definition 4.6. In other words, it induces injective maps
$\Lambda(n,r)\longrightarrow\Xi_{N,2r},\lambda\longmapsto D_{\lambda}:=\text{diag}(\lambda)\;\;\text{ and }\;\;\Lambda^{\textsc{d}}(n,r)\longrightarrow\check{\Xi}(n,r).$
The latter sends, for ${\epsilon}\in\{+,-\}$, $\lambda^{\epsilon}$ to ${}^{\epsilon}D_{\lambda}^{\epsilon}$ if $\lambda_{n+1}=0$, or $\lambda^{\bullet}$ to $\dot{D}_{\lambda}$ if $\lambda_{n+1}\neq 0$.
4.3. The geometric definition of the natural basis
Recall from Lemma 2.2 the bijective $\mathfrak{d}(\lambda,d,\mu)=(|R^{\lambda}_{i}\cap dR^{\mu}_{j}|)_{ij}$. If $\mathfrak{d}(\lambda,d,\mu)=A$, then we write $\mathfrak{d}^{-1}(A)=(\lambda,d_{A},\mu)$, where $\lambda=\text{\rm ro}(A)$, $\mu=\text{\rm co}(A)$. Note that $d_{A}\in{\check{W}}$ if and only if the entries of the upper right corner matrix $A^{\text{\normalsize$\llcorner$}}$ sum to an even number.
We now extend the map ${\mathfrak{d}}$ to the type $D$ case. Observe first that, for $A\in\Xi$ with $a_{n+1,n+1}\neq 0$, the associated distinguished double coset representative $d_{A}$ fixes $r$ and $r+1$. This implies that $s_{r}d_{A}s_{r}=d_{A}$.
Definition 4.11.
Define a map $\eta:\check{\Xi}(n,r)\rightarrow\check{\mathcal{D}}(n,r)$ by the following rules. For $A\in\Xi_{N,2r}$ such that ${\mathfrak{d}}(\lambda,d_{A},\mu)=A$ (so $\lambda=\text{\rm ro}(A)$ and $\mu=\text{\rm co}(A)$), let $W_{\lambda}d_{A}W_{\mu}$ be the associated double coset in $W=W^{\textsc{b}}$. We look for the corresponding double cosets in ${\check{W}}=W^{\textsc{d}}$ associated with $W_{\lambda}d_{A}W_{\mu}$.
  • (0)
    If $A\in\overset{\bullet}{\Xi}$, then we set $\eta(\dot{A})=({\lambda^{\bullet}},d,\mu^{\bullet})$ where
    $d=\begin{cases}\begin{aligned} &d_{A},&\text{ if }d_{A}\in\check{W},\\ &\text{the distinguished double coset representative for }\check{W}_{\lambda^{\bullet}}s_{r}d_{A}\check{W}_{\mu^{\bullet}},&\text{ if }d_{A}\notin\check{W}.\end{aligned}\end{cases}$
  • (1)
    If $A\in\overset{\circ\!\circ\!\circ}{\Xi}$, then, for $d_{A}\in{\check{W}}$, we set $\eta({{}^{+}{A}^{+}})=(\lambda^{+},d_{1},\mu^{+})$ and $\eta({{}^{-}{A}^{-}})=(\lambda^{-},d_{2},\mu^{-})$, where $d_{1}=d_{A}$ and $d_{2}$ is the distinguished double coset representative in $\check{W}_{\lambda^{-}}s_{r}d_{A}s_{r}\check{W}_{\mu^{-}}=s_{r}({\check{W}}_{\lambda}d_{A}{\check{W}}_{\mu})s_{r}$, and, for $d_{A}\not\in{\check{W}}$, we set $\eta({{}^{+}{A}^{-}})=(\lambda^{+},d_{1},\mu^{-})$ and $\eta({{}^{-}{A}^{+}})=(\lambda^{-},d_{2},\mu^{+})$, where $d_{1}$ and $d_{2}$ is the distinguished double coset representatives in $\check{W}_{\lambda^{+}}d_{A}s_{r}\check{W}_{\mu^{-}}$ and $\check{W}_{\lambda^{-}}s_{r}d_{A}\check{W}_{\mu^{+}}$, respectively.
  • (2)
    If $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$ then, for $d_{A}\in{\check{W}}$, we set $\eta({{}^{+}{A}^{+}})=(\lambda^{+},d_{1},\mu^{\bullet})$ and $\eta({{}^{-}{A}^{-}})=(\lambda^{-},d_{2},\mu^{\bullet})$, where $d_{1}=d_{A}$ and $d_{2}$ is the distinguished double coset representative for $\check{W}_{\lambda^{-}}s_{r}g_{A}s_{r}\check{W}_{\mu^{{\bullet}}}$, and, for $d_{A}\not\in{\check{W}}$, we set $\eta({{}^{+}{A}^{-}})=(\lambda^{+},d_{1},\mu^{\bullet})$ and $\eta({{}^{-}{A}^{+}})=(\lambda^{-},d_{2},\mu^{\bullet})$ where $d_{1}$ and $d_{2}$ is the distinguished double coset representative for $\check{W}_{\lambda^{+}}d_{A}s_{r}\check{W}_{\mu^{{\bullet}}}$ and $\check{W}_{\lambda^{-}}s_{r}d_{A}\check{W}_{\mu^{{\bullet}}}$ respectively.
  • (3)
    If $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$, then the definition is symmetric, replacing $\lambda^{+},\lambda^{-}$ by $\lambda^{\bullet}$ and $\mu^{\bullet}$ by $\mu^{+},\mu^{-}$ in (2).
  • (4)
    If $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$ then, for $d_{A}\in{\check{W}}$, we set $\eta({{}^{+}{A}^{+}})=({\lambda^{\bullet}},d_{1},\mu^{\bullet})$ and $\eta({{}^{-}{A}^{-}})=({\lambda^{\bullet}},d_{2},\mu^{\bullet})$, where $d_{1}=d_{A}$ and $d_{2}$ is the distinguished double coset representative for $\check{W}_{\lambda^{{\bullet}}}s_{r}d_{A}s_{r}\check{W}_{\mu^{{\bullet}}}$, and for $d_{A}\not\in{\check{W}}$, we set $\eta({{}^{+}{A}^{-}})=({\lambda^{\bullet}},d_{1},\mu^{\bullet})$ and $\eta({{}^{-}{A}^{+}})=({\lambda^{\bullet}},d_{2},\mu^{\bullet})$, where $d_{1}$ and $d_{2}$ is the distinguished double coset representative for $\check{W}_{\lambda^{{\bullet}}}d_{A}s_{r}\check{W}_{\mu^{{\bullet}}}$ and $\check{W}_{\lambda^{{\bullet}}}s_{r}d_{A}\check{W}_{\mu^{{\bullet}}}$ respectively.
The next lemma shows that $\eta$ is well-defined and bijective. See a different bijection from $\check{\mathcal{D}}(n,r)$ to $\check{\Xi}(n,r)$ in [19, Lem. A.4.1].
Lemma 4.12.
The map $\eta$ is well-defined and is a bijective map. Hence, we obtain a bijection
$\check{\mathfrak{d}}:=\eta^{-1}:\check{\mathcal{D}}(n,r)\longrightarrow\check{\Xi}(n,r).$
Proof.
We first prove that $\eta$ is a well-defined injective map.
(0) If $A\in\overset{\bullet}{\Xi}$, then, by Lemma 2.2(3) and Corollary 2.7(3), $s_{r}d_{A}=d_{A}s_{r}$, ${\check{W}}_{\lambda}=W_{\lambda}\cap{\check{W}}={\check{W}}_{\lambda^{\bullet}}$, and ${\check{W}}_{\mu}={\check{W}}_{\mu^{\bullet}}$. Thus,
$W_{\lambda}d_{A}W_{\mu}={\check{W}}_{\lambda^{\bullet}}d_{A}{\check{W}}_{\mu^{\bullet}}\sqcup s_{r}{\check{W}}_{\lambda^{\bullet}}d_{A}{\check{W}}_{\mu^{\bullet}}.$ (4.12.1)
One of the disjoint subsets is a double coset in ${\check{W}}$.
(1) If $A\in\overset{\circ\!\circ\!\circ}{\Xi}$ then, by Corollary 2.7(1) and (2.6.4), ${\check{W}}_{\lambda}={\check{W}}_{\lambda^{+}}=W_{\lambda}$, ${\check{W}}_{\mu}={\check{W}}_{\mu^{+}}=W_{\mu}$, and
$W_{\lambda}d_{A}W_{\mu}={\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{+}}.$ (4.12.2)
Thus, if $d_{A}\in{\check{W}}$ (resp., $d_{A}\not\in{\check{W}}$), then ${\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{+}}$ and ${\mathfrak{f}}({\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{+}})={\check{W}}_{\lambda^{-}}s_{r}d_{A}s_{r}{\check{W}}_{\mu^{-}}$ (resp., ${\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{+}}s_{r}={\check{W}}_{\lambda^{+}}d_{A}s_{r}{\check{W}}_{\mu^{-}}$ and ${\mathfrak{f}}({\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{+}}s_{r})={\check{W}}_{\lambda^{-}}s_{r}d_{A}{\check{W}}_{\mu^{+}}$) are distinct double cosets in ${\check{W}}$.
(2) If $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$ then, by Corollary 2.7, $W_{\lambda}={\check{W}}_{\lambda^{+}}$ and ${\check{W}}_{\mu}={\check{W}}_{\mu^{\bullet}}$,
$W_{\lambda}d_{A}W_{\mu}=({\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{\bullet}})\sqcup({\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{\bullet}}s_{r}).$ (4.12.3)
Thus, if $d_{A}\in{\check{W}}$ (resp., $d_{A}\not\in{\check{W}}$), then ${\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{\bullet}}$ and ${\mathfrak{f}}({\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{\bullet}})={\check{W}}_{\lambda^{-}}s_{r}d_{A}s_{r}{\check{W}}_{\mu^{\bullet}}$ (resp., ${\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{\bullet}}s_{r}={\check{W}}_{\lambda^{+}}d_{A}s_{r}{\check{W}}_{\mu^{\bullet}}$ and ${\mathfrak{f}}({\check{W}}_{\lambda^{+}}d_{A}{\check{W}}_{\mu^{\bullet}}s_{r})={\check{W}}_{\lambda^{-}}s_{r}d_{A}{\check{W}}_{\mu^{\bullet}}$) are distinct double cosets in ${\check{W}}$.
(3) This case is symmetric to (2).
(4) If $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$ then, by Corollary 2.7(3),
$W_{\lambda}d_{A}W_{\mu}=({\check{W}}_{\lambda^{\bullet}}d_{A}{\check{W}}_{\mu^{\bullet}})\sqcup({\check{W}}_{\lambda^{\bullet}}s_{r}d_{A}s_{r}{\check{W}}_{\mu^{\bullet}})\sqcup({\check{W}}_{\lambda^{\bullet}}d_{A}s_{r}{\check{W}}_{\mu^{\bullet}})\sqcup({\check{W}}_{\lambda^{\bullet}}s_{r}d_{A}{\check{W}}_{\mu^{\bullet}}).$ (4.12.4)
Thus, the two distinct double cosets in ${\check{W}}$ can be seen easily.
We now prove that $\eta$ is surjective.
Let $(\lambda^{{\epsilon}},d,\mu^{\iota})\in\check{\mathcal{D}}(n,r)$, where ${\epsilon},\iota\in\{{\bullet},+,-\}$ and $d\in\check{\mathcal{D}}_{\lambda^{{\epsilon}},\mu^{\iota}}$. We now show the existence of $A\in\check{\Xi}(n,r)$ satisfying $\eta(A)=(\lambda^{{\epsilon}},d,\mu^{\iota})$ corresponding to the cases in Definition 4.11.
Case (0)&(4). We first assume ${\epsilon}=\iota={\bullet}$ and consider the double coset ${\check{W}}_{\lambda^{\bullet}}d{\check{W}}_{\mu^{\bullet}}={\check{W}}_{\lambda}d{\check{W}}_{\mu}$ in ${\check{W}}$ which is contained in the double coset $W_{\lambda}dW_{\mu}$ in $W$. Suppose $A\in\Xi$ defines the double coset $W_{\lambda}dW_{\mu}=W_{\lambda}d_{A}W_{\mu}$. Then $\lambda=\text{\rm ro}(A)$, $\mu=\text{\rm co}(A)$, and $d_{A}$ is the distinguished double coset representative.
If $d_{A}s_{r}=s_{r}d_{A}$ then, by Lemma 2.2 (2), $a_{n+1,n+1}>0$. Thus, $A\in\overset{\bullet}{\Xi}$ and one of the two disjoint sets in (4.12.1) is a double coset in ${\check{W}}$, depending on $d_{A}\in{\check{W}}$ or not.
If $d_{A}s_{r}\neq s_{r}d_{A}$, then $a_{n+1,n+1}=0$. Thus, $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$, $W_{\lambda}d_{A}W_{\mu}$ has a decomposition as in (4.12.4) (and one of the two disjoint sets in (4.12.4) is a double coset, depending on $d_{A}\in{\check{W}}$ or not).
The remaining cases, where Case (1) for ${\epsilon},\iota\in\{+,-\}$ and Cases (2)&(3) for ${\epsilon}={\bullet}$ or $\iota={\bullet}$, but not both, we may use (4.12.2) or (4.12.3) to prove the existence similarly. ∎
Similar to the row/column sum vectors $\text{\rm ro}(A),\text{\rm co}(A)$ defined in (2.3.2), we now define row weights and column weights of elements in $\Xi(n,r)$:
$\text{\rm rw},\text{\rm cw}:\check{\Xi}(n,r)\longrightarrow\Lambda^{\textsc{d}}(n,r)$ (4.12.5)
by setting $\eta({\check{\mathbb{A}}})=(\text{\rm rw}({\check{\mathbb{A}}}),d,\text{\rm cw}({\check{\mathbb{A}}}))$, for any ${\check{\mathbb{A}}}\in\check{\Xi}(n,r)$.
Lemma 4.13.
Let ${\epsilon},{\epsilon}^{\prime}\in\{+,-\}$.
  1. (0)
    If ${\check{\mathbb{A}}}$ is associated with $A\in\overset{\bullet}{\Xi}$, then $\text{\rm rw}({\check{\mathbb{A}}}),\text{\rm cw}({\check{\mathbb{A}}})\in\Lambda^{\bullet}(n+1,r)$.
  2. (1)
    If ${\check{\mathbb{A}}}={}^{\epsilon}\!A^{{\epsilon}^{\prime}}$ with $A\in\overset{\circ\!\circ\!\circ}{\Xi}$, then $\text{\rm rw}({}^{\epsilon}\!A^{{\epsilon}^{\prime}})\in\Lambda^{\circ}_{\epsilon}(n+1,r)$ and $\text{\rm cw}({}^{\epsilon}\!A^{{\epsilon}^{\prime}})\in\Lambda^{\circ}_{{\epsilon}^{\prime}}(n+1,r)$.
  3. (2)
    If ${\check{\mathbb{A}}}={}^{\epsilon}\!A^{{\epsilon}^{\prime}}$ with $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$, then $\text{\rm rw}({}^{\epsilon}\!A^{{\epsilon}^{\prime}})\in\Lambda^{\circ}_{\epsilon}(n+1,r)$ and $\text{\rm cw}({}^{\epsilon}\!A^{{\epsilon}^{\prime}})\in\Lambda^{\bullet}(n+1,r)$.
  4. (3)
    If ${\check{\mathbb{A}}}={}^{\epsilon}\!A^{{\epsilon}^{\prime}}$ with $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$, then $\text{\rm rw}({}^{\epsilon}\!A^{{\epsilon}^{\prime}})\in\Lambda^{\bullet}(n+1,r)$ and $\text{\rm cw}({}^{\epsilon}\!A^{{\epsilon}^{\prime}})\in\Lambda^{\circ}_{\epsilon}(n+1,r)$.
  5. (4)
    If ${\check{\mathbb{A}}}={}^{\epsilon}\!A^{{\epsilon}^{\prime}}$ with $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$, then $\text{\rm rw}({\check{\mathbb{A}}}),\text{\rm cw}({\check{\mathbb{A}}})\in\Lambda^{\bullet}(n+1,r)$.
Note that only in case (0) and (1), the signs ($\pm$ or $\cdot$) on $A$ are inherited by the row/column weights.
Recalled the natural basis $\{\phi_{\alpha,\beta}^{d}\mid(\alpha,d,\beta)\in\check{\mathcal{D}}(n,r)\}$ for $S_{{\boldsymbol{q}}}^{\textsc{d}}(n,r)$ given in Corollary 4.3. With the bijection $\eta$ in Lemma 4.12, we may also label the basis element by signed matrix notation:
$\{\phi_{{\check{\mathbb{A}}}}\mid{\check{\mathbb{A}}}\in\check{\Xi}(n,r)\}\;\;\text{ where }\phi_{{\check{\mathbb{A}}}}=\phi_{\alpha\beta}^{d}\;\;\text{ if }\eta(\alpha,d,\beta)={\check{\mathbb{A}}}.$
We have also, for ${\check{\mathbb{A}}},{\check{\mathbb{B}}}\in\check{\Xi}(n,r)$,
$\phi_{\check{\mathbb{A}}}\phi_{\check{\mathbb{B}}}=0\;\;\text{ unless }\text{\rm cw}({\check{\mathbb{A}}})=\text{\rm rw}({\check{\mathbb{B}}}).$ (4.13.1)
Remark 4.14.
For $A\in\overset{\circ}{\Xi}$ such that $A={\mathfrak{d}}(\lambda,d_{A},\mu)$, denote ${\mathsf{m}}_{1}={\mathsf{m}}(\lambda)$ is the maximal index $i\leq n+1$ satisfying $\lambda_{i}\neq 0$. Let ${\mathsf{m}}_{2}={\mathsf{m}}(\mu)$. Define ${\mathsf{m}}$ to be the maximal index $j$ satisfying $a_{{\mathsf{m}}_{1},{\mathsf{m}}}\neq 0$ if ${\mathsf{m}}_{1}<n+1$, or $a_{n+1,{\mathsf{m}}}\neq 0$ if ${\mathsf{m}}_{1}=n+1$, and ${\mathsf{m}}\leq n$ (such ${\mathsf{m}}$ exists since $\lambda_{n+1}\neq 0$). We now use the subsets ${\check{R}}^{\lambda^{\epsilon}}_{i}$ defined in (2.5.1) to justify Definition 4.11 above.
(1) For $A\in\overset{\circ\!\circ\!\circ}{\Xi}$ and ${\epsilon}\in\{+,-\}$, if $A$ defines ${}^{+}{A}^{{\epsilon}},^{-}{A}^{-{\epsilon}}\in\overset{\circ\!\circ\!\circ}{\Xi}_{+}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}_{-}$, then $\eta(^{+}{A}^{{\epsilon}})=(\lambda^{+},d_{1},\mu^{\epsilon})$ and $\eta(^{-}{A}^{-{\epsilon}})=(\lambda^{-},d_{2},\mu^{-{\epsilon}})$ (see (6.1.1) for notational convention and Definition 4.11 for the definition of $d_{1},d_{2}$) which define distinct natural basis elements $\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{{}^{-}{A}^{-{\epsilon}}}$ according their row or column weights given in (4.12.5). This distinction can be reflected by the distinction of the ${\mathsf{m}}_{1}\times{\mathsf{m}}_{2}$ subset matrices $\big{(}{\check{R}}_{i}^{\lambda^{+}}\cap d_{1}{\check{R}}_{j}^{\mu^{\epsilon}}\big{)}$ and $\big{(}{\check{R}}_{i}^{\lambda^{-}}\cap d_{2}{\check{R}}_{j}^{\mu^{-{\epsilon}}}\big{)}$ since
${\check{R}}_{{\mathsf{m}}_{1}}^{\lambda^{+}}\cap d_{1}{\check{R}}_{{\mathsf{m}}}^{\mu^{\epsilon}}=I_{{\mathsf{m}}_{1},{\mathsf{m}}}\quad\text{ and }\quad{\check{R}}_{{\mathsf{m}}_{1}}^{\lambda^{-}}\cap d_{2}{\check{R}}_{{\mathsf{m}}}^{\mu^{-{\epsilon}}}=(I_{{\mathsf{m}}_{1},{\mathsf{m}}}\setminus\{r\})\cup\{r+1\}$
are distinct. However, as subsets of ${\check{W}}$, the double cosets ${\check{W}}_{\lambda^{+}}d_{1}{\check{W}}_{\mu^{\epsilon}}$ and ${\check{W}}_{\lambda^{-}}d_{2}{\check{W}}_{\mu^{-{\epsilon}}}$ as well as their conjugate intersections
${\check{W}}_{\lambda^{+}}\cap d_{1}{\check{W}}_{\mu^{\epsilon}}{d_{1}^{-1}}$ $=\bigcap_{i,j}\text{Stab}_{{\check{W}}}({\check{R}}_{i}^{\lambda^{+}}\cap d_{1}{\check{R}}_{j}^{\mu^{\epsilon}}),$
${\check{W}}_{\lambda^{-}}\cap d_{2}{\check{W}}_{\mu^{-{\epsilon}}}{d_{2}^{-1}}$ $=\bigcap_{i,j}\text{Stab}_{{\check{W}}}({\check{R}}_{i}^{\lambda^{-}}\cap d_{2}{\check{R}}_{j}^{\mu^{-{\epsilon}}}),$
may be the same.
Similar statement holds for $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$ with all $\mu^{\epsilon}$ and $\mu^{-{\epsilon}}$ replaced by $\mu^{\bullet}$. This also gives a similar statement for $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$ by symmetry (with ${\mathsf{m}}_{1}=n+1$ and ${\mathsf{m}}\leq n$).
(2) For $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$, we also have ${}^{+}{A}^{{\epsilon}},^{-}{A}^{-{\epsilon}}\in\overset{\bullet\!\circ\!\bullet}{\Xi}_{+}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}_{-}$. In this case, the natural basis elements $\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{{}^{-}{A}^{-{\epsilon}}}$ have the same domain and codomain, but are defined by distinct double cosets ${\check{W}}_{\lambda^{\bullet}}d_{1}{\check{W}}_{\mu^{\bullet}}$ and ${\check{W}}_{\lambda^{\bullet}}d_{2}{\check{W}}_{\mu^{\bullet}}$. Of course, the conjugate intersections ${\check{W}}_{\lambda^{\bullet}}\cap({\check{W}}_{\mu^{\bullet}})^{d_{1}^{-1}}$ and ${\check{W}}_{\lambda^{\bullet}}\cap({\check{W}}_{\mu^{\bullet}})^{d_{2}^{-1}}$ are also distinct in this case.
We are now ready to compute the dimension #$\#\check{\Xi}(n,r)$ of $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$.
Theorem 4.15.
Let ${\mathcal{X}}=\mathcal{F}_{n,2r}^{\jmath}$ and ${\check{G}}={\mathrm{SO}}_{2r}(q)$. Then there is an algebra isomorphism
$S^{\textsc{d}}_{q}(n,r):=S_{\boldsymbol{q}}^{\textsc{d}}(n,r)|_{{\boldsymbol{q}}=q}\cong\mathfrak{F}_{{\check{G}}}({\mathcal{X}}\times{\mathcal{X}})$
sending $\phi_{\check{\mathbb{A}}}|_{{\boldsymbol{q}}=q}$ to $f_{\check{\mathcal{O}}({\check{\mathbb{A}}})}$. Moreover, the following dimension formula holds
#$\#\check{\Xi}(n,r)=\binom{2n^{2}+2n+r}{r}+\binom{2n^{2}+2n+r-1}{r}.$ (4.15.1)
Proof.
By Lemma 3.1, the convolution algebra $\mathfrak{F}_{{\mathrm{SO}}_{2r}(q)}({\mathcal{X}}\times{\mathcal{X}})$ is isomorphic to the endomorphism algebra $\text{\rm End}_{{\check{G}}(q)}(\mathbb{Z}{\mathcal{X}})^{\text{\rm op}}$ of the permutation ${\check{G}}(q)$-module $\mathbb{Z}{\mathcal{X}}$.
We now calculate #$\#\check{\Xi}(n,r)$. By the definition of $\check{\Xi}(n,r)$ in (4.9.2), we have, for $N=2n+1$,
#$\#\check{\Xi}(n,r)-\#\Xi_{N,2r}=\#\{A\in\Xi_{N,2r}\mid a_{n+1,n+1}=0\}=\binom{2n^{2}+2n+r-1}{r}=\binom{n(N+1)+r-1}{r},$
(see Lemma 3.3). Moreover,
#$\#\Xi_{N,2r}$ =#$=\#\Big{\{}A=(a_{ij})\in\Xi_{N,2r}\;\Big{|}\;\sum\limits_{i\leq n:j}a_{ij}+\sum\limits_{j\leq n}a_{n+1,j}=r-\frac{a_{n+1,n+1}}{2}\Big{\}}$
$=\sum\limits_{l=0}^{r}\binom{2n^{2}+2n+r-l-1}{2n^{2}+2n-1}=\binom{2n^{2}+2n+r}{r}.$
Combining two equations gives #$\#\check{\Xi}(n,r)=\binom{2n^{2}+2n+r-1}{r}+\binom{2n^{2}+2n+r}{r}.$
Remark 4.16.
(1). The theorem shows that the ${\mathcal{A}}$-algebra $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ is the “quantumization” (in the sense of [8, p.17]) of the algebras $\{\mathfrak{F}_{{\mathrm{SO}}_{2r}(q)}({\mathcal{X}}\times{\mathcal{X}})\}_{q\in\mathcal{P}}$ which is the same as the algebra defined in [14, (5)]. It should be pointed out that the dimension formula #$\#\Xi_{\textsc{d}}$ given in [14, Lem. 4.2.1] is incorrect as the definition of the index set $\Xi_{\textsc{d}}=\Xi^{+}\sqcup\Xi^{0}\sqcup\Xi^{-}$ mistakenly used $\Xi^{0}$ instead of $\overset{\bullet}{\Xi}$. In fact, we have $\Xi^{0}=\overset{\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$. Consequently,
#$\#\check{\Xi}(n,r)=\#\Xi_{\textsc{d}}+\#\overset{\bullet\!\circ\!\bullet}{\Xi}.$ (4.16.1)
Indeed, this can be seen from $\overset{\circ}{\Xi}=\overset{\bullet\!\circ\!\bullet}{\Xi}\sqcup(\overset{\circ\!\circ\text{-}}{\Xi}\cup\overset{\text{-}\circ\!\circ}{\Xi})$, where $\overset{\circ\!\circ\text{-}}{\Xi}=\{A\in\Xi\mid\lambda_{n+1}=0=a_{n+1,n+1}\}$. Thus, #$\#\overset{\bullet\!\circ\!\bullet}{\Xi}=\#\overset{\circ}{\Xi}-2\overset{\circ\!\circ\text{-}}{\Xi}+\#\overset{\circ\!\circ\!\circ}{\Xi}=\binom{2n^{2}+2n+r-1}{r}-2\binom{2n^{2}+n+r-1}{r}+\binom{2n^{2}+r-1}{r}$, giving (4.16.1).
(2). The anonymous referee pointed out to us that by using coordinate algebra approach developed in [18], the dimension formula (4.15.1) of $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ was also obtained by Ziqing Xiang in an unpublished manuscript.
5. Multiplication formulas in the $({\boldsymbol{q}},1)$-Schur algebra $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$
We first follow the idea in [5] or [4] to derive the multiplication formulas in $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$. Due to the unequal parameter nature, we will see the differences between the structure constants here and those occurring in the (equal parameter) multiplication formulas for $S^{\jmath}$ in [4].
Let $\text{\rm iGr}(1,2r)=\text{\rm iGr}_{1}(\mathbb{F}_{q}^{2r})$ be the Grassmannian of isotropic lines in $\mathbb{F}_{q}^{2r}$:
$\text{\rm iGr}(1,2r)=\{\langle x\rangle\subseteq\mathbb{F}_{q}^{2r}\mid x\neq 0,\langle x,x\rangle_{\mathsf{J}}=0\}.$
The following result is known. Part (1) is stated in [14, Lem. 3.1.2]; while part (2) is [14, Lem. 3.1.3]. For completeness and later use in Corollary 5.3, we include a proof for part (1).
Lemma 5.1.
(1) The cardinality of the set $\text{\rm iGr}(1,2r)$ of isotropic lines in $\mathbb{F}_{q}^{2r}$ is
#$\#\text{\rm iGr}(1,2r)=\frac{\left(q^{r}-1\right)\left(q^{r-1}+1\right)}{q-1}=\frac{q^{2r-1}-1}{q-1}+q^{r-1}.$
(2) Let $F=({{F}_{i}})_{1\leq i\leq 5}\in\mathcal{F}_{2,r}^{\jmath}$ be a $2$-step isotropic flag with $\text{\bf dim}(\hat{F})=(a_{1},a_{2},a_{3},a_{4},a_{5})$ and, for $i=3,4$, let $Z_{i}=\{L\subset{F}_{i}\mid L\in\text{\rm iGr}(1,2r),L\not\subseteq{F}_{i-1}\}$. Then
$({i})\ \#Z_{3}=q^{a_{1}+a_{2}}\Big{(}\frac{q^{a_{3}-1}-1}{q-1}+q^{\frac{a_{3}}{2}-1}\Big{)},\qquad({ii})\ \#Z_{4}=q^{a_{1}+a_{2}+a_{3}-1}\frac{q^{a_{4}}-1}{q-1}.$
Proof.
We prove (1) (and (2) is given in loc. cit.). For $x=x_{1}e_{1}+\cdots+x_{r}e_{r}+x_{r+1}e_{r+1}+\cdots+x_{2r}e_{2r}\in\mathbb{F}_{q}^{2r}-\{0\}$, let
${\textbf{a}}={\bf a}(x)=(x_{1},x_{2},\cdots,x_{r}),\;\;{\textbf{b}}={\bf b}(x)=(x_{2r},x_{2r-1},\cdots,x_{r+1})\in\mathbb{F}_{q}^{r}.$
Then, $\mathbb{F}_{q}$ has odd characteristic implies
$\langle x,x\rangle_{\mathsf{J}}=0$ $\iff{\bf a}{\bf b}^{t}=0$
$\iff\text{either }{\bf a}=0,\text{ or }{\bf a}\neq 0,\text{ but }{\textbf{b}}^{t}\in\text{Ker }{\textbf{a}}.$
Thus,
$\text{\rm iGr}(1,2r)=\{\langle x\rangle\subseteq\mathbb{F}_{q}^{2r}\mid{\bf a}=0,{\bf b}\in\mathbb{F}_{q}^{r}-\{0\}\}\sqcup\{\langle x\rangle\subseteq\mathbb{F}_{q}^{2r}\mid{\bf a}\in\mathbb{F}_{q}^{r}-\{0\},{\bf b}^{t}\in\text{Ker}({\bf a})\}.$
Clearly, we have
#$\#\{\langle x\rangle\subseteq\mathbb{F}_{q}^{2r}\mid{\bf a}=0,{\bf b}\in\mathbb{F}_{q}^{r}-\{0\}\}=\frac{q^{r}-1}{q-1}.$
On the other hand, for ${\bf a}\neq 0$, since Ker $\text{Ker }{\textbf{a}}\cong\mathbb{F}_{q}^{r-1}$, it follows that
#$\#\{\langle x\rangle\subseteq\mathbb{F}_{q}^{2r}\mid{\bf a}\in\mathbb{F}_{q}^{r}-\{0\},{\bf b}^{t}\in\text{Ker}({\bf a})\}=\frac{q^{r}-1}{q-1}q^{r-1}.$
Hence,
#$\#\text{\rm iGr}(1,2r)=\frac{q^{r}-1}{q-1}+\frac{q^{r}-1}{q-1}\cdot q^{r-1}=\frac{\left(q^{r}-1\right)\left(q^{r-1}+1\right)}{q-1},$
as desired. ∎
Let $e_{A}=f_{{\mathcal{O}}_{A}}$ be the characteristic function of the ${\mathrm{O}_{2r}}(q)$-orbit corresponding to $A\in\Xi_{N,2r}$. Therefore, $\{e_{A}|A\in\Xi_{N,2r}\}$ forms a basis of $S^{\jmath}(n,r)$. For convenience, set $e_{A}=0,\text{ if }A\notin\Xi_{N,2r}.$
Now we formulate some multiplication formulas in $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$. These formulas are similar to (but different from) those in [4, Lem. 3.2], (cf. [14, 4.3.2(c)], [19, A.5.16]). Also, the proof, motivated from that of [14, Prop. 4.3.2], is much shorter than the one in [4]. Recall the matrices ${}_{h}\!A_{p}$ and ${}_{h}\!A_{\overline{p}}$ associated with $A$ defined in (3.7.1).
Theorem 5.2.
Maintain the notation above and suppose that $h\in[1,n]$, $N=2n+1$, and $A=(a_{i,j})\in\Xi_{N,2r}$.
  • (1)
    If $B=E_{h,h+1}^{\theta}+{\widehat{\lambda}}\in\Xi_{N,2r}$, for some $\lambda\in\Lambda(n+1,r)$, and $\operatorname{ro}(A)=\operatorname{co}(B)$, then we have in $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$
    $e_{B}*e_{A}=\sum_{p\in[1,N],a_{h+1,p}>0}{\boldsymbol{q}}^{\sum_{j>p}a_{h,j}}\left[\!\left[a_{h,p}+1\right]\!\right]e_{{}_{h}\!A_{p}}.$
  • (2)
    If $C=E_{h+1,h}^{\theta}+{\widehat{\mu}}\in\Xi_{N,2r}$, for some $\mu\in\Lambda(n+1,r)$, and $\operatorname{ro}(A)=\operatorname{co}(C)$, then, for $\delta=1-\delta_{a_{n,n+1},0}$, we have in $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$
    $e_{C}*e_{A}=\begin{cases}\sum\limits_{p\in[1,N],a_{h,p}>0}{\boldsymbol{q}}^{\sum_{j<p}a_{h+1,j}}\left[\!\left[a_{h+1,p}+1\right]\!\right]e_{{}_{h}\!A_{\overline{p}}},&\text{ if }h\neq n;\\ \sum\limits_{p\in[1,N],p\neq n+1\atop a_{n,p}>0}{\boldsymbol{q}}^{\sum\limits_{j<p}a_{n+1,j}}\left[\!\left[a_{n+1,p}+1\right]\!\right]e_{{}_{h}\!A_{\overline{p}}}+\delta{\boldsymbol{q}}^{\sum\limits_{j<n+1}a_{n+1,j}}\big{(}\!\left[\!\left[a_{n+1,n+1}+1\right]\!\right]+{\boldsymbol{q}}^{\frac{a_{n+1,n+1}}{2}}\big{)}e_{{}_{n}\!A_{\overline{n+1}}},&\text{ if }h=n.\end{cases}$
Proof.
We only prove (2). The proof consists of two cases: (1) $h<n$ or $h=n,p\leq n$ and (2) $h=n,p>n$. The first case is similar to the original type $A$ proof given in [5, Lem. 3.2], while the second case is parallel to that of [4, Lem. 3.2]. Note that the proof of (1) is similar to the first case in (2).
By the isomorphism $e_{A}|_{{\boldsymbol{q}}=q}\mapsto f_{A}=f_{{\mathcal{O}}_{A}}$ in Theorem 3.10, it suffices to compute the coefficients in the multiplication formula
$f_{C}*f_{A}=\sum_{A^{\prime}}\#Z_{C,A,A^{\prime}}f_{A^{\prime}}$ (5.2.1)
in (3.0.2), where, for $(F,F^{\prime})\in{\mathcal{O}}_{A^{\prime}}$,
$Z_{C,A,A^{\prime}}=\{E\in{\mathcal{F}}_{n,r}^{\jmath}\mid(F,E)\in{\mathcal{O}}_{C},(E,F^{\prime})\in{\mathcal{O}}_{A}\}.$ (5.2.2)
By Corollary 3.6(2), if #$\#Z_{C,A,A^{\prime}}\neq 0$ then $A^{\prime}={}_{h}\!A_{\overline{p}}=(a^{\prime}_{i,j})$ for some $p\in[1,N]$. We now compute the numbers #$\#Z_{C,A,A^{\prime}}.$
Let $Z_{h,p}=Z_{C,A,A^{\prime}}$. Suppose ${\mathfrak{m}}(F,F^{\prime})\in{\mathcal{O}}_{{}_{h}\!A_{\overline{p}}}$ and let $Z_{h}$ be the set of all isotropic subspaces $S$ such that $F_{h}\stackrel{{\scriptstyle 1}}{{\subset}}S\subseteq F_{h+1}$. Note that if $h<n$ then $F_{h+1}$ is isotropic. Thus, every subspace of $F_{h+1}$ is isotropic. Hence, $h<n$ implies #$\#Z_{h}=\#\{\text{1-dimensional subspaces of }F_{h+1}/F_{h}\}$.
[…]
$Z^{\prime}_{h,p}=\{S\in Z_{h}\mid F_{h}\cap F_{j}^{{}^{\prime}}=S\cap F_{j}^{{}^{\prime}}\text{ if }j<p\text{ and }F_{h}\cap F_{j}^{{}^{\prime}}\neq S\cap F_{j}^{{}^{\prime}}\text{ if }j\geq p\}.$
In particular, we have #$\#Z_{h,p}=\#Z^{\prime}_{h,p}$.
Note that if we write $S=F_{h}+\langle v\rangle$, then $F_{h}\cap F_{p-1}^{{}^{\prime}}=S\cap F_{p-1}^{{}^{\prime}}\iff v\not\in S\cap F_{p-1}^{{}^{\prime}}$ which is equivalent to $S\not\subseteq F_{h}+F_{h+1}\cap F^{\prime}_{p-1}$. Similarly, $F_{h}\cap F_{p}^{{}^{\prime}}\neq S\cap F_{p}^{{}^{\prime}}\iff v\in S\cap F_{p}^{{}^{\prime}}\iff S\subseteq F_{h}+F_{h+1}\cap F^{\prime}_{p}$. Thus, with the notation $F_{i,j}:=F_{i-1}+F_{i}\cap F^{\prime}_{j}$ used in the $n(N+1)$-step isotropic flag (3.2.2), we have
$Z^{\prime}_{h,p}=\{S\in Z_{h}\mid S\subseteq F_{h+1,p},S\not\subseteq F_{h+1,p-1}\}.$ (5.2.3)
If $h<n$, then $A$ and $A^{\prime}$ differ only at $a^{\prime}_{h,p}=a_{h,p}-1$, $a^{\prime}_{h+1,p}=a_{h+1,p}+1$. Thus,
#$\#Z_{h,p}$ =#$=\#\{S\in Z_{h}\mid S\subseteq F_{h+1,p}\}-\#\{S\mid S\subseteq F_{h+1,p-1}\}$
$=(q-1)^{-1}(q^{\text{\rm dim\,}(F_{h}+F_{h+1}\cap F_{p}^{\prime}/F_{h})}-q^{\text{\rm dim\,}(F_{h}+F_{h+1}\cap F_{p-1}^{\prime}/F_{h})})$
$=(q-1)^{-1}(q^{\sum_{j\leq p}a^{\prime}_{h+1,j}}-q^{\sum_{j\leq p-1}a^{\prime}_{h+1,j}})=q^{\sum_{j<p}a_{h+1,j}}\frac{q^{a_{h+1,p}+1}-1}{q-1}.$
If $h=n$, then $S\in Z_{n}$ may not be isotropic. However, $p\leq n$ implies $F_{n+1,p}=F_{n}+F_{n}^{\perp}\cap F^{\prime}_{p}$ is isotropic. Thus, every $S\in Z^{\prime}_{n,p}$ is isotropic. Hence, the counting formula above continues to hold for $h=n$ and $p\leq n$.
It remains to prove the case for $h=n$ and $p\geq n+1$. We extract the central section (the $(n+1)$th row) from (3.2.2):
$F_{n}\subseteq F_{n+1,1}\subseteq F_{n+1,2}\subseteq\cdots\subseteq F_{n+1,n}\subseteq F_{n+1,n+1}\subseteq\cdots\subseteq F_{n+1,N}=F_{n+1}.$
By Lemma 3.3, we have $F_{n+1,i}^{\perp}=F_{n+1,N-i}$, for all $i\in[1,n]$. Since $F_{n}^{\perp}=F_{n+1}$ and, for $W:=F_{n+1}/F_{n}$, $\langle\quad\;\rangle_{\mathsf{J}}|_{W}$ is equivalent to $\langle\quad\;\rangle_{{\mathsf{J}}_{2r^{\prime}}}$ by [14, Lem. 3.1.1], where $2r^{\prime}=\text{\rm dim\,}W$, the above filtration induces an $n$-step isotropic flag in $W$:
$0\subseteq W_{1}\subseteq W_{2}\subseteq\cdots\subseteq W_{N-1}\subseteq W_{N}=W\cong\mathbb{F}_{q}^{2r^{\prime}},$
where $W_{i}=\frac{F_{n}+F_{n+1}\cap F_{i}^{\prime}}{F_{n}}$ for $i\in[1,N]$.
Now we finish the computation with simple applications of Lemma  5.1(2).
If $p=n+1$, then consider the $2$-step isotropic flag
$0\overset{a_{1}}{\subseteq}0\overset{a_{2}}{\subseteq}W_{n}\overset{a_{3}}{\subseteq}W_{n+1}\overset{a_{4}}{\subseteq}W\overset{a_{5}}{\subseteq}W$
in $W\cong\mathcal{F}_{2,r^{\prime}}^{\jmath}$. Clearly, since $A^{\prime}=A-E^{\theta}_{n,n+1}+E^{\theta}_{n+1,n+1}$, we have $a_{1}=0$,
$a_{2}=\text{dim }W_{n}=\sum_{j<n+1}a_{n+1,j}^{\prime}=\sum_{j<n+1}a_{n+1,j},\;\text{ and }\;a_{3}=\text{dim }(W_{n+1}/W_{n})=a_{n+1,n+1}^{\prime}=a_{n+1,n+1}+2.$
Thus, (5.2.3) becomes in this case $Z^{\prime}_{n,n+1}=\{S\in Z_{n}\mid S\subseteq F_{n+1,n+1},S\not\subseteq F_{n+1,n}\}$. Hence, #$\#Z^{\prime}_{n,n+1}=\#Z_{3}$ where $Z_{3}=\{S\subset W_{n+1}\mid S\in\text{\rm iGr}(1,2r^{\prime}),S\not\subseteq W_{n}\}$. By Lemma  5.1(2)(i), we obtain
#$\#Z_{n,n+1}=q^{\sum_{j<n+1}a_{n+1,j}}\Big{(}\frac{q^{a_{n+1,n+1}+1}-1}{q-1}+q^{\frac{a_{n+1,n+1}}{2}}\Big{)},$
as desired.
Finally, suppose $p>n+1$. The computation is similar. Consider the $2$-step isotropic flag in $W$:
$0\overset{a_{1}}{\subseteq}W_{N-p}\overset{a_{2}}{\subseteq}W_{N-p+1}\overset{a_{3}}{\subseteq}W_{p-1}\overset{a_{4}}{\subseteq}W_{p}\overset{a_{5}}{\subseteq}W_{N}\cong\mathbb{F}_{q}^{2r^{\prime}}.$
Here, with $A^{\prime}=A-E^{\theta}_{n,p}+E^{\theta}_{n+1,p}$ ($p>n+1$), we have
$a_{1}+a_{2}+a_{3}=\text{dim }W_{p-1}=\sum_{j<p}a_{n+1,j}^{\prime}=\sum_{j<p}a_{n+1,j}\;\text{ and }\;a_{4}=\text{dim }(W_{p}/W_{p-1})=a_{n+1,p}^{\prime}=a_{n+1,p}+1.$
Thus, by Lemma 5.1(2)(ii), #$\#Z^{\prime}_{n,p}=\#Z_{4}$, where $Z_{4}=\{S\subset W_{p}\mid S\in\text{\rm iGr}(1,2r^{\prime}),S\not\subseteq W_{p-1}\}$. Hence, for $p>n+1$,
#$\#Z_{n,p}=q^{\sum_{j<p}a_{n+1,j}}\frac{q^{a_{n+1,p}+1}-1}{q-1}.$
The theorem is proven. ∎
The proof above implies immediately the following which will be used in §7.
Corollary 5.3.
Maintain the notation on $A,C,A^{\prime}={}_{h}\!A_{\overline{p}}$ etc. set in Theorem 3.8. Suppose $(F,F^{\prime})\in{\mathcal{O}}_{A^{\prime}}$ and $\alpha=\text{\rm ro}(A^{\prime})$. If $\{v_{1},v_{2},\ldots,v_{2r}\}$ forms a basis for $\mathbb{F}^{2r}$ such that $F_{i}=\langle v_{1},v_{2},\ldots,v_{\widetilde{\alpha}_{i}}\rangle$ for all $i\in[1,N]$, then the set $Z_{C,A,A^{\prime}}$ defined in (5.2.2) has the following description
$Z_{C,A,A^{\prime}}=\begin{cases}\{F_{h}+\langle v\rangle\mid v=\sum_{i=1}^{m}b_{i}v_{\widetilde{\alpha}_{h}+i}+\sum_{j=1}^{a_{h+1,p}}b^{\prime}_{j}v_{\widetilde{\alpha}_{h}+m+j}\},&\text{ if }h<n\text{ or }h=n,p\neq n+1;\\ \{F_{h}+\langle v\rangle\mid v=\sum_{i=1}^{m}c_{i}v_{\widetilde{\alpha}_{n}+i}+\sum_{j=1}^{a_{h+1,n+1}}c^{\prime}_{j}v_{\widetilde{\alpha}_{n}+m+j}\},&\text{ if }h=n,p=n+1,\end{cases}$
where $m=\begin{cases}a_{h+1,1}\cdots+a_{h+1,p-1},&\text{ if }h<n\text{ or }h=n,p\neq n+1;\\ a_{n+1,1}\cdots+a_{n+1,n},&\text{ if }h=n,p=n+1,\end{cases}$ and the coefficients $b^{\prime}_{j}$ satisfy that not all $b_{j}^{\prime}=0$, and $c^{\prime}_{j}$ satisfy that either ${\bf c}=(c^{\prime}_{r},c^{\prime}_{r-1},\ldots,c^{\prime}_{r-\frac{a}{2}+1})\neq{\bf 0}$ $(a=a_{n+1,n+1})$ with $(c^{\prime}_{r+1},\ldots,c^{\prime}_{r+\frac{a}{2}})\in\text{\rm ker}({\bf c})$, or ${\bf c}={\bf 0}$ but not all $c^{\prime}_{r+1},\ldots,c^{\prime}_{r+\frac{a}{2}}=0$.
6. Multiplication formulas in the ${\boldsymbol{q}}$-Schur algebra $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$, I
We now use the multiplication formulas given in Theorem 5.2 to derive the corresponding multiplication formulas in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$. We first look for the counterpart of Theorem 5.2(1)
$e_{E^{\theta}_{h,h+1}+{\widehat{\gamma}}}*e_{A}=\delta_{{\widehat{\gamma}},\text{\rm ro}(A)-e^{\theta}_{h+1}}\sum_{1\leq p\leq N}g_{h,A,p}e_{{}_{h}\!A_{p}},$ (6.0.1)
where $h\in[1,n]$, $\gamma\in\Lambda(n+1,r)$, $A=(a_{i,j})\in\Xi_{N,2r}$, ${}_{h}\!A_{p}=A+E_{h,p}^{\theta}-E^{\theta}_{h+1,p}$ for $p\in[1,N]$, and
$g_{h,A,p}=\begin{cases}{\boldsymbol{q}}^{\sum_{j>p}a_{h,j}}\left[\!\left[a_{h,p}+1\right]\!\right],&\text{ if }a_{h+1,p}>0;\\ 0,&\text{ if }a_{h+1,p}=0.\end{cases}$ (6.0.2)
Remark 6.1.
Observe from (4.6.1) that, if an ${\text{\rm O}}_{2r}$-orbit ${\mathcal{O}}_{A}$ in ${\mathcal{X}}\times{\mathcal{X}}$ splits into two ${\text{\rm SO}}_{2r}$-orbits $\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\sqcup\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}})$, then the orbit function $f_{A}=f_{{}^{+}\!A^{\epsilon}}+f_{{}^{-}\!A^{-{\epsilon}}}\in\mathfrak{F}_{\check{G}}({\mathcal{X}}\times{\mathcal{X}})$. Thus, we have a $\mathbb{Z}$-algebra embedding $\mathfrak{F}_{G}({\mathcal{X}}\times{\mathcal{X}})\subset\mathfrak{F}_{\check{G}}({\mathcal{X}}\times{\mathcal{X}})$ which is the specialization of the ${\mathcal{A}}$-algebra embedding $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\subset S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ by Theorem 3.10. Composing with the isomorphism in Proposition 4.2 gives an algebra embedding $\iota:S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\hookrightarrow S_{{\boldsymbol{q}}}^{\textsc{d}}(n,r)$. We make the convention that, when we write $e_{A}=\phi_{{}^{+}\!A^{\epsilon}}+\phi_{{}^{-}\!A^{-{\epsilon}}}$ for ${\epsilon}\in\{+,-\}$, we mean that $\iota(e_{A})=\phi_{{}^{+}\!A^{\epsilon}}+\phi_{{}^{-}\!A^{-{\epsilon}}}$. Here and below we use the sign convention
++=+=,+==+.++=+=--$\qquad$+-=-=-+.+ + = + = - - , + - = - = - + . (6.1.1)
Now, in (6.0.1), let $B=E^{\theta}_{h,h+1}+{\widehat{\gamma}}$ and assume $\lambda=\text{\rm ro}(B)$ and $\mu=\operatorname{ro}(A)=\operatorname{co}(B)$. Then the upper right corner matrix $B^{\text{\normalsize$\llcorner$}}=0$ (and so, $\text{sgn}(B^{\text{\normalsize$\llcorner$}})=+$). Thus, by Proposition 4.5 and notations in Definitions 4.6 and 4.8, we have
${\mathcal{O}}_{B}=\begin{cases}\begin{aligned} \check{\mathcal{O}}(^{+}{B}^{+})\sqcup\check{\mathcal{O}}(^{-}{B}^{-}),&\text{ if }b:=b_{n+1,n+1}=0,\\ \check{\mathcal{O}}(\dot{B}),\qquad\ &\text{ otherwise.}\end{aligned}\end{cases}$
Hence, we have
$\begin{cases}B\in\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet}{\Xi},&\text{ if }h<n;\\ B\in\overset{\circ\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet}{\Xi},&\text{ if }h=n.\end{cases}$
Since we only need to consider $A=(a_{i,j})\in\Xi_{N,2r}$ with $\mu=\text{\rm ro}(A)$ in (6.0.1), the three selections of $B$ above determine the column weights $\text{\rm cw}({\check{\mathbb{B}}})$ of ${\check{\mathbb{B}}}\in\{^{+}\!B^{+},{}^{-}\!B^{-},\dot{B}\}$, which further determine certain selections of $A$ and, consequently, of ${}_{h}\!A_{p}$. We now derive multiplication formulas in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ associated with the three selections of $B$ in three subsections below.
Observe first the following general relations. Since ${}_{h}\!A_{p}=A+E^{\theta}_{h,p}-E^{\theta}_{h+1,p}$ (and ${}_{h}\!A_{\overline{p}}=A-E_{h,p}^{\theta}+E^{\theta}_{h+1,p}$), it follows that
$\text{sgn}({}_{h}\!A_{p}^{\text{\normalsize$\llcorner$}})(=\text{sgn}({}_{h}\!A_{\overline{p}}^{\text{\normalsize$\llcorner$}}))=\begin{cases}\text{sgn}(A^{\text{\normalsize$\llcorner$}}),\ &\text{ if $h<n$ or $h=n$ and $1\leq p\leq n+1$},\\ -\text{sgn}(A^{\text{\normalsize$\llcorner$}}),\ &\text{ if $h=n$ and $p>n+1$}.\end{cases}$ (6.1.2)
Further, if $a:=a_{n+1,n+1}=0$ and ${\epsilon}=\text{sgn }A^{\text{\normalsize$\llcorner$}}$, then we have $G$-orbits ${\mathcal{O}}_{A},{\mathcal{O}}_{{}_{h}\!A_{p}}$ splitting into ${\check{G}}$-orbits:
$(1)\qquad{\mathcal{O}}_{A}$ $=\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\sqcup\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}}),$ (6.1.3)
$(2)\;\,\quad{\mathcal{O}}_{{}_{h}\!A_{p}}$ =$=\begin{cases}\check{\mathcal{O}}(^{+}{{}_{h}\!A_{p}}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}),&\text{ if $h<n$ or $h=n$ and $1\leq p\leq n+1$},\\ \check{\mathcal{O}}(^{+}{{}_{h}\!A_{p}}^{-{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{h}\!A_{p}}^{{\epsilon}}),&\text{ if $h=n$ and $p>n+1$}.\end{cases}$
For simplicity, we fix the following notational abbreviation for the central entries of $A$ and $B$:
$a:=a_{n+1,n+1}=0,\qquad b:=b_{n+1,n+1}.$
6.1. The $B\in\overset{\bullet}{\Xi}$ case
Since ${\mathcal{O}}_{B}=\check{\mathcal{O}}(\dot{B})$ in this case and $\text{\rm cw}(\dot{B})=\mu^{\bullet}\in\Lambda^{\bullet}(n+1,r)$, (4.13.1) and Lemma 4.13 imply that it suffices to consider those $A$ with $A\in\overset{\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$.
Theorem 6.2.
Let $A,B\in\Xi_{N,2r}$ with $B=E^{\theta}_{h,h+1}+{\widehat{\gamma}}$, and assume ${\widehat{\lambda}}=\text{\rm ro}(B)$, ${\widehat{\mu}}=\text{\rm co}(B)=\text{\rm ro}(A)$, and ${\widehat{\nu}}=\text{\rm co}(A)$, where $h\in[1,n]$, and $\gamma,\lambda,\mu,\nu\in\Lambda(n+1,r)$. Let ${\epsilon}=\text{\rm sgn }(A^{\text{\normalsize$\llcorner$}})$. If $B\in\overset{\bullet}{\Xi}$, then ${\mathcal{O}}_{B}=\check{\mathcal{O}}(\dot{B})$ and the following multiplication formulas hold in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$:
  • (1)
    If $A\in\overset{\bullet}{\Xi}$, then ${\mathcal{O}}_{A}=\check{\mathcal{O}}(\dot{A})$ and
    $\phi_{\dot{B}}*\phi_{\dot{A}}=\begin{cases}\sum_{1\leqslant p\leqslant N}g_{h,A,p}\phi_{{}_{h}\!{\dot{A}}_{p}},&\text{ if $h<n$ or $h=n$ and $a\neq 2$},\\ \\ \sum\limits_{1\leq p\leq N\atop p\neq n+1}g_{h,A,p}\phi_{{}_{h}\!{\dot{A}}_{p}}+g_{n,A,n+1}(\phi_{{}^{+}{(_{n}\!A_{n+1})}^{{\epsilon}}}+\phi_{{}^{-}{(_{n}\!A_{n+1})}^{-{\epsilon}}}),&\text{ if $h=n$ and $a=2$};\end{cases}$
  • (2)
    If $A\in\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$, then
    $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}=$$\begin{cases}\sum_{1\leq p\leq N}g_{h,A,p}\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}},&\text{ if }h<n,\\ \sum\limits_{1\leq p<n+1}\!g_{h,A,p}\phi_{{}^{+}{({}_{h}\!A_{p})}^{{\epsilon}}}+\!\sum\limits_{p>n+1}\!g_{h,A,p}\phi_{{}^{-}{({}_{h}\!A_{p})}^{{\epsilon}}},&\text{ if }h=n,\end{cases}$ (6.2.1)
    $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=$$\begin{cases}\sum_{1\leq p\leq N}g_{h,A,p}\phi_{{}^{-}{({}_{h}\!A_{p})}^{-\epsilon}},&\text{ if }h<n,\\ \sum\limits_{1\leq p<n+1}\!g_{h,A,p}\phi_{{}^{-}{({}_{h}\!A_{p})}^{-{\epsilon}}}+\sum\limits_{p>n+1}g_{h,A,p}\phi_{{}^{+}{({}_{h}\!A_{p})}^{-{\epsilon}}},&\text{ if }h=n.\end{cases}$
Here all coefficients $g_{h,A,p}$ are given in (6.0.2).
Proof.
(1) For $A\in\overset{\bullet}{\Xi}$, by the hypotheses, we have $(\text{\rm rw}(\dot{A}),\text{\rm cw}(\dot{A}))=(\mu^{\bullet},\nu^{\bullet})$. Since, for $h<n$ or $h=n$ and $a_{n+1,n+1}\neq 2$, the central entry of ${}_{h}\!A_{p}$ is nonzero, it follows that ${\mathcal{O}}_{{}_{h}\!A_{p}}=\check{\mathcal{O}}(\dot{}_{h}\!A_{p})$ is non-split. Hence, we have $e_{B}=\phi_{\dot{B}}$, $e_{A}=\phi_{\dot{A}}$, and $e_{{}_{h}\!A_{p}}=\phi_{\dot{}_{h}\!A_{p}}$, the first case in (1) follows immediately from (6.0.1).
If $h=n$ and $a_{n+1,n+1}=2$, then all $G(q)$-orbits involved are non-split except ${\mathcal{O}}_{{}_{n}\!A_{n+1}}$ which splits into two ${\check{G}}(q)$-orbits by Proposition 4.5. Since $\text{sgn}({{}_{n}A_{n+1}}^{\text{\normalsize$\llcorner$}})=\text{sgn}(A^{\llcorner})$ by (6.1.2), we have ${\mathcal{O}}_{{}_{n}\!A_{n+1}}=\check{\mathcal{O}}({{}^{+}{{}_{n}A_{n+1}}^{{\epsilon}}})\sqcup\check{\mathcal{O}}({{}^{-}{{}_{n}A_{n+1}}^{-{\epsilon}}})$ and $\text{\rm cw}({{}^{+}{{}_{n}A_{n+1}}^{{\epsilon}}})=\nu^{\bullet}=\text{\rm cw}({{}^{-}{{}_{n}A_{n+1}}^{-{\epsilon}}})$. Thus, $e_{{}_{n}\!A_{n+1}}=\phi_{{{}^{+}{{}_{n}A_{n+1}}^{{\epsilon}}}}+\phi_{{{}^{-}{{}_{n}A_{n+1}}^{-{\epsilon}}}}$ in $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$. Substituting gives the second case in (1).
(2) For $A\in\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$, applying the orbit splittings in (6.1.3) to (6.0.1) yields
$\phi_{\dot{B}}*(\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{A}^{-{\epsilon}}})=\begin{cases}\sum_{1\leq p\leq N}g_{h,A,p}(\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}+\phi_{{}^{-}{({}_{h}\!A_{p})}^{-\epsilon}}),&\text{ if }h<n,\\ \sum\limits_{1\leq p<n+1}\!g_{h,A,p}(\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}+\phi_{{}^{-}{({}_{h}\!A_{p})}^{-\epsilon}})+\!\sum\limits_{p>n+1}\!g_{h,A,p}(\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}+\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}),&\text{ if }h=n,\end{cases}$ (6.2.2)
where we omitted the $p=n+1$ term in the $h=n$ case since the central entry of ${}_{n}A_{n+1}=-2$. We want to separate (6.2.2) into two formulas for $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$.
The case for $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$ is easy since this implies all ${}_{h}\!A_{p}\in\overset{\bullet\!\circ\!\circ}{\Xi}$. Thus,
$\text{\rm cw}(^{+}{A}^{{\epsilon}})=\text{\rm cw}(^{+}{({}_{h}\!A_{p})}^{\epsilon})=\text{\rm cw}(^{-}{{}_{h}\!A_{p}}^{{\epsilon}})=\nu^{\epsilon},\;\;\;\text{\rm cw}(^{-}{A}^{-{\epsilon}})=\text{\rm cw}(^{-}{{}_{h}\!A_{p}}^{-{\epsilon}})=\text{\rm cw}(^{+}{{}_{h}\!A_{p}}^{-{\epsilon}})=\nu^{-{\epsilon}}.$
Hence, by (4.13.1), multiplying the idempotent $1_{\nu^{\epsilon}}$ (resp., $1_{\nu^{-{\epsilon}}}$) to the right hand side of (6.2.2) yields the required formula for $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ (resp., $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$).
The case for $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$ cannot be obtained by an idempotent argument as above since all terms on the right hand side of (6.2.2) have a sole column weight $\nu^{\bullet}$. We need to determine the terms by the definition of convolution products in (3.0.2). In other words, to compute $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ and $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$, it requires to compute the convolution products $f_{\check{\mathcal{O}}(\dot{B})}*f_{\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})}$ and $f_{\check{\mathcal{O}}(\dot{B})}*f_{\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}})}$ in $\mathfrak{F}^{\jmath}_{{\check{G}}(q)}({\mathcal{X}}\times{\mathcal{X}})$ via (3.0.2). The following claim is sufficient to complete the proof.
  • Claim:
    For $h\in[1,n]$ and $p\in[1,N]$, if $F,F^{\prime},E\in\mathcal{F}_{n,r}^{\jmath}$ satisfy ${\mathfrak{m}}({F},E)=B$, ${\mathfrak{m}}(E,F^{\prime})=A$, and ${\mathfrak{m}}({F},F^{\prime})={}_{h}\!A_{p}$, then
    $(E,F^{\prime})\in{\check{\mathcal{O}}(^{{\epsilon}_{1}}\!A^{{\epsilon}_{2}})}\implies(F,F^{\prime})\in\begin{cases}\check{\mathcal{O}}(^{{\epsilon}_{1}}\!{{}_{h}\!A_{p}}^{{\epsilon}_{2}}),&\text{ if $h<n$ or $h=n$ and $p\leq n$},\\ \check{\mathcal{O}}(^{-{\epsilon}_{1}}\!{{}_{h}\!A_{p}}^{{\epsilon}_{2}}),&\text{ if $h=n$ and $p\geq n+2$}.\end{cases}$ (6.2.3)
By the claim, we see that $\phi_{\dot{B}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ (resp., $\phi_{\dot{B}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$) is a linear combination of $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ (resp., $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$) if $h<n$, or a linear combination of $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ and $\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}$ (resp., $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$ and $\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}$) if $h=n$. Note that, as natural basis elements in $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined by distinct $({\check{W}}_{\lambda^{\bullet}},{\check{W}}_{\mu^{\bullet}})$ double cosets (see Definition 4.11(4)), $\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}}$ and $\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}$ (or $\{\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}},\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}\}$ and $\{\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}},\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}\}$) are linearly independent. Hence, equating with (6.2.2) gives (6.2.1).
It remains to prove the claim. Since $a=a_{n+1,n+1}=0$ and ${\mathfrak{m}}(E,F^{\prime})=A$, $(E,F^{\prime})\in{\check{\mathcal{O}}(^{{\epsilon}_{1}}\!A^{{\epsilon}_{2}})}$ implies, by Lemma 4.7, that ${\epsilon}_{1}1=(-1)^{d_{E,F^{\prime}}-r}$ and ${\epsilon}_{2}1=(-1)^{d_{F^{\prime},E}-r}$.
[…]
$F_{n}+F_{n+1}\cap F_{n}^{{}^{\prime}}=E_{n}+E_{n+1}\cap F_{n}^{{}^{\prime}},\ F^{{}^{\prime}}_{n}+F^{{}^{\prime}}_{n+1}\cap F_{n}={F}^{{}^{\prime}}_{n}+{F}^{{}^{\prime}}_{n+1}\cap E_{n}.$ (6.2.4)
Thus, $d_{F,F^{\prime}}=d_{E,F^{\prime}}$ and $d_{F^{\prime},F}=d_{F^{\prime},E}$. Hence, $(F,F^{\prime})\in\check{\mathcal{O}}({}^{\epsilon_{1}}{}_{h}\!A_{p}^{\epsilon_{2}})$.
Now assume $h=n$ and $F_{n}=E_{n}+\langle e\rangle$. Then $e\in E_{n+1}$ since $F_{n}^{\perp}=F_{n+1}=E_{n+1}=E_{n}^{\perp}$. Note also that the hypothesis in the Claim implies that the index $p$ satisfies the conditions:
$F_{n}\cap F_{j}^{{}^{\prime}}=E_{n}\cap F_{j}^{{}^{\prime}}\text{ if }j<p\text{ and }F_{n}\cap F_{j}^{{}^{\prime}}\neq E_{n}\cap F_{j}^{{}^{\prime}}\text{ if }j\geq p.$ (6.2.5)
Thus, if $p\leq n$, then $F_{n}\cap F_{n}^{{}^{\prime}}\neq E_{n}\cap F_{n}^{{}^{\prime}}$ and $F_{n}\cap F_{n+1}^{{}^{\prime}}\neq E_{n}\cap F_{n+1}^{{}^{\prime}}$. Hence, $e\in F^{{}^{\prime}}_{n}$ and consequently, (6.2.4) continue to hold. This implies $(F,F^{\prime})\in\check{\mathcal{O}}(^{{\epsilon}_{1}}{{}_{n}\!A_{p}}\!^{{{\epsilon}_{2}}})$ for all $p\leq n$.
Finally, assume $p\geq n+2$. Then $F_{n}\cap F_{n+1}^{{}^{\prime}}=E_{n}\cap F_{n+1}^{{}^{\prime}}$ and so, $e\notin F^{{}^{\prime}}_{n+1}$. Thus, $F^{{}^{\prime}}_{n+1}\cap F_{n}={F}^{{}^{\prime}}_{n+1}\cap E_{n}$. Hence,
$e\in(F_{n}+F_{n+1}\cap F_{n}^{{}^{\prime}})-(E_{n}+E_{n+1}\cap F_{n}^{{}^{\prime}}),\text{ and }F^{{}^{\prime}}_{n}+F^{{}^{\prime}}_{n+1}\cap F_{n}={F}^{{}^{\prime}}_{n}+{F}^{{}^{\prime}}_{n+1}\cap E_{n}.$
Since the central entries of $A$ and ${}_{h}\!A_{p}$ are 0, this display continues to hold if $F^{\prime}_{n}$ is replaced by $F^{\prime}_{n+1}$. Hence, we have $d_{F,F^{\prime}}=d_{E,F^{\prime}}+1$ and $d_{F^{\prime},F}=d_{F^{\prime},E}$. (This can be seen by taking $(F,E)=(F^{\widehat{\lambda}},F^{\widehat{\mu}})$ and $(F,F^{\prime})=(F^{\widehat{\lambda}},F^{A^{\prime}})$. Thus, $F_{n}=E_{n}+\langle e\rangle$ with $e=e_{m},m\in I^{\prime}_{n,p}$; see Proposition 3.8.) Consequently, $(F,F^{\prime})\in\check{\mathcal{O}}(^{-{\epsilon}_{1}}{{}_{n}\!A_{p}}\!^{{{\epsilon}_{2}}})$ for all $p>n+1$.∎
6.2. The case $B\in\overset{\circ\!\circ\!\circ}{\Xi}$ ($\implies h<n$)
In this case, ${\mathcal{O}}_{B}=\check{\mathcal{O}}(^{+}\!B^{+})\sqcup\check{\mathcal{O}}(^{-}\!B^{-})$. Thus, it suffices to consider those $A$ in (6.0.1) with $A\in\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}$.
Theorem 6.3.
Maintain the same assumptions on $A,B,h,\lambda,\mu,\nu$, and ${\epsilon}=\text{\rm sgn }(A^{\text{\normalsize$\llcorner$}})$ as in Theorem 6.2, and assume $B=E^{\theta}_{h,h+1}+{\widehat{\gamma}}\in\overset{\circ\!\circ\!\circ}{\Xi}$. Then $h<n$ and, for $A\in\overset{\circ\!\circ\!\bullet}{\Xi}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}$, the following multiplication formulas hold in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$:
$\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}=\sum_{1\leq p\leq N}g_{h,A,p}\phi_{{}^{+}{({}_{h}\!A_{p})}^{\epsilon}},\qquad\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=\sum_{1\leq p\leq N}g_{h,A,p}\phi_{{}^{-}{({}_{h}\!A_{p})}^{-\epsilon}}.$ (6.3.1)
Proof.
The assertion $h<n$ is clear since the central component of $\text{\rm co}(B)$ is at least 2 when $h=n$. By the hypothesis, we have $(\text{\rm rw}(^{+}{B}^{+}),\text{\rm cw}(^{+}{B}^{+}))=(\lambda^{+},\mu^{+})$ and $(\text{\rm rw}(^{-}{B}^{-}),\text{\rm cw}(^{-}{B}^{-}))=(\lambda^{-},\mu^{-})$. If $A\in\overset{\circ\!\circ\!\bullet}{\Xi}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}$, then ${\mathcal{O}}_{A}=\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}})$ with
$(\text{\rm rw}(^{+}{A}^{{\epsilon}}),\text{\rm cw}(^{+}{A}^{{\epsilon}}))=\begin{cases}(\mu^{+},\nu^{{\epsilon}}),&\text{ if }A\in\overset{\circ\!\circ\!\circ}{\Xi};\\ (\mu^{+},\nu^{\bullet}),&\text{ if }A\in\overset{\circ\!\circ\!\bullet}{\Xi}.\end{cases}\qquad(\text{\rm rw}(\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}}),\text{\rm cw}(\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}})))=\begin{cases}(\mu^{-},\mu^{-{\epsilon}}),&\text{ if }A\in\overset{\circ\!\circ\!\circ}{\Xi};\\ (\mu^{-},\nu^{\bullet}),&\text{ if }A\in\overset{\circ\!\circ\!\bullet}{\Xi}.\end{cases}$
By (6.1.3)(2), (6.0.1) has the following decomposition in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
$(\phi_{{}^{+}{B}^{+}}+\phi_{{}^{-}{B}^{-}})*(\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{A}^{-{\epsilon}}})=\sum_{1\leq p\leq N}g_{h,A,p}(\phi_{{}^{+}{{}_{h}\!A_{p}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}}).$ (6.3.2)
Clearly, LHS$=\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=0=\phi_{{}^{-}{B}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}$. Also, every natural basis element $\phi_{{\check{\mathbb{A}}}^{\prime}}$ on the right hand side has row weight $\text{\rm rw}({\check{\mathbb{A}}}^{\prime})=\lambda^{+}$ or $\lambda^{-}$. Thus, (6.3.1) is obtained by multiplying idempotents $1_{\lambda^{+}}$ or $1_{\lambda^{-}}$ to the left hand side of (6.3.2). ∎
6.3. The case $B\in\overset{\circ\!\circ\!\bullet}{\Xi}$($\implies h=n$)
We have in this case ${\mathcal{O}}_{B}=\check{\mathcal{O}}(^{+}\!B^{+})\sqcup\check{\mathcal{O}}(^{-}\!B^{-})$ with $(\text{\rm rw}(^{+}\!B^{+}),\text{\rm cw}(^{+}\!B^{+}))=(\lambda^{+},\mu^{\bullet})$ and $(\text{\rm rw}(^{-}\!B^{-}),\text{\rm cw}(^{-}\!B^{-}))=(\lambda^{-},\mu^{\bullet})$. Thus, it suffices to consider those $A$ with $A\in\overset{\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}$.
Theorem 6.4.
Let $A,B\in\Xi_{N,2r}$ and ${\epsilon}=\text{\rm sgn }(A^{\text{\normalsize$\llcorner$}})$, and assume $\text{\rm co}(B)=\text{\rm ro}(A)$, $B\in\overset{\circ\!\circ\!\bullet}{\Xi}$, and $B-E^{\theta}_{n,n+1}$ is diagonal. Then the following multiplication formulas hold in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$:
  • (1)
    If $A\in\overset{\bullet}{\Xi}$, then ${\mathcal{O}}_{A}=\check{\mathcal{O}}(\dot{A})$, and
    $\phi_{{}^{+}{B}^{+}}*\phi_{\dot{A}}=g_{{}_{n,A,{n+1}}}\phi_{{}^{+}{{}_{n}A_{n+1}}^{{\epsilon}}},\qquad\phi_{{}^{-}{B}^{-}}*\phi_{\dot{A}}=g_{{}_{n,A,{n+1}}}\phi_{{}^{-}{{}_{n}A_{n+1}}^{-{\epsilon}}}.$
  • (2)
    If $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}$, then ${\mathcal{O}}_{A}=\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}})$ and
    $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ $=\sum\limits_{1\leq p<n+1}\!g_{{}_{n},A,p}\phi_{{}^{+}{{}_{n}\!A_{p}}^{{\epsilon}}},$
    $\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ $=\sum\limits_{1\leq p<n+1}\!g_{n,A,p}\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}},$
    $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ $=\sum\limits_{p>n+1}\!g_{n,A,p}\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}},$
    $\phi_{{}^{-}{B}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ $=\sum\limits_{p>n+1}\!g_{n,A,p}\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}.$
Proof.
Recall in this case $B=E^{\theta}_{n,n+1}+{\widehat{\gamma}}$ with ${\widehat{\lambda}}=\text{\rm ro}(B),{\widehat{\mu}}=\text{\rm co}(B)=\text{\rm ro}(A)$, and $\nu=\text{\rm co}(A)$.
(1) If $A\in\overset{\bullet}{\Xi}$, then $\text{ro}(A)_{n+1}\neq 0$ and $(\text{\rm rw}(\dot{A}),\text{\rm cw}(\dot{A}))=(\mu^{\bullet},\nu^{\bullet}).$ Since $\lambda_{n+1}=0$ and $\text{ro}(A)_{n+1}=\text{co}(B)_{n+1}=2$, it follows that $a_{n+1,n+1}=2$ and $a_{n+1,p}=0$, for all $p\neq n+1$. Thus, the $(n+1,p)$ entry of ${}_{h}\!A_{p}$ is negative, for all $p\neq n+1$. Hence, (6.0.1) becomes $e_{{B}}*e_{A}=g_{n,A,{n+1}}e_{{{}_{n}A_{n+1}}}$. Since the central entry of ${}_{n}A_{n+1}$ is 0, we have ${\mathcal{O}}_{{}_{n}A_{n+1}}=\check{\mathcal{O}}(^{+}{{}_{n}A_{n+1}}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{n}A_{n+1}}^{-{\epsilon}})$ splits into two orbits and $\text{sgn}({{}_{n}A_{n+1}}^{\text{\normalsize$\llcorner$}})=\text{sgn}(A^{\llcorner})$, we have in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
$(\phi_{{}^{+}{B}^{+}}+\phi_{{}^{-}{B}^{-}})*\phi_{\dot{A}}=g_{{}_{B,A,_{n}A_{n+1}}}(\phi_{{}^{+}{{}_{n}A_{n+1}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{n}A_{n+1}}^{-{\epsilon}}}).$
Multiplying idempotents $1_{\lambda^{+}}$ or $1_{\lambda^{-}}$ to the left hand side gives (1).
(2) If $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}$, then the central entry of ${}_{n}A_{n+1}$ is $-2$. Thus, the term $e_{{}_{n}A_{n+1}}$in (6.0.1) is omitted. For every other term ${}_{h}\!A_{p}$, its central entry is 0. Thus, the corresponding $G$-orbits are all split. Applying (6.1.3) to (6.0.1) yields the following formulas in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
$(\phi_{{}^{+}{B}^{+}}+\phi_{{}^{-}{B}^{-}})*$ $(\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{A}^{-{\epsilon}}})=\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{B}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}+\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ (6.4.1)
$=\sum_{1\leq p<n+1}g_{n,A,p}(\phi_{{}^{+}{{}_{h}\!A_{p}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{p}}^{-{\epsilon}}})+\sum_{p>n+1}g_{n,A,p}(\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}),$
If $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$, then both sides have the row weight $\lambda^{+}$ or $\lambda^{-}$ and the column weight $\nu^{+}$ or $\nu^{-}$. Multiplying idempotents $1_{\lambda^{\pm}}$ and $1_{\nu^{\pm}}$ to left and right hand sides of (6.4.1), respectively, gives (2) in this case.
Finally, let $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$. Then both sides still have the row weight $\lambda^{+}$ or $\lambda^{-}$, but a sole column weight $\nu^{\bullet}$. We cannot use only weights to separate them. However, we may combined it with the Claim in the proof of Theorem 6.2(2).
Multiplying $1_{\lambda^{+}}$ to the left hand side of (6.4.1) yields that $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ is a linear combination of $\phi_{{\check{\mathbb{A}}}^{\prime}}$ with
${\check{\mathbb{A}}}^{\prime}\in\{{{}^{+}{{}_{h}\!A_{p}}^{{\epsilon}}}|p\in[1,n+1]\}\cup\{{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}|p>n+1\}.$
By the Claim in the proof of Theorem 6.2, $\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ is a linear combination of $\phi_{{\check{\mathbb{A}}}^{\prime}}$ with
${{\check{\mathbb{A}}}^{\prime}}\in\{{{}^{+}{{}_{h}\!A_{p}}^{{\epsilon}}}|p\in[1,n+1]\}\cup\{{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}|p>n+1\}.$
Hence, we have
$\phi_{{}^{+}{B}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}=\sum\limits_{1\leq p<n+1}\!g_{h,A,p}\phi_{{}^{+}{{}_{h}\!A_{p}}^{{\epsilon}}},$
and, consequently, we have
$\phi_{{}^{+}{B}^{+}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=\sum\limits_{p>n+1}\!g_{h,A,p}\phi_{{}^{+}{{}_{h}\!A_{p}}^{-{\epsilon}}}.$
The case for $\phi_{{}^{-}{B}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ and $\phi_{{}^{-}{B}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ can be proved similarly. ∎
7. Multiplication formulas in the ${\boldsymbol{q}}$-Schur algebra $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$, II
In this section, we derive the multiplication in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ formulas arising from the one in Theorem 5.2(2):
$e_{E_{h+1,h}^{\theta}+{\widehat{\gamma}}}*e_{A}=\delta_{{\widehat{\gamma}},\text{\rm ro}(A)-e^{\theta}_{h}}\sum\limits_{1\leqslant p\leqslant N}g^{\prime}_{{}_{h,A,\overline{p}}}e_{{}_{h}\!A_{\overline{p}}},$ (7.0.1)
where $h\in[1,n]$, $\gamma\in\Lambda(n+1,r),A=(a_{i,j})\in\Xi_{N,2r}$, ${}_{h}\!A_{\overline{p}}=A-E_{h,p}^{\theta}+E^{\theta}_{h+1,p}$, for all $p\in[1,N]$, and
$g^{\prime}_{{}_{h,A,\overline{p}}}=\begin{cases}{\boldsymbol{q}}^{\sum_{j<p}a_{h+1,j}}\left[\!\left[a_{h+1,p}+1\right]\!\right],&\text{ if $h<n$ or }h=n,p\neq n+1\text{ and }a_{h,p}>0;\\ {\boldsymbol{q}}^{\sum\limits_{j<n+1}a_{n+1,j}}\big{(}\left[\!\left[a_{n+1,n+1}+1\right]\!\right]+{\boldsymbol{q}}^{\frac{a_{n+1,n+1}}{2}}\big{)},&\text{ if}h=n,p=n+1,\text{ and }a_{n,n+1}>0;\\ 0,&\text{ if }a_{h,p}=0.\end{cases}$ (7.0.2)
In (7.0.1), let $C=E_{h+1,h}^{\theta}+{\widehat{\gamma}}$ and assume $\lambda=\text{ro}(B)$ and $\mu=\text{ro}(A)=\text{co}(B)$. Then the upper right corner matrix $C^{\text{\normalsize$\llcorner$}}=0$ (and so, $\text{sgn}(C^{\text{\normalsize$\llcorner$}})=+$). Thus, by Proposition 4.5, we have
${\mathcal{O}}_{C}=\begin{cases}\check{\mathcal{O}}(^{+}{C}^{+})\sqcup\check{\mathcal{O}}(^{-}{C}^{-}),&\text{ if }c_{n+1,n+1}=0,\\ \qquad\check{\mathcal{O}}(\dot{C}),&\text{ otherwise}\end{cases}\quad\text{ and }\quad\begin{cases}C\in\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet}{\Xi},&\text{ if }h<n,\\ C\in\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet}{\Xi},&\text{ if }h=n.\end{cases}$
Since ${}_{h}\!A_{\overline{p}}=A-E_{h,p}^{\theta}+E^{\theta}_{h+1,p}$, it follows that
$\text{sgn}({}_{h}\!A_{\overline{p}}^{\text{\normalsize$\llcorner$}})=\begin{cases}\text{sgn}(A^{\text{\normalsize$\llcorner$}}),\ &\text{ if $h<n$ or $h=n$ and $1\leq p\leq n+1$},\\ -\text{sgn}(A^{\text{\normalsize$\llcorner$}}),\ &\text{ if $h=n$ and $p>n+1$}.\end{cases}$ (7.0.3)
If $a_{n+1,n+1}=0$ and ${\epsilon}=\text{sgn }(A^{\text{\normalsize$\llcorner$}})$, then we have $G(q)$-orbits $\mathcal{O}_{A},$ $\mathcal{O}_{{}_{h}\!A_{\overline{p}}}$ splitting into ${\check{G}}(q)$-orbits:
$(1)\quad\mathcal{O}_{A}$ $=\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}}),$ (7.0.4)
$(2)\ \mathcal{O}_{{}_{h}\!A_{\overline{p}}}$ =$=\begin{cases}\check{\mathcal{O}}(^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}),&\text{ if $h<n$ or $h=n$ and }1\leq p\leq n+1,\\ \check{\mathcal{O}}(^{+}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}),&\text{ if $h=n$ and}p>n+1.\end{cases}$
7.1. The $C\in\overset{\bullet}{\Xi}$ case.
Since ${\mathcal{O}}_{C}=\check{\mathcal{O}}(\dot{C})$ in this case and $\text{\rm cw}(\dot{C})=\mu^{\bullet}\in\Lambda^{\bullet}(n+1,r)$, it suffices to consider those $A$ in (7.0.1) with $A\in\overset{\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$. The following result is the counterpart of Theorem 6.2. Note that the two cases $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$ and $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$ are quite different from the corresponding cases there.
Theorem 7.1.
Let $A,C\in\Xi_{N,2r}$ with $C=E_{h+1,h}^{\theta}+{\widehat{\gamma}}$, and assume ${\widehat{\lambda}}=\text{\rm ro}(C)$, ${\widehat{\mu}}=\text{\rm co}(C)=\text{\rm ro}(A)$, and ${\widehat{\nu}}=\text{\rm co}(A)$, where $h\in[1,n]$, and $\gamma,\lambda,\mu,\nu\in\Lambda(n+1,r).$ Let ${\epsilon}=\text{sgn }(A^{\text{\normalsize$\llcorner$}})$. If $C\in\overset{\bullet}{\Xi}$, then ${\mathcal{O}}_{C}=\check{\mathcal{O}}(\dot{C})$ and the following multiplication formulas hold in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$:
  • (1)
    If $A\in\overset{\bullet}{\Xi}$, then ${\mathcal{O}}_{A}=\check{\mathcal{O}}(\dot{A})$ and $\phi_{\dot{C}}*\phi_{\dot{A}}=\sum_{1\leqslant p\leqslant N}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}_{h}\!{\dot{A}}_{\overline{p}}}.$
  • (2)
    If $A\in\overset{\bullet\!\circ\!\circ}{\Xi}\sqcup\overset{\bullet\!\circ\!\bullet}{\Xi}$, then
    $\phi_{\dot{C}}*\phi_{{}^{+}{A}^{{\epsilon}}}=$$\begin{cases}\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}},&\text{ if }h<n,\\ \sum\limits_{1\leq p<n+1}\!g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}+\frac{1}{2}{g^{\prime}_{n,A,{\overline{n+1}}}}\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}+\!\sum\limits_{p>n+1}\!g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}},&\text{ if }h=n,\end{cases}$
    $\phi_{\dot{C}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=$$\begin{cases}\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{-}{({}_{h}\!A_{p})}^{-\epsilon}},&\text{ if }h<n,\\ \sum\limits_{1\leq p<n+1}\!g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}}+\frac{1}{2}{g^{\prime}_{n,A,{\overline{n+1}}}}\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}+\sum\limits_{p>n+1}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}},&\text{ if }h=n.\end{cases}$
Here all coefficients $g^{\prime}_{{}_{h,A,\overline{p}}}$ are given in (7.0.2).
Proof.
(1) Unlike the case in Theorem 6.2(1), we have all ${}_{h}\!A_{\overline{p}}\in\overset{\bullet}{\Xi}$ in this case. Thus, its proof is almost identical to the first paragraph of the corresponding proof there.
(2) If $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}\sqcup\overset{\bullet\!\circ\!\circ}{\Xi}$, then ${}_{n}A_{\overline{n+1}}=A-E^{\theta}_{n,n+1}+E^{\theta}_{n+1,n+1}$ has 2 at the central entry and a negative $(n,n+1)$ entry if $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$. All other ${}_{h}\!A_{\overline{p}}$ has a 0 central entry. Thus, applying (7.0.4) to (7.0.1) yields
$\phi_{\dot{C}}*(\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{A}^{-{\epsilon}}})=\begin{cases}\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}}),&\text{ if }h<n,\\ \sum\limits_{1\leq p<n+1}\!g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}})+g^{\prime}_{n,A,\overline{n+1}}\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}&\\ \qquad\qquad\qquad+\sum\limits_{p>n+1}\!g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}+\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}}),&\text{ if }h=n.\end{cases}$ (7.1.1)
If $A\in\overset{\bullet\!\circ\!\circ}{\Xi}$, then $g^{\prime}_{n,A,\overline{n+1}}=0$. Thus, this case is proved by using an idempotent argument as in the proof of Theorem 6.2(2).
The case for $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$ is a bit more complicated just like its counterpart in Theorem 6.2. First, we need to modify the Claim there as follows:
  • Claim
    1.   For $h\in[1,n]$ and $p\in[1,N]$, if $F,F^{\prime},E\in\mathcal{F}_{n,r}^{\jmath}$ satisfy ${\mathfrak{m}}({F},E)=C$, ${\mathfrak{m}}(E,F^{\prime})=A$ and ${\mathfrak{m}}({F},F^{\prime})={}_{h}\!A_{\overline{p}}$, then
    $(E,F^{\prime})\in\mathcal{\check{O}}(^{{\epsilon}_{1}}\!A^{{\epsilon}_{2}})\Longrightarrow(F,F^{\prime})\in\begin{cases}\mathcal{\check{O}}(^{{\epsilon}_{1}}\!{{}_{h}\!A_{\overline{p}}}^{{\epsilon}_{2}}),&\text{ if $h<n$ or $h=n$ and $p\leq n$};\\ \mathcal{\check{O}}({{}_{n}{\dot{A}}_{\overline{n+1}}}),&\text{ if $h=n$, $p=n+1$ (and $a_{n,n+1}\geq 1$)};\\ \mathcal{\check{O}}(^{-{\epsilon}_{1}}\!{{}_{h}\!A_{\overline{p}}}^{{\epsilon}_{2}}),&\text{ if $h=n$ and $p\geq n+2$}.\end{cases}$
The proof of the claim is similar to that of the Claim in subsection 6.1 with the roles of $F$ and $E$ swapped.
Next, we need to split the coefficient $g^{\prime}_{n,A,\overline{n+1}}$ in (7.1.1).
  • Claim
    2.   For $(F,F^{\prime})\in{\mathcal{O}}_{{}_{n}\!A_{\overline{n+1}}}$, the set =${\mathcal{F}}=\{E\in{\mathcal{F}}^{\jmath}_{n,r}\mid(F,E)\in{\mathcal{O}}_{C},(E,F^{\prime})\in{\mathcal{O}}_{{}_{n}\!A_{\overline{n+1}}}\}$, which defines $g^{\prime}_{n,A,n+1}$, is a disjoint union $^{\mathcal{F}}={\mathcal{E}}\sqcup\hat{\mathcal{E}}$ of two subsets with equal cardinality, where
    ${\mathcal{E}}$ =$=\{E\in{\mathcal{F}}^{\jmath}_{n,r}\mid(F,E)\in\check{\mathcal{O}}(\dot{C}),(E,F^{\prime})\in\check{\mathcal{O}}({}^{+}\!A^{\epsilon})\},$
    ${\mathcal{E}}^{\prime}$ =$=\{E\in{\mathcal{F}}^{\jmath}_{n,r}\mid(F,E)\in\check{\mathcal{O}}(\dot{C}),(E,F^{\prime})\in\check{\mathcal{O}}({}^{-}\!A^{-{\epsilon}})\}.$
Indeed, the disjoint union is clear. To see the assertion of equal cardinality, we construct a bijection
$f:{\mathcal{E}}\longrightarrow{\mathcal{E}}^{\prime},\;\;\;E\longmapsto E^{\prime}.$
We now follow the notation used in Proposition 3.8. Thus, $F_{i}=\langle v_{1},\ldots,v_{\widetilde{\alpha}_{i}}\rangle$ $(\alpha={\widehat{\lambda}})$ and $F^{\prime}=F^{A^{\prime}}$, where $A^{\prime}=_{n}\!A_{\overline{n+1}}$ has central entry 2. For $E\in{\mathcal{F}}$, by Corollary 5.3, we have $E_{n}=F_{n}+\langle v\rangle$, where $v$ has the form $v=c_{1}v_{\widetilde{\lambda}_{n}+1}+\cdots+c_{m}v_{\widetilde{\lambda}_{n}+m}+c_{r}v_{r}+c_{r+1}v_{r+1}$ with $m=a_{n+1,1}+\cdots+a_{n+1,n}$ and either $c_{r}=0$ or $c_{r+1}=0$, but not both 0.
Suppose now $E\in{\mathcal{E}}$ with $c_{r}\neq 0$ (we may assume $c_{r}=1$) and define $E_{n+1}$ by removing $v_{r+1}$ from $F_{n+1}$, and $E_{i}=F_{i}$, for all $i\neq n,n+1$. Then $E_{n}=F_{n}+\langle v\rangle\subset E_{n+1}$. Note that $E$ is an $n$-step isotropic flag since
$E_{n}^{\perp}=F_{n}^{\perp}\cap\langle v\rangle^{\perp}=F_{n+1}\cap\langle v_{k},v\mid k\in[1,N]\setminus\{r,r+1\}\rangle=\langle v_{k},v\mid k\in[1,\widetilde{\alpha}_{n+1}]\setminus\{r+1\}\rangle=E_{n+1}.$
Let $v^{\prime}=c_{1}v_{\widetilde{\lambda}_{n}+1}+\cdots+c_{m}v_{\widetilde{\lambda}_{n}+m}+v_{r+1}$ and define $E^{\prime}\in{\mathcal{F}}$ by setting
$E^{\prime}_{n}=F_{n}+\langle v^{\prime}\rangle,\;E^{\prime}_{n+1}=\langle v_{k},v^{\prime}\mid k\in[1,\widetilde{\alpha}_{n+1}]\setminus\{r\}\rangle,\;\text{ and }\;E^{\prime}_{i}=F_{i},\;\forall i\neq n.$
Then one checks easily that $E^{\prime}$ is $n$-step isotropic and $E^{\prime}\in{\mathcal{E}}^{\prime}$. It is clear that the map $E\mapsto E^{\prime}$ is bijective, proving Claim 2.
By Claim 1, $\phi_{\dot{C}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ is a linear combination of $\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}$ if $h<n$, or a linear combination of $\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}$, $\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}$ and $\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}$ if $h=n$. By Claim 2, the coefficient $g^{\prime}_{n,A,n+1}$ is halved. The argument for $\phi_{\dot{C}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ is similar. The proof is complete. ∎
7.2. The case $C\in\overset{\circ\!\circ\!\circ}{\Xi}$ ($\implies h<n$).
In this case, ${\mathcal{O}}_{C}=\check{\mathcal{O}}(^{+}{C}^{+})\sqcup\check{\mathcal{O}}(^{-}{C}^{-}).$ Enough to consider those $A$ with $A\in\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}.$ This case is entirely similar to Theorem 6.3. We omit the proof.
Theorem 7.2.
Maintain the same assumptions on $A,C,h,\gamma,\lambda,\mu,\nu,$ and ${\epsilon}=\text{sgn }(A^{\text{\normalsize$\llcorner$}})$ as in Theorem 7.1, and assume $C=E_{h+1,h}^{\theta}+\check{\gamma}\in\overset{\circ\!\circ\!\circ}{\Xi}$. Then $h<n$, and, for $A\in\overset{\circ\!\circ\!\bullet}{\Xi}\sqcup\overset{\circ\!\circ\!\circ}{\Xi}$, the following multiplication formulas hold in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$:
$\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ $=\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{+}\!{({}_{h}\!A_{\overline{p}})}^{{\epsilon}}},\qquad\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}=\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{-}\!{({}_{h}\!A_{\overline{p}})}^{-{\epsilon}}}.$
7.3. The case $C\in\overset{\bullet\!\circ\!\circ}{\Xi}$ ($\implies h=n$).
We have in this case ${\mathcal{O}}_{C}=\check{\mathcal{O}}(^{+}{C}^{+})\sqcup\check{\mathcal{O}}(^{-}{C}^{-})$ with $(\text{rw}(^{+}{C}^{+}),\text{cw}(^{+}{C}^{+}))=(\lambda^{\bullet},\mu^{+})$ and $(\text{rw}(^{-}{C}^{-}),\text{cw}(^{-}{C}^{-}))=(\lambda^{\bullet},\mu^{-})$. Thus, enough to consider those $A$ with $A\in\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}.$
Theorem 7.3.
Let $A,C\in\Xi_{N,2r}$ and ${\epsilon}=\text{\rm sgn }(A^{\text{\normalsize$\llcorner$}})$, and assume $\text{\rm co}(C)=\text{\rm ro}(A)$, $C\in\overset{\bullet\!\circ\!\circ}{\Xi}$, and $C-E^{\theta}_{n+1,n}$ is diagonal. Then, for $A\in\overset{\circ\!\circ\!\circ}{\Xi}\sqcup\overset{\circ\!\circ\!\bullet}{\Xi}$, $\mathcal{O}_{A}=\check{\mathcal{O}}(^{+}{A}^{{\epsilon}})\sqcup\check{\mathcal{O}}(^{-}{A}^{-{\epsilon}})$ and the following multiplication formulas hold in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$:
$\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ $=\sum\limits_{1\leq p<n+1}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{+}\!{({}_{h}\!A_{\overline{p}})}^{{\epsilon}}}+\frac{1}{2}{g^{\prime}_{n,A,\overline{n+1}}}\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}+\sum\limits_{p>n+1}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{-}\!{({}_{h}\!A_{\overline{p}})}^{{\epsilon}}},\quad$
$\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ $=\sum\limits_{1\leq p<n+1}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{-}\!{({}_{h}\!A_{\overline{p}})}^{-{\epsilon}}}+\frac{1}{2}{g_{n,A,{\overline{n+1}}}}\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}+\sum\limits_{p>n+1}g^{\prime}_{{}_{h,A,\overline{p}}}\phi_{{}^{+}\!{({}_{h}\!A_{\overline{p}})}^{-{\epsilon}}}.$
Proof.
If $A\in\overset{\circ\!\circ\!\circ}{\Xi}$, then $(\text{rw}(^{+}{A}^{{\epsilon}}),\text{cw}(^{+}{A}^{{\epsilon}}))=(\mu^{+},\nu^{\epsilon})$, $(\text{rw}(^{-}{A}^{-{\epsilon}}),\text{cw}(^{-}{A}^{-{\epsilon}}))=(\mu^{-},\nu^{-{\epsilon}})$, and the central (or $(n+1)$th) component of $\text{\rm co}(A)$ is 0. Thus, the $(n,n+1)$ entry of ${}_{n}{A}_{\overline{n+1}}$ is negative. Hence, $\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}$ does not occur in this case and so we will omit the $p=n+1$ term.
By (7.0.4)(2), (7.0.1) has the following decomposition in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
$(\phi_{{}^{+}{C}^{+}}+\phi_{{}^{-}{C}^{-}})*(\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{A}^{-{\epsilon}}})=\sum\limits_{1\leq p<n+1}g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}})+\sum\limits_{p>n+1}g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}).$ (7.3.1)
Clearly, LHS=$\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{-}{A}^{{\epsilon}}}=\phi_{{}^{-}{C}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}=0.$ Moreover, every natural basis element $\phi_{\check{\mathbb{A}}^{\prime}}$ on the right hand side satisfies cw$(\check{\mathbb{A}}^{\prime})=\nu^{\epsilon}$ or $\nu^{-{\epsilon}}$. Multiplying idempotents $1_{\nu^{\epsilon}}(\text{resp., }1_{\nu^{-{\epsilon}}})$ to the right hand sides of (7.3.1) yields the required formulas for $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}(\text{resp., }\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}})$.
If $A\in\overset{\circ\!\circ\!\bullet}{\Xi}$, then $(\text{rw}(^{+}{A}^{{\epsilon}}),\text{cw}(^{+}{A}^{{\epsilon}}))=(\mu^{+},\nu^{\bullet})$ and $(\text{rw}(^{-}{A}^{-{\epsilon}}),\text{cw}(^{-}{A}^{-{\epsilon}}))=(\mu^{-},\nu^{\bullet}).$ By (7.0.4)(2), (7.0.1) has the following decomposition in $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$,
$(\phi_{{}^{+}{C}^{+}}+\phi_{{}^{-}{C}^{-}})*(\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{A}^{-{\epsilon}}})=$ $\sum\limits_{1\leq p<n+1}g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}})$ (7.3.2)
$+{g^{\prime}_{n,A,{\overline{n+1}}}}\phi_{{}_{n}{\dot{A}}_{\overline{n+1}}}+\sum\limits_{p>n+1}g^{\prime}_{{}_{h,A,\overline{p}}}(\phi_{{}^{+}{{}_{h}\!A_{\overline{p}}}^{-{\epsilon}}}+\phi_{{}^{-}{{}_{h}\!A_{\overline{p}}}^{{\epsilon}}}),$
where the central entry of ${}_{h}\!A_{\overline{p}}$ is 00 except ${}_{n}{A}_{\overline{n+1}}$ which is $2$
Again, LHS=$\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}+\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ since $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{-}{A}^{{\epsilon}}}=\phi_{{}^{-}{C}^{-}}*\phi_{{}^{+}{A}^{{\epsilon}}}=0.$ We cannot only use weights to separate the rest two items. Instead, by Claim 1 and Claim 2 in the proof of Theorem 7.1(2)&(3), $\phi_{{}^{+}{C}^{+}}*\phi_{{}^{+}{A}^{{\epsilon}}}$ is a linear combination of $\phi_{\check{\mathbb{A}}^{\prime}}$ with
${\check{\mathbb{A}}^{\prime}}\in\{{{}^{+}{{}_{h}\!A_{p}}^{{\epsilon}}}\mid p\in[1,n]\}\cup\{{{}_{n}{\dot{A}}_{\overline{n+1}}}\}\cup\{{{}^{-}{{}_{h}\!A_{p}}^{{\epsilon}}}\mid p>n+1\},$
which gives the first equation in (2). The case for $\phi_{{}^{-}{C}^{-}}*\phi_{{}^{-}{A}^{-{\epsilon}}}$ can be similarly obtained. ∎
Remark 7.4.
In [14, Prop. 4.3.2], Z. Fan and the second-named author gave three sets of multiplication formulas in $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ which do not cover all the cases discussed in §§6–7. For example, the case when $A\in\overset{\bullet\!\circ\!\bullet}{\Xi}$ is not mentioned at all. Even for those presented there, the $h=n$ subcase is also missing.